### v1

# Phase Transition for Budgeted Multi-Agent Synergy

Abstract
Multi-agent systems can improve reliability, yet under a fixed inference budget they often help, saturate, or even collapse. We develop a deliberately minimal and calibratable theory that predicts these regimes from three binding constraints of modern agent stacks: finite context windows, lossy inter-agent communication, and shared failures among similar agents. Each leaf agent is summarized by a compute–performance scaling exponent $\beta$; communication is captured by a message-length fidelity curve $\gamma(m)$; dependence is captured by an effective shared-error correlation $\rho$; and a context window $W$ imposes hard fan-in limits that make hierarchy necessary. For binary success/failure tasks with majority aggregation, we prove a sharp phase transition for deep $b$-ary trees with correlated inputs and lossy communication: a single scalar $\alpha_{\rho}$ (combining $\gamma(m)$, $\rho$, and fan-in $b$) determines whether weak signal is amplified to a nontrivial fixed point or washed out to chance. In the amplifying regime, we derive an organization exponent $s$ and show that budgeted synergy, i.e., outperforming the best single agent under the same total budget, occurs exactly when $s>\beta$, yielding closed-form compute allocation rules and explicit budget thresholds. We further characterize saturation via a mixing depth and provide a conservative clipped predictor that remains accurate across growth and saturation. A continuous-performance warm-up gives closed-form risks for star, chain, and tree organizations, making correlation- and communication-induced floors explicit and exposing the core design trade-offs in a smooth setting. Finally, we validate the predicted phase boundaries in controlled synthetic simulations and show how the same mechanisms explain the dominant bottlenecks reported in recent large-scale matched-budget studies of LLM agent-system scaling, including context saturation, subcritical error cascades, and diminishing returns at strong baselines.
Keywords: Budgeted Multi-Agent Synergy, Phase Transition, Shared-failure Correlation, Communication Bottlenecks, Finite Context Windows
1 Introduction
Multi-agent systems are often sold as a reliability primitive: if a single agent is fallible, run many agents, let them interact, and aggregate their outputs. Under a fixed inference budget, however, multi-agent coordination is not reliably beneficial. The same additional agents that sometimes improve performance can also yield saturation or outright degradation. This paper develops a predictive theory of this brittleness: when multi-agent organization produces budgeted synergy (i.e., outperforming any single agent under the same budget) and when scaling out must fail.
The LLM era has made the phenomenon impossible to ignore. It is now routine to instantiate multiple LLM-based agents, assign roles, and coordinate them through discussion, critique, or voting in “agent societies” and orchestration toolkits (Brown et al., 2020; Yao et al., 2023; Wu et al., 2023; Li et al., 2023; Chen et al., 2024; Hong et al., 2024). The underlying intuition is classical: ensembling can reduce error (Breiman, 1996; Freund and Schapire, 1997; Dietterich, 2000), and interaction protocols such as debate or amplification aim to turn many weak judgments into a stronger one (Irving et al., 2018; Christiano et al., 2018; Du et al., 2024). At the same time, controlled matched-budget evaluations increasingly report strong task- and topology-dependence, including regimes where multi-agent variants degrade once coordination overhead and information loss dominate (Kim et al., 2025).
Our starting point is that these failures are not implementation accidents but structural consequences of three constraints that repeatedly bind in modern agent stacks. First, agent errors are not independent: many instances derived from the same base model and prompt share failure modes, producing “groupthink” rather than error cancellation. Second, coordination is communication-limited: as systems scale, agents must compress reasoning into bounded-length messages, and this loss compounds across layers. Third, coordination is context-limited: a central aggregator cannot read arbitrarily many messages under a finite context window, so naive scale-out saturates structurally. Recent work on large-population collaboration and topology learning highlights the same bottleneck: token and context limits are often the binding resource (Qian et al., 2025; Zhang et al., 2024).
Today, agent-system design still largely proceeds by heuristics: choose a topology, message format, and number of agents, then iterate. In spirit, this resembles the hand-crafted feature engineering era of machine learning: powerful components exist, but system-level behavior depends on ad hoc choices and expensive trial-and-error. Our goal is to push multi-agent design toward an “agentic intelligence from first principles” style: make the dominant constraints explicit, explain why scaling out works or fails, and derive sharp, quantitative regime boundaries—in particular, a phase transition for budgeted multi-agent synergy.
A minimal, calibratable coordination theory.
We develop a deliberately small framework that isolates the coordination bottlenecks above while remaining analytically tractable. The term agent is used abstractly: a black-box solver that, given a task instance and an allocated compute budget, produces an output that another component can consume (a decision, a score, or an $m$-token message). This deliberately includes LLM-based agents as a primary motivating instance, because (i) tokens provide a natural budget unit, (ii) finite context windows induce hard fan-in constraints, and (iii) single-agent capability is often well described by compute scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022). At the same time, the abstraction is broader than LLMs: none of the results relies on language as such, only on (a) how single-agent quality scales with compute, (b) how much usable information survives bounded-length communication, and (c) how strongly agents’ failures co-move.
By calibratable we do not mean a theory “about measurement.” We mean a coordination theory parameterized by a small set of environment-dependent effective quantities that can be estimated once (via targeted experiments) and then used to predict qualitative regimes. Concretely, the framework is driven by single-agent scaling (exponent $\beta$), communication fidelity as a function of message length (e.g., $\gamma(m)$), shared-failure correlation ($\rho$), and the context window constraint ($W$), together with a total budget $B$ that fixes the scale-up/scale-out trade-off. These parameters are not claimed to capture every detail of a real agent stack; they form a minimal interface for diagnosing which constraint is binding and for predicting when organization helps, saturates, or fails.
Phase transition for budgeted multi-agent synergy.
Within this framework we analyze three canonical organizations: star (central aggregation), chain (sequential handoff), and hierarchical tree. Our main results characterize when hierarchy is viable and when it is provably counterproductive. For binary success/failure tasks with majority aggregation, we prove a sharp amplification–collapse transition for deep trees: a scalar $\alpha_{\rho}$ (combining communication fidelity, correlation, and fan-in) determines whether weak signal amplifies across layers or is washed out to chance, in the spirit of reconstruction thresholds in noisy broadcasting on trees (Evans et al., 2000; Mossel, 2001; Kesten and Stigum, 1966). In the amplifying regime, we derive an organization exponent $s$ governing growth with the number of leaves and show that budgeted synergy in the growth range occurs exactly when $s>\beta$, yielding closed-form compute allocation and explicit budget thresholds. We also characterize finite-depth saturation via a nontrivial fixed point and mixing-depth guarantees. Finally, we make precise why topology becomes decisive only because constraints bind: absent context limits, central aggregation is information-dominant (data processing inequality) (Cover and Thomas, 2006), and hierarchy is valuable only insofar as it bypasses fan-in limits without crossing into the collapsing regime.
Beyond the binary core, a continuous-performance warm-up yields closed-form risks for star/chain/tree organizations and makes correlation and communication floors explicit. We also translate theorems into design diagnostics (e.g., monotone communication design curves) and connect the predicted bottlenecks to recent controlled large-scale agent-scaling studies (Kim et al., 2025). Detailed calibration templates are collected in Appendix J.
Methodological stance: theory vs. benchmarking.
This is a theory paper: our aim is to delineate regimes of behavior and identify the mechanisms that govern when multi-agent systems help, saturate, or collapse under realistic constraints. Large-scale matched-budget evaluations with real LLM agent stacks are expensive and can be confounded by prompt engineering and rapidly changing model/tool ecosystems. Accordingly, we validate the sharp regime boundaries via controlled synthetic simulations (Section 7.2) that probe the theory under its own assumptions, and we use external controlled evidence (Kim et al., 2025) to show that the bottlenecks isolated by the theory (context saturation, subcritical collapse, and diminishing returns) are salient in modern deployments.
Organization.
Section 2 defines the framework, measurable parameters, and the budget and context constraints. Section 3 develops the continuous warm-up. Section 4 presents the binary theory, including the phase transition, small-signal amplification, and saturation. Section 5 compares topologies under context constraints and derives budget thresholds. Section 6 turns the theory into design diagnostics. Section 7 provides empirical touchpoints via external controlled evidence and synthetic sanity checks. We place detailed proofs and optional calibration/evaluation templates in the appendices.
2 Framework and Measurable Parameters
This section formalizes the abstraction used throughout the paper. We keep only three constraints that repeatedly bind in practice: finite context windows, lossy communication, and shared failures, and summarize the environment by a small set of quantities. These quantities play distinct roles: $\beta$ captures how a single agent improves with compute, $\gamma(m)$ (or $\sigma_{c}^{2}(m)$) captures what survives an $m$-token message, and $\rho$ captures how correlated different agents’ errors are. Together with a total budget $B$ and a context window $W$, they determine whether scaling out amplifies signal, saturates at a floor, or collapses.
We model a multi-agent system as a directed acyclic computation graph. Leaf nodes (agents) produce signals about a latent task variable $Y$; internal nodes receive messages from their children (possibly through a lossy channel) and apply an aggregation rule; the root output is the system prediction. Throughout, an agent is a black-box solver with an adjustable compute budget $x$ that emits a bounded-length output (a decision, a score, or an $m$-token message). LLM-based agents are the primary motivating instance, but the framework only assumes a single-agent scaling curve, a communication-fidelity curve, and a shared-failure correlation.
2.1 Tasks and performance metrics
We study two task primitives that cover many MAS evaluations and isolate the coordination bottlenecks above. Binary success/failure captures accuracy-like outcomes (pass rates, accept/reject decisions). Continuous performance on $[0,1]$ captures graded scoring or calibrated estimates and provides a smooth warm-up where many expressions admit closed forms. The continuous setting is not a substitute for the binary theory; it makes correlation and communication floors explicit without the nonlinearity introduced by majority vote.
Binary tasks.
[…]
$\mu\;:=\;\mathbb{E}[\widehat{Y}Y]\in[-1,1],\qquad\Pr(\widehat{Y}=Y)=\frac{1+\mu}{2}.$
The bias $\mu$ is convenient for multi-layer analysis because it composes cleanly under simple channel models and aggregation maps.
Continuous tasks.
Let $Y\in[0,1]$ and a system output $\widehat{Y}\in[0,1]$. We use mean squared error (MSE)
$v\;:=\;\mathbb{E}\big[(\widehat{Y}-Y)^{2}\big],$
and occasionally report a bounded performance score such as $\mathrm{Perf}=1/(1+v)$, which is monotone in $v$. The continuous analysis yields explicit floors; the binary analysis additionally introduces a sharp amplification–collapse transition.
2.2 Single-agent capability and scaling
Our central comparison is fixed-budget: the same total budget $B$ can be spent on fewer stronger agents (scale up) or more weaker agents (scale out). We model per-leaf compute by a single knob $x\geq 0$. In an LLM agent, $x$ can represent a token budget for reasoning, the number of samples or self-consistency votes, the number of tool calls, or any additive resource that trades off against parallelism.
We summarize the effect of compute via a single-agent scaling law, motivated by empirical power-law relations between performance and compute observed for large language models and related systems (Kaplan et al., 2020; Hoffmann et al., 2022). For binary tasks, a leaf with compute $x$ produces a vote $\widehat{Y}$ with bias
$\mu_{0}(x)=g(x),$
where $g(\cdot)$ is increasing and typically exhibits diminishing returns. For continuous tasks, a leaf produces an estimate $X$ with conditional variance $v_{0}(x)$, where $v_{0}(\cdot)$ decreases with $x$.
In our main theorems, we work in a regime where these functions are well-approximated by power laws:
Binary: $g(x)\approx k\,x^{\beta}\quad\text{(small-signal regime)},$ (1)
Continuous: $v_{0}(x)\approx c\,x^{-\beta},$ (2)
with constants $k,c>0$ and a single-agent scaling exponent $\beta>0$. The exponent $\beta$ is not a universal constant. It depends on the model, prompting, tools, and the task family, and it is intended to be measured.
2.3 Communication as a controllable lossy channel
Agents coordinate through messages. Our analysis uses a compact abstraction of how message length affects what downstream nodes can reliably use.
We model each edge as a channel controlled by a message length parameter $m$ (in tokens). Longer messages generally preserve more usable information, but they also cost more budget and reduce feasible fan-in under a finite context window.
Binary: an effective bit channel.
For binary tasks and majority-style protocols, it is natural to treat each child as intending to transmit a single bit of information (its vote or decision), encoded into a message of length $m$. We model the receiver’s decoded bit as passing through a binary symmetric channel (BSC) with reliability $\gamma(m)\in(0,1]$: if a child has bias $\mu$ relative to $Y$, then after transmission the received bit has bias $\gamma(m)\mu$. Equivalently, $\gamma(m)=1-2\epsilon(m)$ where $\epsilon(m)$ is an effective flip probability.
Continuous: additive distortion.
For continuous tasks, we model transmission as additive zero-mean distortion,
$\widetilde{X}=X+\eta,\qquad\mathbb{E}[\eta]=0,\quad\mathrm{Var}(\eta)=\sigma_{c}^{2}(m),$
where $\sigma_{c}^{2}(m)$ decreases with $m$. This captures compression loss, imperfect interpretation, and degradation introduced by summarization or constrained prompts.
The intent is not to claim that real messages are literally bits or Gaussian perturbations, but to represent an effective fidelity curve as a function of $m$. In particular, $\gamma(m)$ or $\sigma_{c}^{2}(m)$ can be estimated by simple encode–decode experiments at the same message length.
2.4 Shared-failure correlation and the $\rho$ model
A key departure from idealized ensemble analyses is that agent errors are not independent, even though many ensemble gains are easiest to analyze under weak dependence assumptions (Dietterich, 2000; Germain et al., 2015). Agents instantiated from the same base model and prompt often share systematic error modes, which can dominate any aggregation gain. We capture this effect with a single, measurable correlation parameter $\rho\in[0,1)$.
Binary: correlation of signed correctness.
For a leaf vote $\widehat{Y}_{i}$, define the signed correctness variable
$S_{i}:=\widehat{Y}_{i}Y\in\{-1,+1\}.$
Then $\mathbb{E}[S_{i}]=\mu_{0}$, and we define
$\rho\;:=\;\mathrm{Corr}(S_{i},S_{j}),\qquad i\neq j,$
estimated empirically by averaging pairwise correlations across tasks and agent instances. Here $i$ and $j$ index two agent instances whose outputs are aggregated together: for a star this is any pair of leaves; for a tree it should be interpreted as the sibling correlation at an internal node. In practice the effective correlation can depend on depth, heterogeneity, or the message schema; for tractability we treat $\rho$ as a depth-independent effective parameter and return to extensions in Section 8. Intuitively, $\rho=0$ corresponds to independent errors, while $\rho$ close to 1 indicates near-perfect groupthink.
For analysis, we use a concrete generative model that matches this definition exactly. At each aggregation node, with probability $\rho$ the children share a common “mode” (their signed correctness variables are identical), and with probability $1-\rho$ they are conditionally independent given $Y$. This model produces the same pairwise correlation $\rho$ while remaining analytically tractable and yields a closed-form correlated aggregation map used throughout the paper.
Continuous: correlated residuals.
For continuous outputs $X_{i}$, define residuals $E_{i}:=X_{i}-Y$. We use the standard equal-correlation model
$\mathrm{Var}(E_{i})=v,\qquad\mathrm{Cov}(E_{i},E_{j})=\rho v,\quad i\neq j,$
which makes correlation floors explicit and leads to closed-form recursions in Section 3. The same $\rho$ can be estimated by pairwise residual correlations.
Why a single parameter is useful.
Real systems may have richer dependence structure than a single coefficient. The intent here is not to model all dependence, but to isolate the dominant failure direction: shared errors. As we show later, this single parameter already shifts phase boundaries and imposes aggregation floors that are otherwise difficult to diagnose.
2.5 Budget and context constraints
Our central comparison is fixed-budget: a multi-agent system should be judged against the best single-agent baseline under the same total cost. We therefore make computation and communication costs explicit.
Budget.
Let $B$ denote the total available budget. In token-based LLM systems, a natural accounting unit is tokens, including both generated tokens and tokens processed as input. Our theory is compatible with other budget units (time, API cost, calls), as long as they are additive and comparable across designs.
We separate two costs:
  • Leaf computation: each leaf agent is allocated compute $x$.
  • Communication: each edge transmits a message of length $m$ tokens.
To keep the model simple, we account a per-edge cost proportional to $m$. A useful approximation in token-based systems is
$\text{edge cost}\approx 2m,$
representing $m$ tokens produced by the sender and $m$ tokens read by the receiver.
Context window as a fan-in constraint.
Let $W$ denote the maximum number of tokens that an internal node can read and process. If each incoming message has length $m$, then any aggregation node can read at most
$\text{fan-in}\leq\Big\lfloor\frac{W}{m}\Big\rfloor.$
This inequality is the reason depth matters: a star is capped by $N\leq\lfloor W/m\rfloor$ leaves, while a tree can scale to $N=b^{L}$ leaves by distributing the fan-in constraint across levels.
Topology-specific cost summaries.
Let $N$ be the number of leaves. Under the accounting above:
  • Star: $C_{\text{star}}(N,x,m)\approx N(x+2m)$, with feasibility constraint $Nm\leq W$.
  • Chain: for length $L$, $C_{\text{chain}}(L,x,m)\approx(L+1)x+2mL$ (assuming each handoff step runs an agent with compute x, e.g., one model call), with feasibility constraint $m\leq W$ (usually nonbinding).
  • Tree: for a full $b$-ary tree of depth $L$ with $N=b^{L}$ leaves, the number of edges is
    $E=b\cdot\frac{N-1}{b-1},$
    so $C_{\text{tree}}(b,L,x,m)\approx Nx+2mE$. It is convenient to write this as a per-leaf cost:
    $C_{\text{tree}}(b,L,x,m)\approx N\big(x+c_{0}(b,m)\big),\qquad c_{0}(b,m):=2m\frac{b}{b-1},$
    subject to the fan-in feasibility $bm\leq W$. Alternative accounting (templating overheads, parsing costs, asymmetric protocols) rescales $c_{0}(b,m)$ and can be treated as an implementation constant.
This approximation isolates the basic trade-off: increasing $m$ can improve fidelity, but it also increases coordination cost and reduces the number of leaves affordable under $B$.
2.6 Topologies and protocols
A topology specifies how information flows; a protocol specifies what each node does with the information it receives. We keep protocols deliberately simple to make the effect of constraints and scaling transparent. In particular, we focus on three canonical organizations that recur in practice and isolate the key trade-offs: star, chain, and hierarchical tree (Figure 1).
[…]
Figure 1: Canonical organizations under a context window constraint $W$. A star requires the root to read all messages, so feasibility enforces $Nm\leq W$. A hierarchical tree enforces the constraint locally ($bm\leq W$), enabling many more leaves at the cost of multi-hop communication and repeated aggregation. Chains avoid fan-in bottlenecks but compound communication loss.
Star.
Leaves transmit their outputs directly to a central aggregator. Star avoids multi-hop loss, but it concentrates fan-in at one node and therefore saturates at $N\approx W/m$ leaves regardless of budget.
Chain.
Outputs are passed sequentially through $L$ steps. Chains avoid fan-in bottlenecks but repeatedly retransmit information; without new evidence injected along the path, each step can only preserve or degrade the signal.
Hierarchical tree.
A tree distributes aggregation across many small nodes. Each internal node aggregates $b$ child messages and forwards a summary upward, enforcing $bm\leq W$ locally and enabling many more leaves. The cost is multi-hop communication and repeated aggregation, so performance hinges on whether amplification dominates compounding loss and correlation.
Why focus on star, chain, and tree?
These motifs capture the dominant structural trade-offs induced by finite context and imperfect coordination. Star minimizes depth but is capped by a global fan-in constraint $Nm\leq W$. Chain removes fan-in limits but maximizes the number of lossy transmissions. Trees are the simplest family that bypasses the fan-in cap ($bm\leq W$ locally) while still aggregating many signals; this is precisely the setting in which phase transitions and budget thresholds emerge.
It is also useful to view more general feed-forward organizations through these motifs. With a finite context window, every node can only aggregate a bounded number of incoming messages, i.e., each local computation is an in-star with fan-in at most $\lfloor W/m\rfloor$. Globally, any connected directed acyclic communication graph contains a spanning arborescence that routes leaf information to the output; star and chain are the two extreme cases (depth $1$ versus fan-in $1$), and bounded-fan-in trees are the minimal structure that scales beyond the star’s context cap. Additional edges or multi-round interaction are valuable mainly when they effectively change the measurable parameters, for example by improving effective fidelity through redundancy or by reducing effective correlation through independent verification, rather than by rerouting the same information along multiple dependent paths.
Aggregation rules.
We analyze majority vote at internal nodes for binary tasks (ties are avoided by choosing odd fan-in) and averaging for continuous tasks (the MSE-optimal linear unbiased aggregator under equal-variance equal-correlation assumptions). More sophisticated protocols (critique, verification, debate) can be interpreted in our framework as mechanisms that effectively increase communication fidelity or reduce shared-error correlation. We return to this connection in the discussion.
2.7 What is measurable and how to measure it
The framework is parameterized by a small set of environment-dependent quantities. Table 1 summarizes their meaning and outlines how one might estimate them in a specific model–task setting. These estimates are optional: the theoretical results that follow hold given $\beta$, $\rho$, and the communication curves, while calibration is a way to connect the regime predictions back to concrete agent stacks at a coarse level.
Quantity Meaning in the framework How to estimate in practice
$\beta$ Single-agent scaling exponent: how fast a single agent improves with compute $x$. Sweep $x$ and fit a power law to bias $g(x)$ (binary) or MSE $v_{0}(x)$ (continuous) in the non-saturated regime.
$\gamma(m)$ / $\sigma_{c}^{2}(m)$ Communication fidelity as a function of message length $m$. Binary uses an effective bit reliability $\gamma(m)$; continuous uses a distortion variance $\sigma_{c}^{2}(m)$. Encode known labels/values into $m$ tokens and measure decode flip rate (binary) or MSE (continuous).
$\rho$ Shared-error correlation (groupthink strength) among agents within an aggregation group. Run multiple agents on the same tasks and estimate pairwise correlation of signed correctness $S_{i}=\widehat{Y}_{i}Y$ (binary) or residuals $E_{i}=X_{i}-Y$ (continuous).
$W$ Context window constraint controlling fan-in: each node can read at most $W$ tokens. System constraint minus templating/tool overhead; verify effective $W$ in the deployed scaffold.
$B$ Total budget (tokens, time, calls) to be allocated across agents and communication. Define a consistent additive accounting unit and record $B$ from logs.
Table 1: Key parameters of the framework and indicative estimation strategies. The theory treats these quantities as inputs; calibration is optional and necessarily coarse.
[…]
Figure 2: A possible use of the framework. Estimate $\beta$ (single-agent scaling), $\rho$ (shared failures), and $\gamma(m)$ or $\sigma_{c}^{2}(m)$ (communication fidelity) in a given setting, then use the theory to map budgets $B$ and context windows $W$ to plausible organizations and message lengths.
Sections 36 use these parameters to derive correlation/communication floors, a binary amplification–collapse phase transition, and budget-dependent topology comparisons.
3 Continuous Warm-up: Averaging Under Correlation and Communication Noise
We begin with a continuous-performance setting in which agents estimate a scalar target $Y\in[0,1]$ and performance is measured by mean squared error (MSE). This warm-up makes the basic organization trade-offs transparent in closed form: averaging reduces independent noise, shared errors impose floors, and communication loss introduces additional distortion. Throughout we adopt the framework of Section 2 and use averaging at internal nodes, which is the natural choice under squared loss and symmetric noise assumptions.
Although we phrase the warm-up as each leaf producing an estimate of the same latent $Y$, it should be read as an abstraction of role-specialized pipelines. A child agent may solve a subtask and transmit an $m$-token message; $X_{i}$ denotes the effective scalar contribution about the downstream target that the parent can decode from that message. Role heterogeneity then enters through the effective parameters ($\rho$ and $\sigma_{c}^{2}(m)$ or $\gamma(m)$); we return to richer semantics and extensions in Section 8.
3.1 Model assumptions for the continuous setting
Let $Y\in[0,1]$ denote the latent truth and let each leaf output an estimate $X\in[0,1]$. To keep the warm-up focused on organization rather than estimator bias, we assume leaf estimates are conditionally unbiased:
$\mathbb{E}[X\mid Y]=Y.$
We write the leaf residual as E:=XYE:=X-Y, so $\mathbb{E}[E\mid Y]=0$. Leaf quality is summarized by the conditional MSE,
$v_{0}(x):=\mathbb{E}\big[(X-Y)^{2}\big]=\mathbb{E}\big[E^{2}\big],$
which depends on per-leaf compute $x$. In the scaling regime of interest we model
$v_{0}(x)\approx c\,x^{-\beta},$ (3)
with constants $c>0$ and $\beta>0$, as introduced in Section 2.2.
Communication distortion.
An estimate sent across an edge as a message of length $m$ is received with additive distortion:
$\widetilde{X}=X+\eta,\qquad\mathbb{E}[\eta]=0,\quad\mathrm{Var}(\eta)=\sigma_{c}^{2}(m),$ (4)
where $\sigma_{c}^{2}(m)$ decreases with $m$. We treat the channel noise variables $\eta$ as independent across edges and independent of $Y$. This is a compact abstraction of imperfect summarization, constrained prompts, and the general fact that receivers do not reconstruct sender information perfectly.
Shared errors as residual correlation.
Within any aggregation group, child residuals are not assumed independent. Instead, we use an equal-correlation model: for children $i\neq j$ with residuals $E_{i},E_{j}$,
$\mathrm{Var}(E_{i})=v,\qquad\mathrm{Cov}(E_{i},E_{j})=\rho v,$ (5)
where $\rho\in[0,1)$ is the correlation coefficient. This matches the intended interpretation of $\rho$ as a measurable “shared-failure” strength: $\rho=0$ corresponds to independent residuals, while $\rho$ near 1 indicates strong groupthink.
Aggregation by averaging.
An internal node with fan-in $b$ receives $\widetilde{X}_{1},\dots,\widetilde{X}_{b}$ and outputs the average
$\widehat{Y}=\frac{1}{b}\sum_{i=1}^{b}\widetilde{X}_{i}.$ (6)
Under the symmetry assumptions above, this is the MSE-optimal linear unbiased aggregator.
3.2 Closed-form risk for star, chain, and tree
We now derive exact MSE expressions (or recursions) for the three canonical topologies. These formulas already reveal three key messages: star saturates under context constraints, chains compound loss, and trees trade scale for multi-hop noise.
Star.
Consider a star with $N$ leaves feeding a single aggregator. Each leaf sends a length-$m$ message, so the aggregator is feasible only if $Nm\leq W$. Let $v_{0}$ denote the leaf MSE (for a given compute allocation $x$) and let $\rho$ denote the pairwise residual correlation within this aggregation group. After one hop, each received estimate has residual $\widetilde{E}_{i}=(X_{i}-Y)+\eta_{i}$, with
$\mathrm{Var}(\widetilde{E}_{i})=v_{0}+\sigma_{c}^{2}(m),\qquad\mathrm{Cov}(\widetilde{E}_{i},\widetilde{E}_{j})=\rho v_{0}.$
The star output is the average of the received values, so the MSE is
$v_{\text{star}}(N)=\mathbb{E}\big[(\widehat{Y}-Y)^{2}\big]=\frac{v_{0}\big(1+(N-1)\rho\big)}{N}+\frac{\sigma_{c}^{2}(m)}{N}.$ (7)
If $N$ could grow arbitrarily, the channel noise term $\sigma_{c}^{2}(m)/N$ would vanish. However, correlation produces an irreducible floor:
$v_{\text{star}}(N)\ \to\ \rho v_{0}\qquad\text{as }N\to\infty\text{ (for fixed }v_{0}\text{)}.$
Thus, even perfect communication cannot eliminate a globally shared residual component. In practice, the context constraint $Nm\leq W$ prevents $N$ from growing with budget once $m$ is fixed, which is why star organizations often stop improving beyond a modest scale.
Chain.
Consider a chain of length $L$, where an estimate is handed off sequentially through $L$ communication steps, each of length $m$, without introducing new independent evidence. Starting from a leaf estimate with MSE $v_{0}$, each hop adds independent distortion variance $\sigma_{c}^{2}(m)$. Thus,
$v_{\text{chain}}(L)=v_{0}+L\,\sigma_{c}^{2}(m).$ (8)
Equation (8) isolates the unavoidable cost of multi-hop relay communication: if intermediate nodes primarily re-encode the same information (no new independent evidence), each hop injects distortion and MSE grows linearly in $L$. Pipeline-style chains can still be beneficial when each stage performs additional computation or incorporates new information; in that case, the estimation error may decrease across stages, but the same $\sigma_{c}^{2}(m)$ term captures a communication bottleneck that accumulates with depth and must be outweighed by per-stage gains.
Hierarchical tree.
Consider a full $b$-ary tree with depth $L$, where each internal node averages its $b$ children. Let $v_{t}$ denote the MSE of a node output at level $t$ (leaves are $t=0$; the root is $t=L$). We assume that at each aggregation node, the $b$ child residuals satisfy the equal-correlation model (5) with coefficient $\rho$, and that channel noise on incoming edges is independent.
A parent receives $\widetilde{X}_{i}=X_{i}+\eta_{i}$ from each child, so each received residual has variance $v_{t}+\sigma_{c}^{2}(m)$, while covariances remain $\rho v_{t}$ because the channel noises are independent. Averaging (6) yields the recursion
$v_{t+1}=a\,v_{t}+\frac{\sigma_{c}^{2}(m)}{b},\qquad a:=\frac{1+(b-1)\rho}{b}=\rho+\frac{1-\rho}{b}.$ (9)
Unrolling gives the closed form
$v_{L}=a^{L}v_{0}+\frac{\sigma_{c}^{2}(m)}{b}\cdot\frac{1-a^{L}}{1-a}.$ (10)
The coefficient $a\in(0,1)$ when $\rho<1$, so the first term $a^{L}v_{0}$ shrinks exponentially in depth. This is the statistical benefit of hierarchy: averaging repeatedly reduces the variance of what is being passed upward. The second term is the cost: each level injects fresh communication distortion, which accumulates into an error floor described next.
3.3 Saturation floors and mixing depth
Equation (10) exposes a sharp and practical phenomenon: deeper trees do not improve indefinitely. Even if leaves become arbitrarily accurate, repeated communication loss produces a floor. Figure 3 illustrates the typical behavior: $v_{L}$ shrinks rapidly with depth at first (roughly at rate $a^{L}$) and then saturates at the floor $v^{\star}$ determined by $\sigma_{c}^{2}(m)$, $b$, and $\rho$.
Tree floor.
For $\rho<1$ we have $a<1$, and therefore $a^{L}\to 0$ as $L\to\infty$. Taking limits in (10) yields
$v^{\star}(b,m,\rho):=\lim_{L\to\infty}v_{L}=\frac{\sigma_{c}^{2}(m)}{b(1-a)}=\frac{\sigma_{c}^{2}(m)}{(b-1)(1-\rho)}.$ (11)
This expression cleanly separates the levers available to a designer:
  • Increasing message length $m$ decreases $\sigma_{c}^{2}(m)$, lowering the floor.
  • Increasing fan-in $b$ lowers the floor approximately as 1/(b1)1/(b-1), but $b$ is limited by $bm\leq W$.
  • Reducing shared correlation $\rho$ lowers the floor and also accelerates convergence (since $a$ decreases).
Mixing depth.
The recursion (9) is linear, so convergence to the floor is explicit:
$v_{L}-v^{\star}=a^{L}\,(v_{0}-v^{\star}).$ (12)
Thus, to achieve a relative tolerance $v_{L}\leq(1+\varepsilon)v^{\star}$ when $v_{0}>v^{\star}$, it suffices to choose
[…]
Equation (13) is the continuous analogue of the “mixing depth” results we prove for binary trees later. It is also a direct design rule: beyond a certain depth, additional leaves yield negligible improvement unless the designer reduces the floor by improving communication or reducing correlation.
A note on when deeper is worse.
If leaves are already more accurate than the floor (i.e., $v_{0}<v^{\star}$), then (10) implies that increasing depth moves error upward toward $v^{\star}$. In that regime, deeper hierarchies are counterproductive unless they come with reduced $\sigma_{c}^{2}(m)$ or reduced $\rho$. This observation will reappear in the binary setting as a saturation effect: once the system has reached its fixed point, additional scale should be spent on changing the communication or correlation regime rather than adding depth.
depth LLMSE $v_{L}$$v^{\star}(b,m,\rho)$$v^{\star}(b,m^{\prime},\rho)$shorter messageslonger messagesexponential shrink until saturation
Figure 3: Conceptual behavior of the tree recursion (9). Increasing depth reduces error rapidly at first (by repeatedly averaging), but improvement saturates at a floor $v^{\star}$ determined by communication distortion and shared correlation. Longer messages (larger $m$) reduce $\sigma_{c}^{2}(m)$, lowering the floor, but also reduce feasible fan-in due to the context constraint $bm\leq W$.
3.4 Scale-out versus scale-up under a fixed budget
The closed forms above become design tools once we connect them to budget. We use the budget model from Section 2.5, where each leaf receives compute $x$ and each edge uses message length $m$. For a full $b$-ary tree of depth $L$ with $N=b^{L}$ leaves, the total cost is approximately
$B\;\approx\;N\big(x+c_{0}(b,m)\big),\qquad c_{0}(b,m)=2m\frac{b}{b-1},$
so the number of leaves is $N\approx B/(x+c_{0})$, and depth is $L\approx\log_{b}N$.
A growth-regime approximation.
When the tree has not yet hit the communication floor, the dominant term in (10) is the shrinking term $a^{L}v_{0}(x)$. Using $N=b^{L}$ and $a^{L}=N^{-t}$ with
$t(b,\rho):=-\frac{\log a}{\log b}=-\frac{\log\big(\rho+(1-\rho)/b\big)}{\log b},$ (14)
we obtain the approximation
$v_{L}\;\approx\;v_{0}(x)\,N^{-t(b,\rho)}.$ (15)
The exponent $t\in[0,1]$ quantifies how efficiently scaling out reduces error in the continuous setting. If $\rho=0$, then $a=1/b$ and $t=1$, recovering the familiar $1/N$ averaging benefit. As $\rho$ increases, $a$ approaches 1 and $t$ approaches 0, reflecting that shared residuals diminish the value of adding more agents.
Compute allocation and a simple threshold.
Substituting $v_{0}(x)\approx cx^{-\beta}$ and $N\approx B/(x+c_{0})$ into (15) yields
$v_{L}\;\approx\;c\,x^{-\beta}\left(\frac{B}{x+c_{0}}\right)^{-t}=c\,B^{-t}\,x^{-\beta}(x+c_{0})^{t}.$
For a fixed budget $B$, minimizing this expression over $x$ reduces to minimizing $x^{-\beta}(x+c_{0})^{t}$. A short calculus argument shows a qualitative threshold:
  • If $t>\beta$, there is a unique interior optimum
    $x^{\star}\;=\;\frac{\beta}{t-\beta}\,c_{0}(b,m).$ (16)
    In this regime, scale-out and scale-up should be balanced: make agents strong enough that their improvement exponent $\beta$ is not wasted, but keep them weak enough that adding more of them leverages the organization exponent $t$.
  • If $t\leq\beta$, the objective decreases as $x$ increases, suggesting that under the growth approximation it is better to spend budget on fewer stronger agents rather than more weaker ones. Correlation and context constraints tend to push systems into this regime by reducing $t$ (via higher $\rho$) or restricting $b$, which increases communication overhead $c_{0}(b,m)$ and raises the floor $v^{\star}(b,m,\rho)$, shortening the growth regime.
This is the continuous analogue of our main binary condition $s>\beta$ developed later: synergy from scaling out appears only when the organization exponent exceeds the single-agent scaling exponent.
How the floor changes the budget story.
The growth approximation (15) cannot hold indefinitely because of the communication floor $v^{\star}$. Once the predicted $v_{L}$ approaches $v^{\star}(b,m,\rho)$, additional budget spent on more leaves or more depth yields diminishing returns. At that point, the effective levers are no longer $N$ and $L$, but $m$ and diversity (which affect $\sigma_{c}^{2}(m)$ and $\rho$). This conclusion will reappear in the binary setting through fixed-point saturation and mixing depth.
3.5 Design implications and a bridge to the binary analysis
The continuous warm-up already exposes the structural trade-offs we will reuse in the binary theory. Absent a binding context constraint, star aggregation is information-dominant because it avoids multi-hop loss; trees are useful primarily as a way to bypass the fan-in limit $Nm\leq W$ by enforcing $bm\leq W$ locally, at the cost of repeated communication. Both correlation and communication loss induce explicit performance floors, so scale-out can stall early when $\rho$ is large or messages are too short. In the growth regime under a fixed budget, the key comparison is exponent versus exponent: whether the organization exponent beats the single-agent scaling exponent.
The binary case is qualitatively different. Averaging is linear, so errors evolve smoothly toward an explicit floor. Majority vote discards magnitude and retains only a sign; when the incoming signal is too weak relative to noise and correlation, repeated majority aggregation can wash out information rather than denoise it. Section 4 formalizes this amplification–collapse transition and shows how the same measurable quantities ($\rho$, message length $m$, fan-in $b$, and budget allocation) determine whether organization helps or hurts on success/failure tasks.
4 Binary Core: Majority Aggregation and Phase Transitions
Binary success/failure tasks behave differently from the continuous warm-up because majority vote discards magnitude information and retains only a sign. That nonlinearity can either amplify weak signal across layers or wash it out entirely. As a result, hierarchical organization exhibits a sharp tipping point: below a threshold, deeper trees collapse to chance; above it, depth improves accuracy until it saturates.
Real multi-agent pipelines often mix graded intermediate signals (scores, confidences, rankings) with hard decisions (accept/reject, pass/fail). We therefore view the continuous and binary analyses as two canonical primitives that appear within richer protocols; the same measurable constraints (communication fidelity, shared failures, and context-limited fan-in) govern both.
4.1 Majority maps and correlated aggregation
From leaf votes to bias.
[…]
$\mu:=\mathbb{E}[\widehat{Y}Y],\qquad\Pr(\widehat{Y}=Y)=\frac{1+\mu}{2}.$
In a tree, each internal node receives $b$ child votes and outputs an aggregated vote. To analyze such systems, we need to understand how bias transforms through one aggregation step.
Majority aggregation map.
Let $b\geq 3$ be an odd integer and let $V_{1},\dots,V_{b}\in\{-1,+1\}$ be i.i.d. child votes with
$\mathbb{E}[V_{i}Y]=u\in[-1,1].$
Equivalently, $V_{i}$ is correct with probability $p=(1+u)/2$. Define the majority vote
$\mathrm{Maj}(V_{1},\dots,V_{b}):=\mathrm{sign}\!\Big(\sum_{i=1}^{b}V_{i}\Big)\in\{-1,+1\},$
where ties are impossible because $b$ is odd. The induced majority map is the bias at the output:
$f_{b}(u)\;:=\;\mathbb{E}\big[\mathrm{Maj}(V_{1},\dots,V_{b})\,Y\big]=2\,\Pr\!\Big(\mathrm{Bin}(b,p)\geq\tfrac{b+1}{2}\Big)-1,\qquad p=\frac{1+u}{2}.$ (17)
The function $f_{b}$ summarizes the effect of one noiseless majority aggregation step on the bias.
Lemma 1 (Basic properties of the majority map)
For odd $b\geq 3$, the map $f_{b}:[-1,1]\to[-1,1]$ satisfies:
  1. 1.
    $f_{b}$ is odd and increasing, with $f_{b}(0)=0$ and $f_{b}(1)=1$.
  2. 2.
    $f_{b}$ is concave on $[0,1]$.
  3. 3.
    The derivative at the origin is
    $f_{b}^{\prime}(0)=\frac{b\binom{b-1}{(b-1)/2}}{2^{b-1}}.$ (18)
Lemma 1 is standard for majority maps; we provide a self-contained proof in Appendix A. The quantity $f_{b}^{\prime}(0)$ is especially important. It is the linear gain of the majority when the inputs are only slightly better than chance. As $b$ increases, $f_{b}^{\prime}(0)$ grows sublinearly (in fact on the order of $\sqrt{b}$), so increasing fan-in yields diminishing marginal amplification. Figure 4 visualizes $f_{b}(u)$ for several fan-ins, highlighting its concavity and the increasing (but sublinear) small-signal gain $f_{b}^{\prime}(0)$.
input bias uuoutput bias $f_{b}(u)$f=uf=ub=3b=3b=5b=5b=7b=7b=11b=11
Slope at origin: $f_{b}^{\prime}(0)=\frac{b\binom{b-1}{(b-1)/2}}{2^{b-1}}$    (grows as $O(\sqrt{b})$)
Figure 4: The majority aggregation map $f_{b}(u)$ for odd fan-in $b\in\{3,5,7,11\}$. Each curve shows how input bias $u$ transforms into output bias after majority voting. All curves pass through $(0,0)$ and $(1,1)$, with slope $f_{b}^{\prime}(0)>1$ at the origin—majority amplifies weak signals. Larger fan-in produces stronger amplification, but the gain grows only as $O(\sqrt{b})$.
A tractable shared-failure model.
To capture groupthink, we need a model in which child votes are positively correlated. We use the following local generative model, which matches a measurable correlation coefficient and yields a closed-form aggregation rule.
Definition 2 ($\rho$-shared correlation model)
Fix a target bias $u\in[-1,1]$ and correlation strength $\rho\in[0,1)$. Conditional on $Y$, the child votes $V_{1},\dots,V_{b}\in\{-1,+1\}$ are generated as follows: with probability $\rho$, all children share a common vote $V_{1}=\cdots=V_{b}=:V$ where $\mathbb{E}[VY]=u$; with probability $1-\rho$, the votes are i.i.d. with $\mathbb{E}[V_{i}Y]=u$.
This model has two practical advantages. First, $\rho$ equals the pairwise correlation of signed correctness. Let $S_{i}:=V_{i}Y\in\{-1,+1\}$. Then $\mathrm{Corr}(S_{i},S_{j})=\rho$ for any $i\neq j$, so $\rho$ can be estimated from logs as described in Section 2.4. Second, the induced aggregation map is explicit. Figure 5 provides an intuition for the mixture: with probability $\rho$ the children act as a single shared voter (groupthink), and with probability $1-\rho$ they behave as independent voters.
Lemma 3 (Correlated majority map under the $\rho$-shared model)
Under Definition 2, the output bias of majority aggregation is
$f_{b,\rho}(u)\;:=\;\mathbb{E}\big[\mathrm{Maj}(V_{1},\dots,V_{b})\,Y\big]=\rho\,u+(1-\rho)\,f_{b}(u).$ (19)
In particular, $f_{b,\rho}$ is increasing and concave on $[0,1]$, and
$f_{b,\rho}^{\prime}(0)=\rho+(1-\rho)f_{b}^{\prime}(0).$ (20)
Lemma 3 is immediate from the mixture construction: in the shared mode, majority returns the same vote (bias $u$); in the independent mode, it returns $f_{b}(u)$. The key design message is already visible in (20): increasing $\rho$ moves $f_{b,\rho}$ closer to the identity map, weakening amplification and making deep organization harder.
[…]
Figure 5: The $\rho$-shared correlation model (Definition 2). With probability $\rho$, all $b$ child agents share a common vote $V$ (left), producing perfect within-group correlation. This captures “groupthink” where agents fail together. With probability $1-\rho$, agents vote independently given the truth $Y$ (right), allowing errors to cancel through aggregation. The parameter $\rho$ equals the pairwise correlation of signed correctness and can be estimated from logs.
4.2 Phase transition in deep trees: $\alpha_{\rho}\gtrless 1$
We now incorporate communication and study deep trees. Consider a homogeneous $b$-ary tree in which every internal node aggregates $b$ child votes by majority, and every edge transmits a message of length $m$ tokens. Communication is modeled as an effective bit channel with reliability $\gamma(m)\in(0,1]$ (Section 2.3): if a child vote has bias $\mu$, then the parent receives a vote with bias $\gamma(m)\mu$.
One-layer bias recursion.
Let $\mu_{t}$ denote the bias of a node at depth $t$ from the leaves (so leaves are $t=0$). After one hop, each child vote’s marginal bias is attenuated to $\gamma(m)\mu_{t}$. We model the $b$ received votes entering each majority gate by the $\rho$-shared model (Definition 2) with this marginal bias and within-group correlation $\rho$. (Any additional dependence induced by shared message templates or decoding is absorbed into the effective $\rho$.) Under this abstraction, the bias evolves as
$\mu_{t+1}=T(\mu_{t}),\qquad T(u):=f_{b,\rho}\!\big(\gamma(m)\,u\big).$ (21)
The map $T$ is increasing, concave on $[0,1]$, and satisfies $T(0)=0$.
The effective gain $\alpha_{\rho}$.
The local behavior of $T$ near the origin is governed by the derivative
$\alpha_{\rho}:=T^{\prime}(0)=\gamma(m)\,f_{b,\rho}^{\prime}(0)=\gamma(m)\big(\rho+(1-\rho)f_{b}^{\prime}(0)\big).$ (22)
We interpret $\alpha_{\rho}$ as the effective per-layer gain on weak signal. If leaf votes are only slightly better than random, then one layer of aggregation and communication multiplies their bias by approximately $\alpha_{\rho}$. This single scalar already suggests a threshold: if $\alpha_{\rho}<1$, weak signal shrinks from layer to layer; if $\alpha_{\rho}>1$, it grows.
The next theorem shows that this intuition is exact at the level of global dynamics: $\alpha_{\rho}$ determines whether deep trees amplify accuracy or collapse to chance.
Theorem 4 (Phase transition for deep majority trees)
Fix odd $b\geq 3$, correlation $\rho\in[0,1)$, and channel reliability $\gamma=\gamma(m)\in(0,1]$. Let $T(u)=f_{b,\rho}(\gamma u)$ and $\alpha_{\rho}=T^{\prime}(0)$ as in (21)–(22). Consider the recursion $\mu_{t+1}=T(\mu_{t})$ with $\mu_{0}\in[0,1]$.
  1. 1.
    (Subcritical collapse.) If $\alpha_{\rho}\leq 1$, then $\mu_{t}\to 0$ as $t\to\infty$.
  2. 2.
    (Supercritical amplification and saturation.) If $\alpha_{\rho}>1$, then there exists a unique fixed point $\mu^{\star}\in(0,1]$ such that $T(\mu^{\star})=\mu^{\star}$, and $\mu_{t}\to\mu^{\star}$ for every $\mu_{0}\in(0,1]$. Moreover, $\mu^{\star}=1$ if $\gamma=1$, and $\mu^{\star}<1$ if $\gamma<1$.
We assume odd fan-in $b$ to avoid ties in majority voting; even $b$ can be handled by specifying a tie-breaking rule (e.g., random) with only minor notational changes. Figure 6 illustrates the two regimes: when $\alpha_{\rho}<1$ the recursion map lies below the diagonal and $\mu_{t}$ collapses to 0, while when $\alpha_{\rho}>1$ the map crosses the diagonal at a stable fixed point $\mu^{\star}$, yielding amplification followed by saturation.
Proof sketch.
Because $T$ is concave with $T(0)=0$, the ratio $r(u):=T(u)/u$ is nonincreasing on $u\in(0,1]$. Its limit at the origin is $r(0^{+})=T^{\prime}(0)=\alpha_{\rho}$. If $\alpha_{\rho}\leq 1$, then $T(u)\leq u$ for all $u\in(0,1]$, so the recursion is monotone decreasing and must converge to a fixed point; the only possibility is 0. If $\alpha_{\rho}>1$, then $r(0^{+})>1$, while $r(1)=T(1)\leq 1$ with strict inequality when $\gamma<1$. By monotonicity of $r$, there is a unique $u$ where $r(u)=1$, yielding a unique $\mu^{\star}$. The sign of T(u)uT(u)-u implies the recursion moves toward $\mu^{\star}$ from any starting point. Full details are given in Appendix C. $\square$
Connection to reconstruction thresholds.
In the classical broadcasting problem on trees with a binary symmetric channel (BSC) of second eigenvalue $\theta$, the Kesten–Stigum condition $b\theta^{2}>1$ characterizes when the root remains reconstructable from the leaves (it is tight for the two-state symmetric model and, more generally, gives the Kesten–Stigum bound for Bayes-optimal reconstruction) (Kesten and Stigum, 1966; Evans et al., 2000; Mossel, 2001). Our condition $\alpha_{\rho}>1$ plays an analogous role for the specific majority-based recursion under our correlated-input model and effective reliability $\gamma(m)$: it is exactly the linear gain $T^{\prime}(0)$ of the update map at the origin. We emphasize that $\alpha_{\rho}$ is estimator-dependent (majority rather than Bayes-optimal) and explicitly incorporates shared failures through $\rho$.
What the threshold means in practice.
Theorem 4 clarifies why adding hierarchical structure is risky. In a supercritical regime, a tree can turn many weak votes into a strong decision and is robust to depth until it saturates. In a subcritical regime, depth destroys information: the system becomes less reliable as it grows. Both correlation ($\rho$) and short messages (small $\gamma(m)$) push $\alpha_{\rho}$ downward, making collapse more likely. This matches the practitioner experience that groupthink and aggressive compression are the two most common reasons multi-agent scaling fails.
$\mu_{t}$$T(\mu_{t})$$T(\mu)=\mu$$\mu^{\star}$$\alpha_{\rho}<1$$\alpha_{\rho}>1$
(a) Recursion map (schematic)
depth $t\mu_{t}$$\mu^{\star}$collapseamplify then saturate
(b) Trajectories (schematic)
Figure 6: A phase transition in deep majority-vote trees. Left: the recursion map $T(\mu)=f_{b,\rho}(\gamma\mu)$ compared to the diagonal. If $\alpha_{\rho}=T^{\prime}(0)<1$, the map stays below the diagonal and bias decays to 0. If $\alpha_{\rho}>1$, the map is initially above the diagonal, creating amplification, but boundedness and communication loss yield a stable fixed point $\mu^{\star}$. Right: corresponding trajectories from the same initialization.
4.3 Small-signal amplification and the organization exponent $s$
The phase transition determines whether deep organization is viable at all. When it is viable ($\alpha_{\rho}>1$), the next question is how fast accuracy can grow with the number of leaves. This is where a simple exponent emerges.
Linear regime and a global upper bound.
Concavity of $T$ implies a strong inequality: the tangent line at the origin upper-bounds the entire map on $[0,1]$,
$T(u)\leq T^{\prime}(0)\,u=\alpha_{\rho}\,u,\qquad u\in[0,1].$ (23)
Thus, for a depth-$L$ tree,
$\mu_{L}\leq\mu_{0}\,\alpha_{\rho}^{L}.$ (24)
This bound is informative even when $\alpha_{\rho}>1$: it tells us that exponential growth in depth is the fastest possible behavior in the small-signal regime.
To complement (24) we also need a lower bound showing that the upper bound is tight when signals remain small. Because $T$ is differentiable at 0, its local behavior is well-approximated by its derivative.
Theorem 5 (Small-signal amplification band)
Fix $b,\rho,\gamma$ and let $T(u)=f_{b,\rho}(\gamma u)$ with $\alpha_{\rho}=T^{\prime}(0)$. For any $\eta\in(0,1)$, there exists $\delta=\delta(\eta;b,\rho,\gamma)\in(0,1]$ such that for all $u\in[0,\delta]$,
$(1-\eta)\,\alpha_{\rho}\,u\;\leq\;T(u)\;\leq\;\alpha_{\rho}\,u.$ (25)
Consequently, if $\mu_{0}\,\alpha_{\rho}^{L}\leq\delta$, then the depth-$L$ recursion satisfies
$(1-\eta)^{L}\,\mu_{0}\,\alpha_{\rho}^{L}\;\leq\;\mu_{L}\;\leq\;\mu_{0}\,\alpha_{\rho}^{L}.$ (26)
Proof sketch.
The upper bound is (23). For the lower bound, differentiability of $T$ at 0 implies $T(u)/u\to\alpha_{\rho}$ as $u\downarrow 0$. Thus there exists $\delta$ such that $T(u)\geq(1-\eta)\alpha_{\rho}u$ on $0,\delta]$. If $\mu_{0}\alpha_{\rho}^{L}\leq\delta$, then the upper bound ensures $\mu_{t}\leq\delta$ for all $t\leq L$, so the lower inequality applies at every step and yields (26). Full details appear in Appendix D. $\square$
The organization exponent.
For a full $b$-ary tree of depth $L$, the number of leaves is $N=b^{L}$. Theorem 5 implies that in the growth regime (before saturation),
$\mu_{L}\approx\mu_{0}\,\alpha_{\rho}^{L}=\mu_{0}\,N^{s},\qquad s:=\frac{\log\alpha_{\rho}}{\log b}.$
We call $s$ the organization exponent. It converts the per-layer gain $\alpha_{\rho}$ into a statement about how performance scales with the number of leaves under hierarchical aggregation.
This exponent is the main interface between organization and budget. Later, in Section 5, we compare $s$ to the single-agent scaling exponent $\beta$ to decide whether scale-out can beat scale-up under a fixed budget. At this stage, the key point is conceptual: $s$ increases with better communication ($\gamma(m)$), decreases with shared failures ($\rho$), and is limited by fan-in through $f_{b}^{\prime}(0)$.
4.4 Saturation and finite-depth guarantees: $L_{\mathrm{mix}}$ and clipped objectives
Small-signal amplification does not continue forever. Bias is bounded by 1, and communication loss and correlation prevent arbitrarily deep trees from achieving perfect accuracy unless the channel is lossless. Theorem 4 already tells us that in the supercritical regime the recursion converges to a fixed point $\mu^{\star}$. In practice, we need two additional pieces of information: how quickly $\mu_{t}$ approaches $\mu^{\star}$, and how to build a design objective that remains accurate across both growth and saturation regimes.
Mixing depth.
Because $T$ is concave and $T(\mu^{\star})=\mu^{\star}$, the slope at the fixed point satisfies $T^{\prime}(\mu^{\star})\leq 1$, with strict inequality whenever $\gamma<1$ or $\rho<1$. This creates a locally contracting region around $\mu^{\star}$, which implies geometric convergence once the recursion enters that region.
Theorem 6 (Finite-depth convergence to the fixed point)
Assume $\alpha_{\rho}>1$ and $\rho\in[0,1)$. Let $\mu^{\star}\in(0,1]$ be the unique fixed point of $T$ from Theorem 4. For any $\varepsilon\in(0,1)$, there exists a finite depth $L_{\mathrm{mix}}(\varepsilon)$ such that
$\mu_{L}\geq(1-\varepsilon)\mu^{\star}\qquad\text{for all }L\geq L_{\mathrm{mix}}(\varepsilon)$
whenever $\mu_{0}>0$. Moreover, $L_{\mathrm{mix}}(\varepsilon)$ can be chosen to scale as
$L_{\mathrm{mix}}(\varepsilon)=O\!\left(\log\frac{1}{\mu_{0}}+\log\frac{1}{\varepsilon}\right),$
with constants depending only on $(b,\rho,\gamma)$.
Proof sketch.
There are two phases. First, while $\mu_{t}$ is small, the map grows approximately linearly with factor $\alpha_{\rho}>1$, so reaching a constant-size neighborhood of $\mu^{\star}$ takes $O(\log(1/\mu_{0}))$ steps. Second, concavity implies $T^{\prime}(\mu)\leq\kappa<1$ on a neighborhood below $\mu^{\star}$, yielding a contraction argument that reaches relative error $\varepsilon$ in $O(\log(1/\varepsilon))$ additional steps. An explicit construction and constants appear in Appendix E. $\square$
A clipped objective for design.
The results above suggest a simple approximation that is both interpretable and safe: use the small-signal prediction $\mu_{0}\alpha_{\rho}^{L}$ when it is below the saturation point, and clip it at $\mu^{\star}$ otherwise:
$\widehat{\mu}_{L}:=\min\big\{\mu^{\star},\ \mu_{0}\,\alpha_{\rho}^{L}\big\}.$ (27)
The clipped objective captures the two dominant regimes with a single expression. In the common small-signal regime $\mu_{0}\leq\mu^{\star}$, concavity yields the global upper bound $\mu_{L}\leq\widehat{\mu}_{L}$, while Theorem 5 guarantees that $\mu_{L}$ tracks the growth term when $\widehat{\mu}_{L}=\mu_{0}\alpha_{\rho}^{L}$. Theorem 6 guarantees that $\mu_{L}$ tracks the saturation term when $\widehat{\mu}_{L}=\mu^{\star}$ and depth exceeds $L_{\mathrm{mix}}$.
Design takeaway.
Equation (27) makes a practical point precise. In the supercritical regime, adding depth and leaves helps only until the system approaches $\mu^{\star}$. Beyond that point, additional budget should be spent on changing $\mu^{\star}$ itself, which requires improving communication fidelity (increasing $\gamma(m)$ by using longer messages) or reducing shared-error correlation $\rho$ (increasing diversity or adding verification). In the subcritical regime, adding depth is actively harmful; the only viable route to scale-out is to move the system across the threshold $\alpha_{\rho}>1$ by improving $m$, reducing $\rho$, or increasing fan-in $b$ within the context constraint.
Section 5 uses the organization exponent $s$, the threshold $\alpha_{\rho}>1$, and the clipped objective (27) to derive topology and budget phase diagrams under context limits. Section 6 turns these results into a theory-guided design algorithm that outputs monotone communication design curves and compute allocations under a fixed budget.
5 Topology and Budget Phase Diagrams
Sections 34 characterized how information transforms through a single aggregation layer. This section turns those layer-wise results into system-level guidance: when centralized aggregation (star) is feasible, when hierarchy (tree) is necessary, and how these choices interact with a fixed budget $B$ under a context window constraint $W$. The resulting phase diagrams organize a large design space into regimes governed by the environment parameters of Section 2 (the single-agent scaling exponent $\beta$, shared-failure correlation $\rho$, and communication fidelity $\gamma(m)$) together with the context constraint $W$.
We focus on the binary setting and use the recursion and organization exponent developed in Section 4; when helpful, we connect back to the continuous warm-up to build intuition.
5.1 Star as an upper bound without constraints
A recurring empirical pattern is that hierarchies sometimes outperform naive centralization. It is tempting to interpret this as “hierarchy is intrinsically better.” Without constraints, that interpretation is false. A star is not merely a particular topology; it is the topology that gives the decision rule access to all leaf information in a single place. Any hierarchy that compresses information on the way up can only discard information, not create it (Cover and Thomas, 2006).
We formalize this point in an intentionally general way, independent of the details of the aggregation rule.
Proposition 7 (Centralization dominates under unlimited context)
Let $X=(X_{1},\dots,X_{N})$ denote the collection of all leaf messages (or leaf outputs) generated in response to a task with truth label $Y$. Consider any multi-agent protocol on any directed acyclic topology such that the root ultimately observes a variable $Z$ that is a (possibly randomized) function of $X$ with randomness independent of $Y$. Then for any loss $\ell(\widehat{Y},Y)$,
$\inf_{\widehat{Y}=\phi(Z)}\mathbb{E}[\ell(\widehat{Y},Y)]\;\geq\;\inf_{\widehat{Y}=\psi(X)}\mathbb{E}[\ell(\widehat{Y},Y)].$
In particular, a centralized decision rule $\psi(X)$ that has access to $X$ cannot perform worse than any rule $\phi(Z)$ that only sees $Z$.
Proof.
Let $U$ collect the protocol’s internal randomness, independent of $Y$, so that $Z=g(X,U)$. Any rule $\phi(Z)$ induces a (possibly randomized) rule $\psi(X,U):=\phi(g(X,U))$; hence predictors based on $Z$ are a subset of predictors based on $(X,U)$ (and thus no better than those based on $X$), proving the risk inequality. $\square$
Proposition 7 is not an endorsement of star in constrained settings. It is a sanity check: hierarchies are only needed because some constraint makes full centralization infeasible or too expensive. In this paper, the dominant such constraint is the context window $W$, which caps the fan-in of any node (Section 2.5). Once $W$ binds, a star can only incorporate a bounded number of leaves regardless of budget, and that is precisely where hierarchy becomes a meaningful design choice.
5.2 Why chains degrade without new evidence
Chains are appealing in practice because they resemble deliberation: one agent’s output becomes the next agent’s input. However, in our framework a chain does not introduce new independent evidence; it only retransmits and transforms what already exists. Under lossy communication, repeated retransmission compounds loss.
This is explicit in both settings. In the continuous warm-up, (8) shows MSE increases linearly with chain length. In the binary setting, the effect is even simpler.
Proposition 8 (Bias decays along a chain under lossy communication)
Consider a chain of length $L$ in which a single leaf vote with bias $\mu_{0}$ is transmitted through $L$ independent communication steps, each with reliability $\gamma(m)$, and no new independent evidence enters the chain. Then the final bias is
$\mu_{\text{chain}}(L)=\gamma(m)^{L}\,\mu_{0}.$
Consequently, if $\gamma(m)<1$ and $L\geq 1$, then $\mu_{\text{chain}}(L)<\mu_{0}$.
Proof.
Each hop multiplies bias by $\gamma(m)$ by the channel model in Section 2.3. $\square$
Design takeaway.
Absent new evidence, chains are not a mechanism for synergy. If a task is inherently sequential, chains may still be necessary, but then one must model the arrival of new information explicitly. If intermediate agents inject new information or perform error-correcting verification, the effective bias can increase; Proposition 8 isolates the relay-only communication effect. Our phase diagrams therefore focus on star versus tree, which represent the two competing strategies for aggregating many leaf signals under context constraints.
5.3 Context-limited star versus hierarchical trees
The context window $W$ turns Proposition 7 from a theoretical upper bound into a practical obstacle. A star aggregator that receives $N$ messages of length $m$ must read roughly $Nm$ tokens, so feasibility requires
$Nm\leq W\qquad\Longrightarrow\qquad N\leq N_{\max}(m):=\Big\lfloor\frac{W}{m}\Big\rfloor.$ (28)
For fixed $W$, $N_{\max}(m)$ is a constant independent of budget. Thus, once a star uses as many leaves as it can fit in context, additional budget cannot increase the number of incorporated leaves. The only remaining levers are to strengthen each leaf (increase $x$) or to increase message length (increase $m$) at the cost of reducing $N_{\max}(m)$.
A hierarchical tree avoids the global cap (28) by enforcing the context constraint locally. If each internal node has fan-in $b$ and reads $b$ messages of length $m$, feasibility requires
$bm\leq W,$ (29)
but the number of leaves can grow as $N=b^{L}$ with depth $L$. In other words, a tree converts a global bottleneck ($Nm\leq W$ at one node) into many local bottlenecks ($bm\leq W$ at many nodes). Figure 7 illustrates this trade-off: the star saturates early due to context limits, while the tree can continue scaling until it reaches its fixed-point ceiling $\mu^{\star}$.
[…]
Figure 7: Schematic performance comparison under a fixed context window $W$. Star (orange) aggregates NN leaves but saturates when NN hits the ceiling $N_{\max}=\lfloor W/m\rfloor$. Tree (blue) distributes the constraint locally ($bm\leq W$), enabling $N=b^{L}$ leaves. When $s>\beta$, trees overtake at budget $B_{\mathrm{crit}}$ and continue with slope $s$ until reaching $\mu^{\star}$.
This benefit is real only if information survives the hierarchy. The binary phase transition from Section 4.2 provides the feasibility condition: a deep tree can only maintain nontrivial accuracy when
$\alpha_{\rho}=\gamma(m)\big(\rho+(1-\rho)f_{b}^{\prime}(0)\big)>1.$ (30)
Even when $\alpha_{\rho}>1$, deeper trees saturate at a fixed point $\mu^{\star}$, so trees should be viewed as a tool for scaling up to a regime of strong performance, not for unlimited improvement.
[…]
Figure 8: A schematic binary phase diagram for hierarchical trees at fixed fan-in $b$ and single-agent scaling exponent $\beta$. The lower boundary $\alpha_{\rho}=1$ (Theorem 4) separates regimes where deep hierarchies collapse to chance from regimes where they amplify accuracy. Within the amplifying regime, the stronger condition $\alpha_{\rho}>b^{\beta}$ (equivalently $s>\beta$) marks where scaling out can beat scaling up in the budgeted growth regime. Both boundaries move upward as $\rho$ increases. Since $\gamma(m)\leq 1$, the $\alpha_{\rho}=b^{\beta}$ boundary may exit the feasible range at large $\rho$; the schematic clips it at $\gamma=1$, truncating the synergy region.
Figure 8 summarizes the logic in a single picture. A tree can only be useful if it lies above the amplification threshold $\alpha_{\rho}>1$. Among such trees, budgeted synergy further requires $s>\beta$, which is equivalent to
$s>\beta\qquad\Longleftrightarrow\qquad\alpha_{\rho}>b^{\beta}.$ (31)
The next subsection turns (31) into an explicit budget threshold.
5.4 Budget thresholds for synergy: conditions on $s$ and $\beta$
We now connect topology to budget. The question is not whether a deep tree can amplify weak votes in isolation (Section 4), but whether it can do so more efficiently than simply making a single agent stronger. This is where the comparison between the organization exponent $s$ and the single-agent scaling exponent $\beta$ becomes decisive.
Setup (growth regime).
We work in the small-signal regime where the single-agent scaling law is well approximated by
$\mu_{0}(x)=g(x)\approx kx^{\beta},$
and where the tree has not yet saturated (so Theorem 5 applies). For a full $b$-ary tree of depth $L$ with $N=b^{L}$ leaves, the small-signal prediction gives
$\mu_{L}\approx\mu_{0}(x)\,\alpha_{\rho}^{L}=\mu_{0}(x)\,N^{s},\qquad s=\frac{\log\alpha_{\rho}}{\log b}.$ (32)
Under the budget model of Section 2.5, the total cost is approximately
$B\approx N\big(x+c_{0}(b,m)\big),\qquad c_{0}(b,m)=2m\frac{b}{b-1},$
so $N\approx B/(x+c_{0})$. Substituting into (32) yields the design objective
$\mu_{\mathrm{grow}}(B;b,m,x)\;\approx\;k\,x^{\beta}\left(\frac{B}{x+c_{0}(b,m)}\right)^{s}.$ (33)
This objective is interpretable: $x^{\beta}$ captures scale-up, and $(B/(x+c_{0}))^{s}$ captures scale-out through the number of leaves.
Theorem 9 (Closed-form compute allocation in the growth regime)
Fix $(b,m,\rho)$ and suppose $s(b,m,\rho)>\beta$. Then the growth objective (33) is maximized at
$x^{\star}\;=\;\frac{\beta}{s-\beta}\,c_{0}(b,m),$ (34)
and the resulting optimized growth prediction scales as
$\mu_{\mathrm{grow}}^{\star}(B;b,m)\;\approx\;k\,\kappa(s,\beta)\,c_{0}(b,m)^{\beta-s}\,B^{s},\qquad\kappa(s,\beta):=\beta^{\beta}(s-\beta)^{\,s-\beta}s^{-s}.$ (35)
If $s\leq\beta$, the objective (33) is maximized by pushing $x$ as large as allowed (scale up rather than scale out).
Proof sketch.
For $s>\beta$, maximize $\log\mu_{\mathrm{grow}}=\beta\log x+s\log B-s\log(x+c_{0})$ over $x>0$. Setting the derivative to zero gives (34). Substituting back yields (35). If $s\leq\beta$, the derivative is nonnegative for all $x$, so the objective increases with $x$. Details are routine and given in Appendix D. $\square$
Synergy against a single agent.
A single-agent baseline that spends the full budget on compute achieves (in the same small-signal regime)
$\mu_{\mathrm{single}}(B)\approx k\,B^{\beta}.$ (36)
Comparing (35) and (36) yields a simple conclusion: in the growth regime, a hierarchy can only outpace the single agent when $s>\beta$. Moreover, when $s>\beta$, there is an explicit budget threshold beyond which the optimized tree dominates the single-agent scaling curve.
Corollary 10 (A budget threshold for scale-out synergy (growth regime))
Fix $(b,m,\rho)$ with $s>\beta$, and define $\kappa(s,\beta)$ as in (35). Ignoring saturation, a sufficient condition for the optimized tree prediction $\mu_{\mathrm{grow}}^{\star}(B;b,m)$ to exceed the single-agent prediction $\mu_{\mathrm{single}}(B)$ is
$B\;\geq\;B_{\mathrm{crit}}(b,m)\;:=\;c_{0}(b,m)\,\kappa(s,\beta)^{-\frac{1}{s-\beta}}.$ (37)
Interpretation.
Equation (37) is a concrete, testable statement. It predicts that scale-out synergy is harder when coordination is expensive (large $c_{0}$), when the organization exponent barely beats the scaling exponent (small $s-\beta$), or when the system is close to the amplification boundary (small $s$). In Section 4.4 we introduced clipped objectives to account for saturation; when saturation matters, the threshold still indicates whether there exists a budget range in which scale-out should help before the fixed point dominates.
[…]
Figure 9: A schematic budget phase diagram under a fixed context window $W$. A star cannot incorporate more than $N_{\max}(m)$ leaves, so beyond a point its gains come primarily from making leaves stronger (scale-up), leading to a $B^{\beta}$-type growth until task-level saturation. A supercritical tree with organization exponent $s>\beta$ can exhibit a steeper $B^{s}$ growth regime before saturating at $\mu^{\star}$. The budget threshold $B_{\mathrm{crit}}$ marks where optimized scale-out can overtake single-agent scaling in the growth regime.
5.5 A universal upper bound on amplification exponents
The organization exponent $s$ governs how quickly a majority-vote hierarchy can amplify weak signal as the number of leaves grows. It is therefore important to understand its fundamental limits. Perhaps surprisingly, the majority family admits a clean and universal upper bound: no choice of fan-in can make $s$ exceed $1/2$ in the binary model with one-bit messages.
Theorem 11 (A universal bound: $s\leq\tfrac{1}{2}$)
For every odd $b\geq 3$, the majority map satisfies
$f_{b}^{\prime}(0)\leq\sqrt{b}.$ (38)
Consequently, for any $\rho\in[0,1)$ and any $\gamma(m)\leq 1$,
$\alpha_{\rho}=\gamma(m)\big(\rho+(1-\rho)f_{b}^{\prime}(0)\big)\leq\sqrt{b},\qquad\text{and hence}\qquad s=\frac{\log\alpha_{\rho}}{\log b}\leq\frac{1}{2}.$ (39)
Proof sketch.
Equation (38) follows from the closed form $f_{b}^{\prime}(0)=\frac{b\binom{b-1}{(b-1)/2}}{2^{b-1}}$ (Lemma 1) and a standard upper bound on the central binomial coefficient. Plugging (38) into (22) yields (39). A fully elementary proof is given in Appendix H. $\square$
Implications for synergy.
The bound $s\leq 1/2$ has two immediate consequences. First, it explains why scale-out is fragile on some tasks. If the single-agent scaling exponent satisfies $\beta\geq 1/2$, then the inequality $s>\beta$ is impossible under majority-vote hierarchies with one-bit messages. In that regime, a designer should not expect budget-exponent gains from adding agents; improvement must come from making agents stronger, from richer communication than a single bit, or from protocols that effectively reduce shared failures. Second, the bound clarifies the role of diversity. Even if $\beta<1/2$, a large $\rho$ reduces $\alpha_{\rho}$ and pushes $s$ downward, shrinking or eliminating the synergy region in Figure 8. This is precisely the regime in which practitioners observe that “more agents does not help,” and it suggests a concrete remedy: invest budget not in more of the same agents, but in mechanisms that reduce shared error modes, which in our framework means reducing the effective $\rho$.
Summary of the phase-diagram logic.
Taken together, the results above reduce topology choice to a few quantitative checks. Under context constraints, deep trees are only viable in the supercritical regime $\alpha_{\rho}>1$; exponent-level budgeted synergy in the growth range requires $s>\beta$ (equivalently $\alpha_{\rho}>b^{\beta}$); and within the majority-vote tree family we always have the universal limit $s\leq 1/2$. Even when these conditions hold, depth helps only up to saturation: once the recursion is near $\mu^{\star}$, additional budget is typically better spent improving communication fidelity (larger $\gamma(m)$) or reducing shared failures (smaller $\rho$) than on additional depth. Section 6 turns these checks into a compact design diagnostic within the model.
6 Design Diagnostics Within the Model
The results so far characterize when different organizations can help under a fixed budget: a context-limited star saturates (Section 5.3), chains compound loss (Section 5.2), and trees are viable only in the supercritical regime $\alpha_{\rho}>1$ with small-signal growth governed by an organization exponent $s$ (Section 4). Here we translate these results into design diagnostics within the model: given a budget $B$, a context window $W$, and a small set of effective environment quantities, we ask whether hierarchy is feasible, whether scale-out can outpace scale-up before saturation, and which constraint is binding when it cannot. For LLM-based agent systems, the knobs correspond to token allocations: per-leaf reasoning/compute $x$ and inter-agent message length $m$. The abstraction is deliberately coarse, so the goal is regime prediction and diagnostic guidance, not a faithful end-to-end optimizer for any particular agent stack.
Under a budget $B$ and context window $W$, we parameterize a tree design by
[…]
and ask how to choose these quantities to maximize the model-predicted performance subject to feasibility. The inner problem chooses $x$ for fixed $(b,m)$; the outer problem trades off communication quality and coordination cost by selecting $(b,m)$ as a function of $B$. A key structural takeaway is monotonicity: in the growth regime, the budget-optimal message length can be chosen nondecreasing in $B$, yielding efficient design curves rather than expensive search. We focus on binary success/failure tasks, where hierarchies can either amplify or collapse; the continuous setting admits analogous (and smoother) diagnostics via the closed-form recursions of Section 3.
6.1 Closed-form compute allocation: optimal $x^{\star}$ in the growth regime
Fix a candidate tree architecture specified by a fan-in $b$ and message length $m$, with local feasibility constraint $bm\leq W$. Let $\rho$ be the calibrated shared-failure correlation and $\gamma(m)$ be the calibrated communication reliability. Define the per-layer weak-signal gain
$\alpha_{\rho}(b,m):=\gamma(m)\big(\rho+(1-\rho)f_{b}^{\prime}(0)\big),\qquad s(b,m):=\frac{\log\alpha_{\rho}(b,m)}{\log b}\quad\text{(for fixed calibrated }\rho\text{)}.$
Theorem 4 implies that deep hierarchies are only viable if $\alpha_{\rho}(b,m)>1$. Assume we are in this supercritical regime and consider budgets for which the system remains in the growth regime before saturation. Under the small-signal scaling approximation $g(x)\approx kx^{\beta}$ and the tree budget model $B\approx N(x+c_{0}(b,m))$ with $c_{0}(b,m)=2m\frac{b}{b-1}$, Section 5.4 derived the growth objective
$\mu_{\mathrm{grow}}(B;b,m,x)\;\approx\;k\,x^{\beta}\left(\frac{B}{x+c_{0}(b,m)}\right)^{s(b,m)}.$ (40)
The next fact is the inner workhorse of the design procedure: once $(b,m)$ are fixed, the optimal compute per leaf has a closed form.
Proposition 12 (Growth-regime compute allocation)
Fix $(b,m)$ with $s(b,m)>\beta$. Then the objective (40) is maximized at
$x^{\star}(b,m)=\frac{\beta}{s(b,m)-\beta}\,c_{0}(b,m),$ (41)
and the resulting optimized growth prediction scales as
$\mu_{\mathrm{grow}}^{\star}(B;b,m)\;\approx\;k\,\kappa\big(s(b,m),\beta\big)\,c_{0}(b,m)^{\beta-s(b,m)}\,B^{s(b,m)},$ (42)
where $\kappa(s,\beta)=\beta^{\beta}(s-\beta)^{s-\beta}s^{-s}$. If $s(b,m)\leq\beta$, then (40) is (weakly) increasing in $x$, so the growth approximation favors scale-up over scale-out.
Proposition 12 is a restatement of Theorem 9 in design form. It makes the trade-off explicit: when coordination is expensive (large $c_{0}$), each leaf should be stronger (larger $x^{\star}$); when organization is efficient (large $s-\beta$), each leaf can be weaker, freeing budget to add more leaves.
In practice, given $B$, the recommended number of leaves and depth are
$N^{\star}(B;b,m)\;\approx\;\left\lfloor\frac{B}{x^{\star}(b,m)+c_{0}(b,m)}\right\rfloor,\qquad L^{\star}(B;b,m)\;\approx\;\left\lfloor\log_{b}N^{\star}(B;b,m)\right\rfloor.$ (43)
Depth $L^{\star}$ should not exceed the mixing depth needed to approach saturation (Theorem 6); see Section 6.4.
6.2 Monotone communication design: why $m^{\star}(B)$ increases with budget
Compute allocation solves only the inner problem. We still need to choose message length $m$, which trades off two effects: increasing $m$ improves communication fidelity $\gamma(m)$ and thus increases the gain $\alpha_{\rho}$ (raising the exponent $s$), but increasing $m$ also increases the coordination cost $c_{0}(b,m)$, reducing the number of leaves we can afford.
A natural design question is whether there is structure in the budget-dependent solution. The answer is yes: in the growth regime, the optimal message length is monotone in the budget. This formalizes a useful heuristic in the model: as you spend more, the optimal design communicates less aggressively compressed information.
A growth-regime objective indexed by $m$.
Fix fan-in $b$ and consider the feasible message set
$\mathcal{M}_{b}:=\{m\in\mathbb{Z}_{+}:\ bm\leq W\}.$
For each $m\in\mathcal{M}_{b}$, define the growth-optimal prediction from Proposition 12,
$\underline{\mu}_{b}(m;B):=\mu_{\mathrm{grow}}^{\star}(B;b,m)=k\,\kappa(s_{b,m},\beta)\,c_{0}(b,m)^{\beta-s_{b,m}}\,B^{s_{b,m}},\qquad s_{b,m}:=s(b,m).$ (44)
We only consider candidates with $s_{b,m}>\beta$ and $\alpha_{\rho}(b,m)>1$ since only these can outpace scale-up in exponent terms in the growth regime.
Theorem 13 (Monotone message-length design in the growth regime)
Fix fan-in $b\geq 3$ and correlation $\rho\in[0,1)$. Assume $\gamma(m)$ is strictly increasing in $m$ and $c_{0}(b,m)$ is nondecreasing in $m$. Let
$m_{b}^{\star}(B)\in\arg\max_{m\in\mathcal{M}_{b}}\ \underline{\mu}_{b}(m;B),$
where the maximization ranges over candidates satisfying $s_{b,m}>\beta$ and $\alpha_{\rho}(b,m)>1$. Then for any $B_{2}>B_{1}$, there exists an optimal selection such that
$m_{b}^{\star}(B_{2})\ \geq\ m_{b}^{\star}(B_{1}).$
Equivalently, $m_{b}^{\star}(B)$ can be chosen as a nondecreasing function of $B$.
Why this holds.
Taking logs of (44) yields
$\log\underline{\mu}_{b}(m;B)=s_{b,m}\,\log B+q_{b}(m),$
where $q_{b}(m)=\log k+\log\kappa(s_{b,m},\beta)+(\beta-s_{b,m})\log c_{0}(b,m)$ does not depend on $B$. Thus, for fixed $b$, each message length $m$ defines a line in $\log B$ with slope $s_{b,m}$. Because $\gamma(m)$ increases with $m$, the slopes $s_{b,m}$ are increasing in $m$. When choosing among lines with increasing slopes, the maximizing line index can only move in one direction as $\log B$ increases. A full proof, stated in terms of increasing differences and upper envelopes, is given in Appendix I.
$t=\log B$$\log\underline{\mu}_{b}(m;B)$upper envelopemm smallmm mediummm large
(a) Envelope over $\log B$
$t=\log B$$m_{b}^{\star}(B)$$m_{1}$$m_{2}$$m_{3}$
(b) Monotone design curve
Figure 10: Why the optimal message length increases with budget in the growth regime (fixed fan-in $b$). After optimizing compute allocation $x$, each candidate message length $m$ induces a line $\log\underline{\mu}_{b}(m;B)=s_{b,m}\log B+q_{b}(m)$ with slope $s_{b,m}$ increasing in $m$. The maximum over $m$ is the upper envelope of these lines (left), implying that an optimal $m_{b}^{\star}(B)$ can be chosen as a nondecreasing, piecewise-constant function of $B$ (right).
What the theorem does and does not claim.
Theorem 13 is a statement about the growth-regime objective $\underline{\mu}_{b}(m;B)$ for a fixed fan-in $b$. It does not assert global monotonicity when the optimal fan-in $b$ itself changes with budget, nor does it rely on the saturation fixed point $\mu^{\star}$. In practice, once a design approaches saturation, increasing $m$ often remains beneficial because it increases $\alpha_{\rho}$ and improves $\mu^{\star}$, but the theorem is intentionally limited to a regime where the approximation is provably accurate.
6.3 A linear-time envelope algorithm for design curves
The envelope view in Figure 10 yields an efficient algorithm for generating design curves. For a fixed fan-in $b$, we want to compute $m_{b}^{\star}(B)$ for a range of budgets, without brute-force search over all $m$ at every $B$.
Lines and intersections.
Let $t=\log B$. For each feasible $m\in\mathcal{M}_{b}$ with $s_{b,m}>\beta$ and $\alpha_{\rho}(b,m)>1$, define
$\ell_{b,m}(t):=s_{b,m}\,t+q_{b}(m),$ (45)
where $q_{b}(m)$ is the intercept from (44). The maximizer of $\ell_{b,m}(t)$ over $m$ is the envelope. Because $s_{b,m}$ is increasing in $m$, the envelope can be computed by a standard monotone upper-hull procedure: maintain a stack of candidate lines and the $t$-values where one line overtakes the previous one.
Given two lines $\ell_{1}(t)=s_{1}t+q_{1}$ and $\ell_{2}(t)=s_{2}t+q_{2}$ with $s_{1}<s_{2}$, their intersection occurs at
$t^{\star}=\frac{q_{1}-q_{2}}{s_{2}-s_{1}}.$ (46)
If a newly added line overtakes the previous line earlier than the previous line overtook its predecessor, then the previous line is never optimal and can be removed. This is the same geometric logic used in convex-hull tricks for dynamic programming.
Algorithm 1 Envelope construction for $m_{b}^{\star}(B)$ (fixed $b$).
1:Enumerate candidates (in increasing $m$).
[…]
3:Keep $m$ only if $\alpha_{\rho}(b,m)>1$ and $s_{b,m}>\beta$. For each kept $m$, form $\ell_{b,m}(t)=s_{b,m}t+q_{b}(m)$.
4:(Optional robustness) Remove equal-slope duplicates.
5:If two retained candidates have the same slope $s$, keep only the one with larger intercept $q_{b}(m)$.
6:Build the upper envelope (stack).
7:Maintain a stack of $(m,\ell,\tau)$, where $\tau$ is the activation time.
8:for retained $m$ in increasing order do
9:  if stack empty then
10:    push $(m,\ell_{b,m},-\infty)$
11:  else
12:    let $(m_{\text{top}},\ell_{\text{top}},\tau_{\text{top}})$ be stack top
13:    compute intersection time $t^{\star}$ using $t^{\star}=\frac{q_{\text{top}}-q_{b}(m)}{s_{b,m}-s_{\text{top}}}$
14:    while stack nonempty and $t^{\star}\leq\tau_{\text{top}}$ do
15:     pop stack; update $(\ell_{\text{top}},\tau_{\text{top}})$; recompute $t^{\star}$
16:    end while
17:    push $(m,\ell_{b,m},t^{\star})$
18:  end if
19:end for
20:Query (one pass over budgets).
21:For budgets $B_{1}<\dots<B_{T}$, set $t_{i}=\log B_{i}$.
22:Walk a pointer through breakpoints and output the active $m$ for each $t_{i}$.
Complexity and usage.
For a fixed fan-in $b$, the feasible message set has size $|\mathcal{M}_{b}|=\lfloor W/b\rfloor$. Algorithm 1 scans candidates in increasing $m$ (equivalently increasing slope $s_{b,m}$) and builds the upper envelope with a stack in which each candidate line is pushed once and popped at most once. Thus, envelope construction takes $O(|\mathcal{M}_{b}|)$ time and $O(|\mathcal{M}_{b}|)$ memory. Given budgets $B_{1}<\cdots<B_{T}$, the optimal choices $m_{b}^{\star}(B_{t})$ can then be returned in $O(T)$ additional time by a single pass over the envelope breakpoints. To obtain a global design curve, we run Algorithm 1 for each odd $b\geq 3$ with $\mathcal{M}_{b}\neq\emptyset$ (equivalently $b\leq W$ since $m\geq 1$), and then select the best $(b,m)$ per budget using the clipped objective described next.
6.4 Putting the diagnostics together
We now combine compute allocation, monotone communication design, and saturation into a compact workflow. The inputs are a budget $B$, a context window $W$, and a small set of estimated effective quantities, in particular $\mu_{0}(x)$, $\rho$, and $\gamma(m)$ (Section 2.7). The output is a model-based prediction and a set of diagnostics (which constraint is binding), rather than a claim that the abstraction optimizes real agent stacks end-to-end.
Compare against centralization.
For message length $m$, a star can include at most $N_{\max}(m)=\lfloor W/m\rfloor$ leaves. When the budget regime of interest does not push beyond this cap, centralization is the natural baseline: it aggregates in one hop and avoids multi-hop loss (Proposition 7). When $N_{\max}(m)$ binds, additional budget cannot increase the number of incorporated leaves in a star, and hierarchy becomes the only way to expose more parallel evidence under the same per-node context limit.
Screen tree candidates and allocate compute.
For each odd fan-in $b\geq 3$ and each feasible $m$ with $bm\leq W$, compute $\alpha_{\rho}(b,m)$ and $s(b,m)$. Discard candidates with $\alpha_{\rho}(b,m)\leq 1$, which collapse in depth (Theorem 4). For exponent-level budgeted synergy in the growth regime, also discard $s(b,m)\leq\beta$. For remaining candidates, set the per-leaf compute to the closed form $x^{\star}(b,m)$ in (41) (Proposition 12) and choose the implied leaf count and depth via (43), with depth interpreted through the mixing-depth guarantee of Theorem 6. For fixed $b$, Algorithm 1 yields the monotone budget-indexed communication curve $m_{b}^{\star}(B)$, so the outer search can be performed efficiently.
Clip by saturation.
In the supercritical regime, the binary recursion converges to a fixed point $\mu^{\star}(b,m)$ (computed by iterating $u\mapsto f_{b,\rho}(\gamma(m)u)$). To avoid extrapolating the growth law beyond its range, we evaluate a design using the clipped prediction from Section 4.4:
$\widehat{\mu}(B;b,m):=\min\big\{\mu^{\star}(b,m),\ \mu_{0}(x)\,\alpha_{\rho}(b,m)^{L}\big\}.$ (47)
At each budget $B$, we compare this clipped tree prediction to the best feasible star and to the single-agent baseline under the same budget.
Interpretation as diagnostics.
These checks separate four common reasons why scaling out fails: context saturation of a star ($N_{\max}(m)$ caps centralization), subcriticality ($\alpha_{\rho}\leq 1$), exponent mismatch ($s\leq\beta$), and saturation ($\widehat{\mu}$ close to $\mu^{\star}$). When performance plateaus, the theory points to two levers: improve effective communication fidelity (larger $\gamma(m)$ at the same cost) or reduce shared failures (smaller $\rho$ via diversity or verification), rather than increasing depth.
Continuous tasks.
The same workflow applies in the continuous warm-up with $\gamma(m)$ replaced by $\sigma_{c}^{2}(m)$ and $\mu^{\star}$ replaced by the explicit error floor $v^{\star}$ in (11).
7 Empirical Touchpoints
This paper is primarily theoretical: our goal is to explain why budgeted multi-agent scaling amplifies in some regimes and collapses in others, and to make the relevant regime boundaries explicit. Accordingly, we keep the empirical component modest. We (i) connect our bottlenecks and regime predictions to recent large-scale matched-budget evaluations reported elsewhere, and (ii) provide a minimal synthetic sanity check that recovers the predicted amplification boundary under the assumed model.
7.1 Theoretical Mechanisms in Recent Large-Scale Studies
Large-scale matched-budget evaluations are computationally prohibitive for academic groups, but recent industrial studies provide high-fidelity data that validates our theoretical predictions. Most notably, Kim et al. (2025) conducted an extensive evaluation of scaling laws for agent systems across multiple benchmarks and architectures. Their empirical findings align closely with the qualitative regime boundaries derived in our framework:
  • Context saturation and coordination overhead. Kim et al. (2025) observe that simply concatenating agent traces can degrade performance on tool-heavy tasks when coordination overhead consumes the context window. In our framework, this corresponds to the $Nm\leq W$ bottleneck: once a centralized aggregator becomes context-saturated, adding agents can dilute usable signal.
  • Saturation at strong baselines. They report diminishing (and sometimes negative) returns once the single-agent baseline is already strong. This is consistent with our fixed-point analysis: once a hierarchy approaches its saturation point $\mu^{\star}$, marginal gains from adding depth or leaves are small; improvement requires changing effective parameters (better $\gamma(m)$ or smaller $\rho$), not more aggregation.
  • Shared failures and error cascades. They emphasize that agents can share failure modes and that naive aggregation can amplify errors, while more structured coordination can mitigate this. In our framework, shared failures are summarized by $\rho$: higher $\rho$ pushes $\alpha_{\rho}$ downward and can drive hierarchies into the subcritical regime where depth washes out information.
By providing the mathematical derivation for these empirically observed regimes, our theory explains why scaling fails when it does, suggesting that future improvements must come from altering the effective parameters, specifically increasing $\gamma(m)$ or decreasing $\rho$, rather than simply adding more agents.
7.2 Synthetic sanity check: recovering the amplification boundary
Synthetic experiments provide a controlled setting where the ground-truth parameters are known and the phase boundary can be probed directly. Our goal here is not to optimize absolute performance, but to verify that the qualitative prediction—amplification versus collapse governed by $\alpha_{\rho}$—is recovered under the generative model analyzed in Section 4.
We include one lightweight Monte Carlo study under the $\rho$-shared binary model. We fix $b=5$, $\gamma=0.6$, $\mu_{0}=0.05$, sweep $\rho\in[0,0.8]$, and compare the recursion prediction $\mu_{L}$ at depths $L\in\{10,30\}$ to Monte Carlo estimates (n=50,000n=50{,}000). Figure 11 shows close agreement and a sharpening transition around the predicted boundary $\alpha_{\rho}=1$.
Figure 11: Synthetic Monte Carlo sanity check for the binary majority tree under the $\rho$-shared model and a binary symmetric channel with reliability $\gamma$. We plot the root bias $\mu_{L}$ as a function of $\rho$ for $b=5$, $\gamma=0.6$, and $\mu_{0}=0.05$. Curves show recursion predictions at finite depth; markers show Monte Carlo estimates (n=50,000n=50{,}000). The vertical line indicates the threshold $\alpha_{\rho}=1$, separating amplification ($\alpha_{\rho}>1$) from collapse ($\alpha_{\rho}\leq 1$) in the deep limit.
For continuous tasks, the same role can be played by a simple correlated-Gaussian generator, where the closed-form expressions of Section 3.2 make correlation and communication floors explicit.
8 Discussion and Limitations
We asked when scaling out a budgeted multi-agent system yields synergy and when it fails. The framework is intentionally minimal: it treats an agent as a black-box solver with an allocated compute budget, models inter-agent communication as a controllable lossy channel, and summarizes shared failures through an effective correlation parameter. This minimalism makes the regime boundaries transparent, but it also defines what the theory does and does not claim.
8.1 Relationship to prior results
Our binary phase transition echoes reconstruction thresholds for broadcasting on trees (Kesten and Stigum, 1966; Evans et al., 2000; Mossel, 2001), where information can or cannot propagate from leaves to the root depending on a single scalar. Two differences are central for agent-system design. First, our “channel” is partly under design control: message length trades off fidelity $\gamma(m)$ against budget and context constraints, so the relevant threshold depends on an explicit communication choice. Second, we couple propagation to a fixed-budget allocation problem: an amplifiable hierarchy is synergistic only when its organization exponent outpaces single-agent scaling.
Recent work offers complementary perspectives on information flow in agent systems. He et al. (2025) cast agentic architectures as compressor–predictor pairs and propose mutual information between context and compression as a task-independent proxy for information survival; their finding that scaling the compressor can dominate scaling the predictor aligns with our emphasis on communication fidelity $\gamma(m)$ as a first-order bottleneck, though our framework adds explicit budget constraints and multi-hop aggregation structure. Li (2026) study skill selection in single-agent systems and observe a phase transition: selection accuracy remains stable up to a critical library size then drops sharply, with semantic confusability among skills playing a central role. This echoes our amplification–collapse dichotomy, but at a different interface: where we model information loss across agents, they model selection confusion within a single agent’s skill routing. Together, these results suggest that sharp thresholds, whether in inter-agent communication, intra-agent skill selection, or hierarchical aggregation, may be a recurring motif in bounded-resource AI systems.
8.2 What the framework explains (and what it does not)
The framework is designed to explain qualitative regime changes that are otherwise hard to diagnose: why stars saturate under context limits, why chains compound loss, why trees can either amplify or collapse depending on $\alpha_{\rho}$, and why correlated failures impose floors that make large ensembles behave like “one agent with shared blind spots.” It also clarifies what additional sophistication must accomplish: any useful complexity must effectively increase communication fidelity, reduce shared failures, or relax the fan-in constraint.
At the same time, the model is not a faithful simulator of real agent stacks. Parameters such as $\rho$ and $\gamma(m)$ are effective summaries that can vary with depth, prompts, tools, and task difficulty; real communication is far richer than a one-bit abstraction; and practical budget accounting can differ across implementations. For these reasons, we view the theory as a coordination diagnostic—predicting which constraint is binding and which direction should help—rather than as a source of precise point predictions. Crucially, the alignment between our derived bottlenecks and independent large-scale observations (Kim et al., 2025) suggests that the effective parameters defined here capture the structural essence of the scaling problem, rendering the theory robust to changes in the underlying base models.
8.3 How stronger protocols fit the framework
Protocols such as debate, critique, verification, or self-consistency can be interpreted as mechanisms that change the effective parameters of the model (Irving et al., 2018; Christiano et al., 2018; Du et al., 2024; Wang et al., 2023). For example, structured verification can reduce effective shared failures ($\rho$), while richer message schemas can improve effective fidelity ($\gamma(m)$); both, however, typically increase coordination cost. The theory suggests a concrete way to evaluate such protocols: measure how much they move the effective bottlenecks per unit budget, and check whether they move the system across the supercritical boundary $\alpha_{\rho}>1$ or the synergy boundary $s>\beta$.
8.4 Extensions beyond our scope
Several extensions are natural. Allowing heterogeneity across leaves (different models, prompts, or tools) makes $\rho$ and scaling task- and agent-dependent; a tractable next step is to treat $\rho$ as depth-dependent or to model mixtures of agent types. More general graphs and multi-round interaction can be analyzed by unrolling protocols into computation graphs, but doing so requires accounting for how added edges change effective dependence as well as cost. Finally, our strongest binary bounds rely on one-bit-style messages and majority aggregation; allowing richer messages or confidence reports changes the amplification map and may relax universal exponent limits. In all cases, the guiding principle remains the same: additional structure helps only insofar as it relaxes the binding constraints (context, communication loss, or shared failures) under a fixed budget.
9 Conclusion
Multi-agent systems can exhibit synergy, but under a fixed budget scaling out is not monotone: it can help, saturate, or collapse. We argued that these outcomes become predictable once we make the binding constraints explicit: finite context windows, lossy communication, and shared failures, and summarize an environment by a small set of measurable effective parameters.
In a continuous warm-up we derived closed-form risks for star/chain/tree organizations and exposed correlation and communication floors. For binary success/failure tasks, we proved a sharp amplification–collapse phase transition for deep majority trees and characterized the resulting fixed-point saturation and mixing depth. In the amplifying regime we introduced an organization exponent $s$ governing small-signal growth and showed that budgeted synergy in the growth range occurs exactly when $s>\beta$, yielding closed-form compute allocation and explicit budget thresholds. We also translated these results into simple design diagnostics within the model.
The broader message is that topology is not magic. Hierarchies are useful because context limits can make full centralization infeasible, and they succeed only when communication is sufficiently faithful and shared failures are sufficiently weak. We hope this work helps move agent-system design from ad hoc heuristics toward principled, testable coordination theories.
Acknowledgments and Disclosure of Funding
The authors gratefully acknowledge funding from Canada CIFAR AI Chair Program and the Canada NSERC Discovery Grant (RGPIN-2021-03115).
Proof Roadmap and Dependency Graph
This appendix provides self-contained proofs for the main theoretical statements in the paper. For convenience, we list where each result is proved:
  • Lemma 1 (properties of the majority map): Appendix A.
  • Lemma 3 (closed form under the $\rho$-shared model): Appendix B.
  • Theorem 4 (phase transition for deep majority trees): Appendix C.
  • Theorem 5 (small-signal amplification band), Theorem 9 (growth-regime compute allocation), and Corollary 10 (budget threshold): Appendix D.
  • Theorem 6 (finite-depth convergence / mixing depth): Appendix E.
  • Proposition 7 and Proposition 8 (topology baselines): Appendix F.
  • Theorem 11 (universal upper bound $s\leq\tfrac{1}{2}$): Appendix H.
  • Theorem 13 and Algorithm 1 (envelope construction): Appendix I.
Appendix G collects derivations for the continuous warm-up, and Appendix J provides experimental templates for estimating the effective parameters.
Appendix A Properties of Majority Maps
Throughout, let $b\geq 3$ be an odd integer and write $b=2k+1$ with $k\in\mathbb{Z}_{+}$. For $u\in[-1,1]$, define $p=(1+u)/2\in[0,1]$ and recall the majority map
$f_{b}(u)=2\,\Pr\!\Big(\mathrm{Bin}(b,p)\geq k+1\Big)-1.$
A.1 A derivative identity for binomial tails
We start with a calculus identity that makes monotonicity and concavity transparent.
Lemma 14 (Derivative of the symmetric binomial tail)
Let $b=2k+1$ and define
$F(p):=\Pr\!\big(\mathrm{Bin}(b,p)\geq k+1\big)=\sum_{j=k+1}^{2k+1}\binom{2k+1}{j}p^{j}(1-p)^{2k+1-j}.$
Then $F$ is differentiable on $(0,1)$ and
$F^{\prime}(p)=(2k+1)\binom{2k}{k}\,p^{k}(1-p)^{k}.$ (48)
Proof Differentiate term-by-term for $p\in(0,1)$. Using the identity $j\binom{2k+1}{j}=(2k+1)\binom{2k}{j-1}$ and $(2k+1-j)\binom{2k+1}{j}=(2k+1)\binom{2k}{j}$, we obtain
$\frac{d}{dp}\Big[\binom{2k+1}{j}p^{j}(1-p)^{2k+1-j}\Big]$ $=\binom{2k+1}{j}\Big(jp^{j-1}(1-p)^{2k+1-j}-(2k+1-j)p^{j}(1-p)^{2k-j}\Big)$
$=(2k+1)\Big[\binom{2k}{j-1}p^{j-1}(1-p)^{2k+1-j}-\binom{2k}{j}p^{j}(1-p)^{2k-j}\Big].$
Summing over $j=k+1,\dots,2k+1$ telescopes: the negative term at index $j$ cancels the positive term at index $j+1$. Only the first positive term (at $j=k+1$) remains, yielding
$F^{\prime}(p)=(2k+1)\binom{2k}{k}p^{k}(1-p)^{k}.$
A.2 Proof of Lemma 1
Proof [Proof of Lemma 1] We prove each item.
(1) Oddness, monotonicity, and endpoints.
Let $p=(1+u)/2$. Flipping the sign of the bias corresponds to sending $p\mapsto 1-p$, i.e., swapping “correct” and “incorrect” votes. By symmetry of the binomial distribution,
$\Pr\!\big(\mathrm{Bin}(b,1-p)\geq k+1\big)=\Pr\!\big(\mathrm{Bin}(b,p)\leq k\big)=1-\Pr\!\big(\mathrm{Bin}(b,p)\geq k+1\big),$
so $f_{b}(-u)=2(1-F(p))-1=-f_{b}(u)$, hence $f_{b}$ is odd. Also $f_{b}(0)=0$ and $f_{b}(1)=1$ are immediate from the definition.
To see that $f_{b}$ is increasing on $[0,1]$, note that $u\mapsto p=(1+u)/2$ is increasing and, for fixed $b$ and threshold $k+1$, the binomial tail $p\mapsto F(p)$ is increasing. Formally, Lemma 14 gives $F^{\prime}(p)\geq 0$ on $(0,1)$, hence $F$ is nondecreasing and so is $f_{b}$.
(2) Concavity on $[0,1]$.
Differentiate $f_{b}(u)=2F((1+u)/2)-1$. Since $dp/du=1/2$, we have
$f_{b}^{\prime}(u)=F^{\prime}\!\Big(\frac{1+u}{2}\Big).$ (49)
Plugging (48) into (49) yields, for $u\in(-1,1)$,
$f_{b}^{\prime}(u)=(2k+1)\binom{2k}{k}\Big(\frac{1+u}{2}\Big)^{k}\Big(\frac{1-u}{2}\Big)^{k}=\frac{(2k+1)\binom{2k}{k}}{2^{2k}}\,(1-u^{2})^{k}.$ (50)
The right-hand side is nonnegative and strictly decreasing for $u\in(0,1)$ (because $(1-u^{2})^{k}$ strictly decreases on $(0,1)$). Therefore $f_{b}^{\prime}$ is nonincreasing on $[0,1]$, which implies that $f_{b}$ is concave on $[0,1]$ (and in fact strictly concave on $(0,1)$ for $b\geq 3$).
(3) Derivative at the origin.
Setting $u=0$ in (50) gives
$f_{b}^{\prime}(0)=\frac{(2k+1)\binom{2k}{k}}{2^{2k}}=\frac{b\binom{b-1}{(b-1)/2}}{2^{b-1}},$
which is exactly (18).
A useful corollary: majority improves independent votes.
Concavity on $[0,1]$ with $f_{b}(0)=0$ and $f_{b}(1)=1$ implies $f_{b}(u)\geq u$ for all $u\in[0,1]$, with strict inequality for $u\in(0,1)$ because $f_{b}$ is strictly concave (for $b\geq 3$). We use this fact in Appendix C when treating the $\gamma=1$ case.
Appendix B Correlated Aggregation via the $\rho$-Shared Model
Recall Definition 2. Let $S_{i}:=V_{i}Y\in\{-1,+1\}$ denote signed correctness.
B.1 Relationship to pairwise correlation
Lemma 15 (Pairwise correlation equals $\rho$)
Under Definition 2, for any $i\neq j$,
$\mathrm{Corr}(S_{i},S_{j})=\rho.$
Proof By construction, $\mathbb{E}[S_{i}]=u$ for every $i$, so $\mathrm{Var}(S_{i})=1-u^{2}$. Moreover,
$\mathbb{E}[S_{i}S_{j}]=\rho\cdot 1+(1-\rho)\cdot\mathbb{E}[S_{i}]\mathbb{E}[S_{j}]=\rho+(1-\rho)u^{2}.$
Thus $\mathrm{Cov}(S_{i},S_{j})=\mathbb{E}[S_{i}S_{j}]-u^{2}=\rho(1-u^{2})$, and dividing by $\sqrt{\mathrm{Var}(S_{i})\mathrm{Var}(S_{j})}=1-u^{2}$ yields $\mathrm{Corr}(S_{i},S_{j})=\rho$.
B.2 Proof of Lemma 3
Proof [Proof of Lemma 3] Condition on the mixture event in Definition 2. With probability $\rho$, all child votes equal a common $V$ with $\mathbb{E}[VY]=u$, and majority returns $V$, hence output bias $u$. With probability $1-\rho$, the votes are i.i.d. with bias $u$, so the output bias is $f_{b}(u)$ by definition. Taking expectation over the mixture yields
$f_{b,\rho}(u)=\rho\,u+(1-\rho)\,f_{b}(u),$
which is (19). Since $f_{b}$ is increasing and concave on $[0,1]$ (Lemma 1), the mixture $f_{b,\rho}$ is also increasing and concave on $[0,1]$. Differentiating at the origin gives (20).
Appendix C Binary Phase Transition Proofs
We prove Theorem 4. Throughout, fix odd $b\geq 3$, $\rho\in[0,1)$, and $\gamma\in(0,1]$, and define
$T(u)=f_{b,\rho}(\gamma u)=\rho\gamma u+(1-\rho)f_{b}(\gamma u),\qquad u\in[0,1].$
By Lemma 1 and Lemma 3, $T$ is continuous, increasing, and concave on $[0,1]$, with $T(0)=0$. Because $1-\rho>0$ and $f_{b}$ is strictly concave on $(0,1)$, the map $T$ is strictly concave on $(0,1)$.
C.1 A ratio monotonicity lemma
Lemma 16 (Concavity implies decreasing ratio)
Let $T:[0,1]\to\mathbb{R}_{+}$ be concave with $T(0)=0$. Then the ratio $r(u):=T(u)/u$ is nonincreasing on $(0,1]$.
Proof Fix $0<u<v\leq 1$. Concavity gives
$T(u)=T\Big(\frac{u}{v}v+\Big(1-\frac{u}{v}\Big)0\Big)\geq\frac{u}{v}T(v)+\Big(1-\frac{u}{v}\Big)T(0)=\frac{u}{v}T(v).$
Dividing by $u>0$ yields $T(u)/u\geq T(v)/v$.
C.2 Proof of Theorem 4
Proof [Proof of Theorem 4] Let $\alpha_{\rho}=T^{\prime}(0)=\gamma f_{b,\rho}^{\prime}(0)$ as in (22). Define h(u):=T(u)uh(u):=T(u)-u. Since $T$ is concave, $h$ is concave (and in fact strictly concave on $(0,1)$) with $h(0)=0$ and $h^{\prime}(0)=\alpha_{\rho}-1$.
(1) Subcritical collapse: $\alpha_{\rho}\leq 1$.
If $\alpha_{\rho}\leq 1$, then $h^{\prime}(0)\leq 0$. Because $h$ is concave and $h(0)=0$, for any $u\in[0,1]$ we have the supporting-line bound
$h(u)\leq h(0)+h^{\prime}(0)u\leq 0.$
Thus $T(u)\leq u$ on $[0,1]$. The recursion $\mu_{t+1}=T(\mu_{t})$ is therefore monotone nonincreasing and bounded below by 0, hence it converges to some $\bar{\mu}\in[0,1]$. By continuity of $T$, the limit is a fixed point: $\bar{\mu}=T(\bar{\mu})$, i.e., $h(\bar{\mu})=0$.
We now show $\bar{\mu}=0$. If $\bar{\mu}>0$, then $h(\bar{\mu})=0$ and $h(u)\leq 0$ for all $u\in[0,1]$. Concavity then forces $h(u)=0$ for all $u\in[0,\bar{\mu}]$: indeed, for any $0<u<\bar{\mu}$,
$h(u)\geq\frac{u}{\bar{\mu}}h(\bar{\mu})+\Big(1-\frac{u}{\bar{\mu}}\Big)h(0)=0,$
while also $h(u)\leq 0$, hence $h(u)=0$. This contradicts strict concavity of $h$ on $(0,1)$ unless $\bar{\mu}=0$. Therefore $\mu_{t}\to 0$.
(2) Supercritical amplification: $\alpha_{\rho}>1$.
Assume $\alpha_{\rho}>1$, so $h^{\prime}(0)>0$ and hence $h(u)>0$ for all sufficiently small $u>0$. Next, we show that $h(1)\leq 0$, with strict inequality when $\gamma<1$. Indeed,
$h(1)=T(1)-1=f_{b,\rho}(\gamma)-1.$
If $\gamma=1$, then $f_{b,\rho}(1)=1$ so $h(1)=0$. If $\gamma<1$, then $0\leq\gamma<1$ and $f_{b}(\gamma)<1$ (majority cannot be perfectly correct when each vote has bias <1<1), hence $f_{b,\rho}(\gamma)=\rho\gamma+(1-\rho)f_{b}(\gamma)<\rho+(1-\rho)\cdot 1=1$, so $h(1)<0$.
Since $h$ is continuous, $h(0)=0$, $h(u)>0$ for small $u>0$, and $h(1)\leq 0$, there exists at least one root $\mu^{\star}\in(0,1]$ with $h(\mu^{\star})=0$, i.e., $T(\mu^{\star})=\mu^{\star}$. Uniqueness follows from strict concavity: if there were two distinct roots in $(0,1]$, then $h$ would have to be linear on the interval between them, contradicting strict concavity. Thus the fixed point $\mu^{\star}\in(0,1]$ is unique.
We now show global convergence to $\mu^{\star}$ for any $\mu_{0}\in(0,1]$. By Lemma 16, the ratio $r(u)=T(u)/u$ is nonincreasing on $(0,1]$. Because $r(0^{+})=\alpha_{\rho}>1$ and $r(\mu^{\star})=1$, we have $r(u)>1$ for $u\in(0,\mu^{\star})$ and $r(u)<1$ for $u\in(\mu^{\star},1]$. Equivalently,
$u\in(0,\mu^{\star})\Rightarrow T(u)>u,\qquad u\in(\mu^{\star},1]\Rightarrow T(u)<u.$
If $\mu_{0}\in(0,\mu^{\star})$, then $\{\mu_{t}\}$ is strictly increasing and bounded above by $\mu^{\star}$, hence converges to some $\bar{\mu}\leq\mu^{\star}$; by continuity, $\bar{\mu}$ is a fixed point, so $\bar{\mu}=\mu^{\star}$. If $\mu_{0}\in(\mu^{\star},1]$, then $\{\mu_{t}\}$ is strictly decreasing and bounded below by $\mu^{\star}$, hence converges to a fixed point $\bar{\mu}\geq\mu^{\star}$, again forcing $\bar{\mu}=\mu^{\star}$. This proves $\mu_{t}\to\mu^{\star}$ for all $\mu_{0}\in(0,1]$.
Finally, the statements about $\mu^{\star}$ follow from the above boundary cases: if $\gamma=1$, then $h(1)=0$ and the unique positive fixed point is $\mu^{\star}=1$; if $\gamma<1$, then $h(1)<0$ so $\mu^{\star}<1$.
A stability fact used later.
At the attracting fixed point $\mu^{\star}$, we necessarily have $T^{\prime}(\mu^{\star})<1$. Indeed, since $h(\mu^{\star})=0$ and $h$ is concave with $h(u)>0$ for $u<\mu^{\star}$, the tangent line at $\mu^{\star}$ must have strictly negative slope: $h^{\prime}(\mu^{\star})<0$, i.e., $T^{\prime}(\mu^{\star})<1$. We use this to obtain geometric convergence in Appendix E.
Appendix D Small-Signal Bounds and Budget Thresholds
This section proves Theorem 5, Theorem 9, and Corollary 10.
D.1 Linear-band bounds and proof of Theorem 5
Recall $T(u)=f_{b,\rho}(\gamma u)$, which is concave on $[0,1]$ with $T(0)=0$ and derivative $\alpha_{\rho}=T^{\prime}(0)$.
Lemma 17 (Tangent upper bound at the origin)
If $T$ is concave on $[0,1]$ and differentiable at 0, then
$T(u)\leq T(0)+T^{\prime}(0)\,u=\alpha_{\rho}\,u,\qquad u\in[0,1].$
Proof For concave functions, every tangent line is a global upper bound. In particular, differentiability at 0 implies that the affine function $u\mapsto T(0)+T^{\prime}(0)u$ supports $T$ at 0, hence upper-bounds $T$ on $[0,1]$.
Proof [Proof of Theorem 5] Fix $\eta\in(0,1)$. The upper inequality in (25) is Lemma 17. For the lower inequality, note that $T$ is differentiable at 0, so
$\lim_{u\downarrow 0}\frac{T(u)}{u}=T^{\prime}(0)=\alpha_{\rho}.$
Therefore, there exists $\delta\in(0,1]$ such that for all $u\in(0,\delta]$,
$\frac{T(u)}{u}\geq(1-\eta)\alpha_{\rho},$
which is exactly the lower bound in (25).
To obtain the iterative sandwich (26), assume $\mu_{0}\alpha_{\rho}^{L}\leq\delta$. By Lemma 17 and induction, $\mu_{t}\leq\mu_{0}\alpha_{\rho}^{t}\leq\mu_{0}\alpha_{\rho}^{L}\leq\delta$ for all $t\leq L$. Thus the lower bound in (25) applies at each step: $\mu_{t+1}=T(\mu_{t})\geq(1-\eta)\alpha_{\rho}\,\mu_{t}$, so again by induction $\mu_{t}\geq(1-\eta)^{t}\mu_{0}\alpha_{\rho}^{t}$. Setting $t=L$ gives (26).
D.2 Closed-form compute allocation and proof of Theorem 9
We treat the growth-regime surrogate objective in (33) as an exact function of $x>0$ for fixed $(B,b,m)$, ignoring multiplicative constants independent of $x$:
$\mu_{\mathrm{grow}}(x)\propto x^{\beta}(x+c_{0})^{-s},\qquad c_{0}=c_{0}(b,m)>0.$
Taking logs gives
$\phi(x):=\log\mu_{\mathrm{grow}}(x)=\beta\log x-s\log(x+c_{0})+\text{const}(B,b,m).$
Proof [Proof of Theorem 9] Differentiate $\phi$:
$\phi^{\prime}(x)=\frac{\beta}{x}-\frac{s}{x+c_{0}}=\frac{\beta(x+c_{0})-sx}{x(x+c_{0})}=\frac{\beta c_{0}-(s-\beta)x}{x(x+c_{0})}.$
If $s>\beta$, the unique critical point is at $x^{\star}=\frac{\beta}{s-\beta}c_{0}$, which is (34). A second derivative check shows it is a maximizer:
$\phi^{\prime\prime}(x)=-\frac{\beta}{x^{2}}+\frac{s}{(x+c_{0})^{2}}<0\qquad\text{at }x=x^{\star},$
because at $x^{\star}$ we have $\frac{s}{(x^{\star}+c_{0})^{2}}=\frac{s(s-\beta)^{2}}{s^{2}c_{0}^{2}}<\frac{\beta(s-\beta)^{2}}{\beta^{2}c_{0}^{2}}=\frac{\beta}{(x^{\star})^{2}}$. Equivalently, $\phi$ is strictly concave around $x^{\star}$.
Substituting $x^{\star}=\frac{\beta}{s-\beta}c_{0}$ into (33) and simplifying yields (35). Indeed,
$\frac{x^{\star}}{x^{\star}+c_{0}}=\frac{\beta}{s},\qquad x^{\star}+c_{0}=\frac{s}{s-\beta}c_{0},$
so
$x^{\star\beta}\left(\frac{1}{x^{\star}+c_{0}}\right)^{s}=\left(\frac{\beta}{s-\beta}c_{0}\right)^{\beta}\left(\frac{s-\beta}{s}\frac{1}{c_{0}}\right)^{s}=\beta^{\beta}(s-\beta)^{s-\beta}s^{-s}\,c_{0}^{\beta-s}=\kappa(s,\beta)\,c_{0}^{\beta-s}.$
If $s\leq\beta$, then $\phi^{\prime}(x)\geq 0$ for all $x>0$, so the growth surrogate is nondecreasing in $x$. Under the budget constraint $B\approx N(x+c_{0})$ with $N\geq 1$, the maximum feasible $x$ is on the order of $B-c_{0}$ (corresponding to $N=1$), i.e., “scale up rather than scale out.”
D.3 Proof of Corollary 10
Proof [Proof of Corollary 10] In the growth surrogate, Theorem 9 gives
$\mu_{\mathrm{grow}}^{\star}(B;b,m)\approx k\,\kappa(s,\beta)\,c_{0}^{\beta-s}\,B^{s}.$
The single-agent surrogate is $\mu_{\mathrm{single}}(B)\approx kB^{\beta}$. Requiring $\mu_{\mathrm{grow}}^{\star}(B;b,m)\geq\mu_{\mathrm{single}}(B)$ and canceling $k>0$ yields
$\kappa(s,\beta)\,c_{0}^{\beta-s}\,B^{s-\beta}\geq 1\qquad\Longleftrightarrow\qquad B^{s-\beta}\geq\kappa(s,\beta)^{-1}c_{0}^{s-\beta}.$
Since $s>\beta$, raising both sides to the power $1/(s-\beta)$ gives the sufficient condition
$B\geq c_{0}\,\kappa(s,\beta)^{-\frac{1}{s-\beta}}=B_{\mathrm{crit}}(b,m),$
which is (37).
D.4 A remark on the $s\leq\beta$ regime
When $s\leq\beta$, the growth surrogate cannot deliver an exponent advantage over the single-agent curve. More precisely, optimizing (33) over $x$ yields a prediction that scales at most as $B^{\beta}$ (and typically prefers $N\approx 1$). This is the sense in which the condition $s>\beta$ is necessary for scale-out exponent gains in the growth regime.
Appendix E Saturation and Mixing Depth
We prove Theorem 6. Assume the supercritical regime $\alpha_{\rho}>1$, so by Theorem 4 there is a unique attracting fixed point $\mu^{\star}\in(0,1]$ and $\mu_{t}\to\mu^{\star}$ for every $\mu_{0}>0$.
E.1 Two-phase mixing analysis and proof of Theorem 6
Proof [Proof of Theorem 6] If $\mu_{0}\geq\mu^{\star}$, then $\mu_{t}\geq\mu^{\star}$ for all $t\geq 0$ (the recursion is monotone decreasing toward $\mu^{\star}$ from above), and therefore $\mu_{L}\geq(1-\varepsilon)\mu^{\star}$ holds for every $L$ with $L_{\mathrm{mix}}(\varepsilon)=0$. We therefore focus on the case $\mu_{0}\in(0,\mu^{\star})$, where $\mu_{t}\uparrow\mu^{\star}$.
Phase I: reaching a constant fraction of $\mu^{\star}$.
Define $r(u)=T(u)/u$ for $u>0$, which is continuous on $(0,1]$ and nonincreasing by Lemma 16. Because $r(\mu^{\star})=1$ and $r(u)>1$ for $u\in(0,\mu^{\star})$, pick any reference level $u_{1}\in(0,\mu^{\star})$ (e.g., $u_{1}=\mu^{\star}/2$) and set
$\underline{\alpha}:=r(u_{1})=\frac{T(u_{1})}{u_{1}}>1.$
For any $u\in(0,u_{1}]$, monotonicity of $r$ gives $T(u)/u=r(u)\geq r(u_{1})=\underline{\alpha}$, hence $T(u)\geq\underline{\alpha}u$. As long as $\mu_{t}\leq u_{1}$, we therefore have $\mu_{t+1}\geq\underline{\alpha}\mu_{t}$, and by induction
$\mu_{t}\geq\mu_{0}\,\underline{\alpha}^{\,t}\qquad\text{until the first time }\mu_{t}>u_{1}.$
Thus the hitting time
$t_{1}:=\min\{t:\ \mu_{t}\geq u_{1}\}$
satisfies
$t_{1}\leq\left\lceil\frac{\log(u_{1}/\mu_{0})}{\log\underline{\alpha}}\right\rceil=O\!\left(\log\frac{1}{\mu_{0}}\right),$
with constants depending only on $u_{1}$ (hence on $b,\rho,\gamma$).
Phase II: geometric contraction near $\mu^{\star}$.
As noted at the end of Appendix C, we have $T^{\prime}(\mu^{\star})<1$. By continuity of $T^{\prime}$ and compactness, there exists $u_{0}\in(u_{1},\mu^{\star})$ and a constant $\kappa\in(0,1)$ such that
$\sup_{u\in[u_{0},\mu^{\star}]}T^{\prime}(u)\leq\kappa<1.$ (51)
Once $\mu_{t}\geq u_{0}$, define the gap $d_{t}:=\mu^{\star}-\mu_{t}\in[0,\mu^{\star}-u_{0}]$. By the mean value theorem, for each $t$ with $\mu_{t}\in[u_{0},\mu^{\star}]$ there exists $\xi_{t}\in[\mu_{t},\mu^{\star}]\subseteq[u_{0},\mu^{\star}]$ such that
$d_{t+1}=\mu^{\star}-\mu_{t+1}=T(\mu^{\star})-T(\mu_{t})=T^{\prime}(\xi_{t})\,(\mu^{\star}-\mu_{t})\leq\kappa\,d_{t}.$
Iterating yields $d_{t+n}\leq\kappa^{n}d_{t}$. In particular, once we have reached $u_{0}$, to ensure $d_{t+n}\leq\varepsilon\mu^{\star}$ it suffices that
$n\geq\left\lceil\frac{\log(d_{t}/(\varepsilon\mu^{\star}))}{\log(1/\kappa)}\right\rceil=O\!\left(\log\frac{1}{\varepsilon}\right),$
where the constant depends on $\kappa$ and the initial gap bound $d_{t}\leq\mu^{\star}$.
Combining phases.
Since $\mu_{t}\uparrow\mu^{\star}$, the time to go from $u_{1}$ to $u_{0}$ is finite and depends only on $(b,\rho,\gamma)$; absorb it into constants. Combining the bounds gives a depth
$L_{\mathrm{mix}}(\varepsilon)\leq C_{1}\log\frac{1}{\mu_{0}}+C_{2}\log\frac{1}{\varepsilon}+C_{3},$
with constants depending only on $(b,\rho,\gamma)$, such that $\mu_{L}\geq(1-\varepsilon)\mu^{\star}$ for all $L\geq L_{\mathrm{mix}}(\varepsilon)$. This establishes the claimed scaling.
E.2 Clipped objectives and approximation guarantees
The clipped proxy $\widehat{\mu}_{L}=\min\{\mu^{\star},\mu_{0}\alpha_{\rho}^{L}\}$ in (27) is most naturally interpreted for the regime $\mu_{0}\leq\mu^{\star}$, which includes the small-signal settings used in Section 5.4. In that case, monotonicity and Theorem 4 imply $\mu_{L}\leq\mu^{\star}$ for all $L$, while Lemma 17 implies $\mu_{L}\leq\mu_{0}\alpha_{\rho}^{L}$. Therefore $\mu_{L}\leq\widehat{\mu}_{L}$, justifying the global upper bound claim in Section 4.4. Theorems 5 and 6 provide matching lower bounds in the growth and saturation regimes, respectively.
Appendix F Topology Comparison Proofs
F.1 Star upper bounds under context constraints
A star node that reads $m$ tokens per leaf can aggregate at most $N_{\max}(m)=\lfloor W/m\rfloor$ leaves within a context budget $W$. This is the sense in which a star saturates under a finite context window: beyond $N_{\max}(m)$, additional leaves cannot be included without increasing $m$ (which changes both cost and fidelity).
F.2 Proof of Proposition 7
Proof [Proof of Proposition 7] Any predictor $\widehat{Y}=\phi(Z)$ induces a predictor on $X$ via composition $\psi(X):=\phi(Z(X,U))$, where $U$ denotes any protocol randomness (assumed independent of $Y$). Therefore, the set of predictors measurable with respect to $Z$ is a subset of those measurable with respect to $X$, implying the optimal risk using $X$ is no larger than the optimal risk using $Z$.
F.3 Proof of Proposition 8
Proof [Proof of Proposition 8] In the channel model of Section 2.3, each hop multiplies the incoming bias by $\gamma(m)$. After $L$ independent hops, the bias is $\mu_{0}\gamma(m)^{L}$.
F.4 Tree vs. star dominance beyond a threshold
The comparison between a context-limited star and a supercritical tree in Section 5 is driven by two facts: (i) the star cannot use more than $N_{\max}(m)$ leaves for fixed $m$, hence eventually must spend additional budget on scale-up (which grows like $B^{\beta}$ in the small-signal regime), while (ii) a tree with exponent $s>\beta$ admits a growth regime with $B^{s}$ scaling (Appendix D) before saturating at $\mu^{\star}$ (Appendix E). This yields a provable intermediate budget range where trees can dominate under context constraints, formalized through the sufficient condition $B\geq B_{\mathrm{crit}}$ in Corollary 10.
Appendix G Continuous Extension Proofs
This appendix derives the closed-form mean-squared error (MSE) recursions used in the continuous warm-up (Section 3). Throughout, let $Y\in[0,1]$ be the ground truth and suppose each leaf outputs $\widehat{Y}_{i}=Y+E_{i}$ with $\mathbb{E}[E_{i}]=0$, $\mathbb{E}[E_{i}^{2}]=v_{0}$, and $\mathrm{Corr}(E_{i},E_{j})=\rho$ for $i\neq j$.
G.1 Star, chain, and tree recursions
Star.
In the star topology, each leaf transmits $\widehat{Y}_{i}$ over the additive channel of Section 3, so the aggregator receives $\widehat{Y}_{i}+\eta_{i}$ with $\mathbb{E}[\eta_{i}]=0$ and $\mathbb{E}[\eta_{i}^{2}]=\sigma_{c}^{2}(m)$, independent across leaves and independent of the leaf errors $E_{i}$. The aggregator outputs the average of the received messages,
$\widehat{Y}_{\mathrm{star}}=\frac{1}{N}\sum_{i=1}^{N}(\widehat{Y}_{i}+\eta_{i})=Y+\bar{E}+\bar{\eta},\qquad\bar{E}:=\frac{1}{N}\sum_{i=1}^{N}E_{i},\;\;\bar{\eta}:=\frac{1}{N}\sum_{i=1}^{N}\eta_{i}.$
Because $\bar{E}$ and $\bar{\eta}$ are independent and mean-zero, the MSE is
$\mathbb{E}[(\bar{E}+\bar{\eta})^{2}]=\mathbb{E}[\bar{E}^{2}]+\mathbb{E}[\bar{\eta}^{2}]=v_{0}\Big(\rho+\frac{1-\rho}{N}\Big)+\frac{\sigma_{c}^{2}(m)}{N},$
which is (7).
Chain.
In the chain model of Section 3, each hop adds independent communication noise $C_{\ell}$ with $\mathbb{E}[C_{\ell}]=0$, $\mathbb{E}[C_{\ell}^{2}]=\sigma_{c}^{2}(m)$, so after $L$ hops the error variance is $v_{0}+L\sigma_{c}^{2}(m)$, giving (8).
Tree.
For a $b$-ary tree, each internal node averages the $b$ received child messages. Under the edge channel model of Section 3, child $j$ transmits $\widehat{Y}_{j}$ and the parent receives $\widehat{Y}_{j}+\eta_{j}$, where $\mathbb{E}[\eta_{j}]=0$ and $\mathbb{E}[\eta_{j}^{2}]=\sigma_{c}^{2}(m)$, independent across edges and independent of estimation errors. If the child errors have variance $v_{t}$ and pairwise correlation $\rho$, then averaging the $b$ child errors yields variance $v_{t}(\rho+(1-\rho)/b)$ (the same computation as the star with $N=b$). Averaging the $b$ independent channel noises adds variance $\sigma_{c}^{2}(m)/b$. Thus the recursion is
$v_{t+1}=v_{t}\Big(\rho+\frac{1-\rho}{b}\Big)+\frac{\sigma_{c}^{2}(m)}{b},$
which is (9). Solving this affine recursion gives (10), and taking $t\to\infty$ yields the floor (11).
G.2 Mixing depth in the continuous recursion
Let $\lambda:=\rho+(1-\rho)/b\in(0,1)$. From the closed form (10),
$v_{t}-v^{\star}=\lambda^{t}(v_{0}-v^{\star}),$
so reaching relative error $v_{t}\leq v^{\star}+\varepsilon(v_{0}-v^{\star})$ is equivalent to $\lambda^{t}\leq\varepsilon$, i.e., $t\geq\frac{\log(1/\varepsilon)}{\log(1/\lambda)}$, which is (13). This is the continuous analogue of the binary mixing-depth statement in Theorem 6.
Appendix H Universal Upper Bound on Amplification Exponents
We prove Theorem 11.
H.1 A standard bound on the central binomial coefficient
The proof relies on the classical inequality
$\binom{2k}{k}\leq\frac{4^{k}}{\sqrt{\pi k}},\qquad k\geq 1,$ (52)
which we derive here from an integral bound (avoiding asymptotic approximations).
Lemma 18 (Central binomial coefficient bound)
For every integer $k\geq 1$, $\binom{2k}{k}\leq\frac{4^{k}}{\sqrt{\pi k}}$.
Proof Let $I_{k}:=\int_{0}^{\pi/2}\sin^{2k}\theta\,d\theta$. A standard integration-by-parts recursion gives $I_{k}=\frac{2k-1}{2k}I_{k-1}$ with $I_{0}=\pi/2$, hence
$I_{k}=\frac{\pi}{2}\cdot\frac{(2k-1)!!}{(2k)!!}=\frac{\pi}{2}\cdot\frac{(2k)!}{2^{2k}(k!)^{2}}=\frac{\pi}{2}\cdot\frac{1}{4^{k}}\binom{2k}{k}$.
Equivalently,
$\binom{2k}{k}=\frac{2\cdot 4^{k}}{\pi}\,I_{k}.$ (53)
We next upper bound $I_{k}$. Change variables $x=\frac{\pi}{2}-\theta$ to get $I_{k}=\int_{0}^{\pi/2}\cos^{2k}x\,dx$. For $x\in[0,\pi/2]$, define $g(x)=\log(\cos x)+x^{2}/2$. Then $g(0)=0$ and $g^{\prime}(x)=-\tan x+x\leq 0$ because $\tan x\geq x$ for $x\geq 0$. Thus $g(x)\leq 0$ and $\cos x\leq e^{-x^{2}/2}$ on $0,\pi/2]$. Therefore,
$I_{k}=\int_{0}^{\pi/2}\cos^{2k}x\,dx\leq\int_{0}^{\pi/2}e^{-kx^{2}}\,dx\leq\int_{0}^{\infty}e^{-kx^{2}}\,dx=\frac{\sqrt{\pi}}{2\sqrt{k}}.$
Plugging this into (53) yields
$\binom{2k}{k}\leq\frac{2\cdot 4^{k}}{\pi}\cdot\frac{\sqrt{\pi}}{2\sqrt{k}}=\frac{4^{k}}{\sqrt{\pi k}},$
which is (52).
H.2 Proof of Theorem 11
Proof [Proof of Theorem 11] Let $b=2k+1$ with $k\geq 1$. By Lemma 1,
$f_{b}^{\prime}(0)=\frac{(2k+1)\binom{2k}{k}}{2^{2k}}=(2k+1)\binom{2k}{k}\,4^{-k}.$
Applying Lemma 18 yields
$f_{b}^{\prime}(0)\leq(2k+1)\cdot\frac{4^{k}}{\sqrt{\pi k}}\cdot 4^{-k}=\frac{2k+1}{\sqrt{\pi k}}.$
Since $\pi>3$ and $k\geq 1$, we have $\frac{2k+1}{\sqrt{\pi k}}\leq\sqrt{2k+1}=\sqrt{b}$, proving (38).
For the second claim, note that $\gamma(m)\leq 1$ and $\rho+(1-\rho)f_{b}^{\prime}(0)\leq\rho+(1-\rho)\sqrt{b}\leq\sqrt{b}$ because $\sqrt{b}\geq 1$. Therefore $\alpha_{\rho}=\gamma(m)\big(\rho+(1-\rho)f_{b}^{\prime}(0)\big)\leq\sqrt{b}$, and $s=\frac{\log\alpha_{\rho}}{\log b}\leq\frac{1}{2}$.
Appendix I Monotone Design Curves and Envelope Algorithm
This appendix proves Theorem 13 and justifies Algorithm 1 in Section 6.3.
I.1 Increasing differences and proof of Theorem 13
Fix fan-in $b$ and consider feasible $m\in\mathcal{M}_{b}=\{m\in\mathbb{Z}_{+}:\ bm\leq W\}$. For each feasible $m$ that passes the screening conditions $\alpha_{\rho}(b,m)>1$ and $s_{b,m}>\beta$, define
$\ell_{m}(t):=\log\underline{\mu}_{b}(m;e^{t})=s_{b,m}\,t+q_{b}(m),\qquad t=\log B,$
as in (45)–(44). By assumption, $\gamma(m)$ is strictly increasing in $m$, hence so is $\alpha_{\rho}(b,m)$, and therefore $s_{b,m}=\log\alpha_{\rho}(b,m)/\log b$ is strictly increasing in $m$.
Lemma 19 (Single-crossing / increasing differences)
For any $m_{2}>m_{1}$, the difference $\ell_{m_{2}}(t)-\ell_{m_{1}}(t)$ is strictly increasing in $t$.
Proof Compute
$\ell_{m_{2}}(t)-\ell_{m_{1}}(t)=(s_{b,m_{2}}-s_{b,m_{1}})\,t+\big(q_{b}(m_{2})-q_{b}(m_{1})\big).$
Because $s_{b,m_{2}}>s_{b,m_{1}}$, the coefficient of $t$ is positive, so the difference is strictly increasing in $t$.
Proof [Proof of Theorem 13] Let $t=\log B$ and define the argmax correspondence
$\mathcal{A}(t):=\arg\max_{m\in\mathcal{M}_{b}}\ \ell_{m}(t).$
By finiteness of $\mathcal{M}_{b}$, $\mathcal{A}(t)$ is nonempty. We show that the minimal selection $m^{-}(t):=\min\mathcal{A}(t)$ is nondecreasing in $t$, which implies the theorem.
Take $t_{2}>t_{1}$ and let $m_{1}=m^{-}(t_{1})\in\mathcal{A}(t_{1})$ and $m_{2}=m^{-}(t_{2})\in\mathcal{A}(t_{2})$. Suppose for contradiction that $m_{2}<m_{1}$. Optimality gives
$\ell_{m_{1}}(t_{1})\geq\ell_{m_{2}}(t_{1}),\qquad\ell_{m_{2}}(t_{2})\geq\ell_{m_{1}}(t_{2}).$
Adding these inequalities yields
$\ell_{m_{1}}(t_{1})+\ell_{m_{2}}(t_{2})\ \geq\ \ell_{m_{2}}(t_{1})+\ell_{m_{1}}(t_{2}).$ (54)
On the other hand, Lemma 19 with $m_{1}>m_{2}$ and $t_{2}>t_{1}$ implies
$\big(\ell_{m_{1}}(t_{2})-\ell_{m_{2}}(t_{2})\big)>\big(\ell_{m_{1}}(t_{1})-\ell_{m_{2}}(t_{1})\big),$
which rearranges to
$\ell_{m_{1}}(t_{2})+\ell_{m_{2}}(t_{1})\ >\ \ell_{m_{1}}(t_{1})+\ell_{m_{2}}(t_{2}),$
contradicting (54). Therefore $m_{2}\geq m_{1}$, i.e., $m^{-}(t)$ is nondecreasing. Translating back from $t=\log B$ proves that an optimal message length can be chosen nondecreasing in $B$.
I.2 Upper-envelope construction and correctness of Algorithm 1
We sketch a standard correctness argument for the stack-based envelope construction used in Algorithm 1. Consider a finite family of lines $\ell_{i}(t)=s_{i}t+q_{i}$ with strictly increasing slopes $s_{1}<s_{2}<\cdots<s_{M}$. Define the upper envelope $E(t)=\max_{i}\ell_{i}(t)$.
Lemma 20 (Envelope structure for increasing slopes)
If $s_{i}$ are strictly increasing, then there exist indices $1\leq i_{1}<i_{2}<\cdots<i_{r}\leq M$ and breakpoints $\infty=\tau_{0}<\tau_{1}<\cdots<\tau_{r}=\infty$ such that for every $j\in\{1,\dots,r\}$,
$E(t)=\ell_{i_{j}}(t)\quad\text{for all }t\in[\tau_{j-1},\tau_{j}).$
That is, the maximizing index is nondecreasing in $t$ and changes only at increasing intersection points.
Proof For any pair $i<j$, the difference $\ell_{j}(t)-\ell_{i}(t)$ is an increasing affine function of $t$ (because $s_{j}-s_{i}>0$). Therefore, the set where $\ell_{j}$ dominates $\ell_{i}$ is either empty or a half-line $t^{\star},\infty)$, where $t^{\star}$ is their intersection. This single-crossing property implies the argmax index over $\{\ell_{i}\}$ is nondecreasing in $t$, and the envelope is piecewise linear with breakpoints at pairwise intersections.
Algorithm 1 is a constructive implementation of Lemma 20. Processing lines in increasing slope order, it maintains a stack of candidate lines $(\ell_{i_{1}},\dots,\ell_{i_{r}})$ and the associated increasing breakpoints $(\tau_{1},\dots,\tau_{r-1})$. When a new line $\ell$ is added, if its intersection with the current top occurs to the left of the top’s activation breakpoint, then the top line is never maximal and is removed. Because each line is pushed and popped at most once, the construction runs in linear time. The resulting stack and breakpoint list exactly encode the envelope $E(t)$, and the associated maximizing message length is the piecewise-constant function $m_{b}^{\star}(B)$ in Theorem 13.
Appendix J Calibration and evaluation templates
This appendix collects optional templates for estimating the effective parameters of the framework in a given agent stack and for running small diagnostic evaluations. Because the main paper is theoretical and the modeling assumptions are deliberately simplified, these templates are intended primarily to support qualitative regime checks (e.g., whether $\alpha_{\rho}$ is supercritical, or whether saturation is communication- versus correlation-limited), rather than to provide a faithful simulator of all real-world effects.
J.1 Calibration protocol: estimating $\beta$, $\gamma(m)$ / $\sigma_{c}^{2}(m)$, and $\rho$
This subsection operationalizes the “measurable parameters” promise from Section 2.7. The goal is not to fit a complicated model, but to obtain stable estimates that are accurate enough to place the system in the correct qualitative regime of the phase diagram.
Single-agent scaling ($\beta$).
Fix a leaf-agent template (model, prompt, tools) and define a compute knob $x$ (e.g., reasoning tokens, number of samples, number of tool calls). Evaluate the leaf agent on calibration tasks at a grid $x_{1}<\cdots<x_{K}$ (roughly log-spaced).
For binary tasks, estimate accuracy $\widehat{p}(x_{k})$ and convert to bias $\widehat{\mu}(x_{k})=2\widehat{p}(x_{k})-1$. Fit a power law on a small-signal window where $\widehat{\mu}$ is positive and not saturated, e.g.,
$\log\widehat{\mu}(x_{k})\approx\log k+\beta\log x_{k},\qquad k\in\mathcal{K},$
with $\mathcal{K}$ chosen so that $\widehat{\mu}(x_{k})\in[\mu_{\min},\mu_{\max}]$ for task-appropriate constants $\mu_{\min}>0\ll\mu_{\max}<1$. Report $\widehat{\beta}$ (and optionally $\widehat{k}$) with bootstrap confidence intervals over tasks. For continuous tasks, estimate $v_{0}(x_{k})=\mathbb{E}[(\widehat{Y}-Y)^{2}]$ and fit $\log\widehat{v}_{0}(x_{k})\approx\log c-\beta\log x_{k}$ over a range where scaling is stable.
Communication fidelity ($\gamma(m)$) or distortion ($\sigma_{c}^{2}(m)$).
Communication is calibrated with a purpose-built transmission task that isolates encoding/decoding from problem solving, using the same message schema and truncation rules as the target pipeline.
For binary tasks, sample $Y\in\{-1,+1\}$ uniformly, ask a sender to encode $Y$ into at most $m$ tokens, and ask a receiver to decode $\widehat{Y}$ from the message alone. The empirical flip rate $\widehat{\epsilon}(m)=\Pr(\widehat{Y}\neq Y)$ yields $\widehat{\gamma}(m)=1-2\widehat{\epsilon}(m)$. For continuous tasks, sample $Y\in[0,1]$, transmit under an $m$-token constraint, decode $\widehat{Y}$, and estimate $\widehat{\sigma}_{c}^{2}(m)=\mathbb{E}[(\widehat{Y}-Y)^{2}]$.
Shared-failure correlation ($\rho$).
Correlation should be measured at the level where aggregation occurs (siblings), since that is the dependence that limits cancellation. On tasks $t=1,\dots,T$ with ground truth $Y(t)$, run $M$ leaf agents (different randomness and/or prompts). For binary tasks define signed correctness $S_{i}(t)=\widehat{Y}_{i}(t)Y(t)\in\{-1,+1\}$, compute pairwise sample correlations across tasks
$\widehat{\rho}_{ij}:=\widehat{\mathrm{Corr}}_{t\in[T]}\big(S_{i}(t),S_{j}(t)\big),\qquad 1\leq i<j\leq M,$
and aggregate
$\widehat{\rho}:=\frac{2}{M(M-1)}\sum_{1\leq i<j\leq M}\widehat{\rho}_{ij}.$
For continuous tasks, use residuals $E_{i}(t)=\widehat{Y}_{i}(t)-Y(t)$ and compute residual correlations analogously. In practice it can be informative to stratify tasks by difficulty (e.g., baseline accuracy bins), since $\rho$ may increase on harder tasks.
Budget accounting and effective context.
Finally, record realized costs: tokens generated and tokens read (including system prompts, tool outputs, and formatting overhead), and verify the effective context window $W$ after templating. This prevents a common pitfall: attributing gains to “topology” when the true driver is uneven budget usage.
[…]
Figure 12: Calibration experiments (schematic). Left: estimate the single-agent scaling exponent $\beta$ by fitting a power law in a non-saturated range. Middle: estimate communication fidelity as a function of message length $m$ via a transmission task. Right: estimate shared-error correlation $\rho$ and verify that diversity interventions reduce it.
J.2 LLM diagnostic template: context limits, correlation floors, and saturation
The synthetic setting checks internal consistency under the assumed generative models. The more important question is whether the framework remains useful as an engineering diagnostic when agents are real LLM instances with non-Gaussian errors and prompt effects. Our aim here is therefore qualitative: do calibrated parameters correctly identify which lever is binding and which intervention should help?
We recommend using at least one binary family with unambiguous correctness (e.g., multiple-choice, exact-answer math, pass/fail tests) and one continuous family with graded scores (e.g., rubric grading or scalar estimation with known targets), since the framework makes distinct predictions in each. Implement single-agent, star, tree (and optionally chain) pipelines under matched realized budgets, standardizing prompts and message schemas across topologies. For stars, feasibility is $Nm\leq W$; for trees, feasibility is $bm\leq W$, with depth and leaf count chosen using the design rules in Section 6.
In this setting, the most useful validations are diagnostic. First, under fixed $W$, the best feasible star should saturate in effective parallelism once $N$ hits $\lfloor W/m\rfloor$, so additional budget primarily benefits scale-up. Second, measured $\widehat{\rho}$ should help explain why naive voting sometimes fails: high $\widehat{\rho}$ predicts limited aggregation gains even with many agents. Third, the calibrated quantity $\widehat{\alpha}_{\rho}=\widehat{\gamma}(m)(\widehat{\rho}+(1-\widehat{\rho})f_{b}^{\prime}(0))$ should act as a feasibility check for trees: when $\widehat{\alpha}_{\rho}\leq 1$, deeper hierarchies should become unstable or drift toward chance; when $\widehat{\alpha}_{\rho}>1$, trees should amplify and then saturate. Finally, when performance plateaus with depth or budget, the framework predicts that further gains come primarily from improving communication fidelity (longer messages, better schemas) or reducing shared failures (diversity, verification), rather than adding depth.
Diagnostic interventions (compact ablations).
Three inexpensive interventions connect the theory’s parameters to practice under matched budgets: varying message length $m$ (trading off $\gamma(m)$ against coordination cost), injecting diversity to reduce $\widehat{\rho}$ (e.g., prompt, model, temperature, tool-policy variation or explicit verification roles), and swapping topology (single, star, chain, tree) under fixed schemas. When these interventions move $\widehat{\gamma}(m)$ or $\widehat{\rho}$ in the expected direction, the phase-diagram predictions provide testable hypotheses about whether amplification becomes feasible and whether scale-out can outpace scale-up before saturation.
Logging and reproducibility.
For each run, log realized tokens per node, realized message lengths, and outputs at each level (or hashes if privacy is needed). These logs enable post hoc estimation of $\rho$, auditing of budget fairness, and direct diagnosis of failure modes such as repeated identical errors.
These templates are meant to support qualitative diagnostics: whether context is binding, whether measured correlation suggests limited gains, and whether the calibrated $\widehat{\alpha}_{\rho}$ places a hierarchy in the amplifying or collapsing regime. In that sense, the phase diagrams and design curves from Sections 56 provide concrete hypotheses about which organization should help under matched budgets and which lever (communication fidelity versus shared-failure reduction) is most likely to matter when it does not. Establishing quantitative predictiveness in realistic deployments, where prompts, tools, and dependence structure may drift with depth and task difficulty, is an important direction for future work.

### v2

@@
 Abstract
-Multi-agent systems can improve reliability, yet under a fixed inference budget they often help, saturate, or even collapse.
+Multi-agent systems can improve reliability, yet under a fixed test-time budget they often help, saturate, or even collapse.
 We develop a deliberately minimal and calibratable theory that predicts these regimes from three binding constraints of modern agent stacks: finite context windows, lossy inter-agent communication, and shared failures among similar agents.
@@
 1 Introduction
-Multi-agent systems are often sold as a reliability primitive: if a single agent is fallible, run many agents, let them interact, and aggregate their outputs.
-Under a fixed inference budget, however, multi-agent coordination is not reliably beneficial.
-The same additional agents that sometimes improve performance can also yield saturation or outright degradation.
-This paper develops a predictive theory of this brittleness: when multi-agent organization produces budgeted synergy (i.e., outperforming any single agent under the same budget) and when scaling out must fail.
-The LLM era has made the phenomenon impossible to ignore.
-It is now routine to instantiate multiple LLM-based agents, assign roles, and coordinate them through discussion, critique, or voting in “agent societies” and orchestration toolkits (Brown et al., 2020; Yao et al., 2023; Wu et al., 2023; Li et al., 2023; Chen et al., 2024; Hong et al., 2024).
+Multi-agent systems are often presented as a reliability primitive: if a single agent is fallible, run many agents, let them interact, and aggregate their outputs.
+Under a fixed computational budget per task, however, multi-agent coordination is not reliably beneficial.
+The same additional agents that sometimes improve performance can also lead to saturation or even outright degradation.
+This paper develops a predictive theory of this brittleness: when multi-agent organization produces budgeted synergy, i.e., outperforming the best single agent under the same total computational budget, and when scaling out must fail.
+The LLM era has made this phenomenon impossible to ignore.
+It is now routine to instantiate multiple LLM-based agents, assign roles, and coordinate them through discussion, critique, or voting in so-called “agent societies” and orchestration frameworks (Brown et al., 2020; Yao et al., 2023; Wu et al., 2023; Li et al., 2023; Chen et al., 2024; Hong et al., 2024).
 The underlying intuition is classical: ensembling can reduce error (Breiman, 1996; Freund and Schapire, 1997; Dietterich, 2000), and interaction protocols such as debate or amplification aim to turn many weak judgments into a stronger one (Irving et al., 2018; Christiano et al., 2018; Du et al., 2024).
-At the same time, controlled matched-budget evaluations increasingly report strong task- and topology-dependence, including regimes where multi-agent variants degrade once coordination overhead and information loss dominate (Kim et al., 2025).
-Our starting point is that these failures are not implementation accidents but structural consequences of three constraints that repeatedly bind in modern agent stacks.
-First, agent errors are not independent: many instances derived from the same base model and prompt share failure modes, producing “groupthink” rather than error cancellation.
-Second, coordination is communication-limited: as systems scale, agents must compress reasoning into bounded-length messages, and this loss compounds across layers.
+At the same time, controlled matched-budget evaluations increasingly reveal strong task- and topology-dependence, including regimes in which multi-agent variants degrade once coordination overhead and information loss dominate (Kim et al., 2025).
+Our starting point is that these failures are not implementation accidents but structural consequences of a small number of constraints that repeatedly bind in modern agent stacks.
+First, agent errors are not independent: multiple instances derived from the same base model and prompt often share systematic failure modes, producing “groupthink” rather than error cancellation.
+Second, coordination is communication-limited: as systems scale, agents must compress reasoning into bounded-length messages, and the resulting information loss compounds across layers.
 Third, coordination is context-limited: a central aggregator cannot read arbitrarily many messages under a finite context window, so naive scale-out saturates structurally.
-Recent work on large-population collaboration and topology learning highlights the same bottleneck: token and context limits are often the binding resource (Qian et al., 2025; Zhang et al., 2024).
-Today, agent-system design still largely proceeds by heuristics: choose a topology, message format, and number of agents, then iterate.
+Recent work on large-population collaboration and topology learning highlights the same bottleneck: token and context limits are frequently the binding resource (Qian et al., 2025; Zhang et al., 2024).
+Today, most agent-system design still proceeds by heuristics: choose a topology, message format, and number of agents, then iterate.
 In spirit, this resembles the hand-crafted feature engineering era of machine learning: powerful components exist, but system-level behavior depends on ad hoc choices and expensive trial-and-error.
-Our goal is to push multi-agent design toward an “agentic intelligence from first principles” style: make the dominant constraints explicit, explain why scaling out works or fails, and derive sharp, quantitative regime boundaries—in particular, a phase transition for budgeted multi-agent synergy.
+Our goal is to push multi-agent design toward an “agentic intelligence from first principles” approach: make the dominant constraints explicit, explain why scaling out works or fails under a fixed computational budget, and derive sharp, quantitative regime boundaries, most notably, a phase transition for budgeted multi-agent synergy.
 A minimal, calibratable coordination theory.
@@
 At the same time, the abstraction is broader than LLMs: none of the results relies on language as such, only on (a) how single-agent quality scales with compute, (b) how much usable information survives bounded-length communication, and (c) how strongly agents’ failures co-move.
+Figure 1: When does multi-agent organization produce budgeted synergy?
+An overview of our theoretical framework and core contribution.
+Left: scaling out multi-agent systems under different topologies (chain, star, and hierarchical tree) subject to a finite context window $W$, which constrains fan-in globally for stars ($Nm\leq W$) and locally for trees ($bm\leq W$).
+Bottom: three binding constraints of modern agent stacks: finite context windows, lossy communication with fidelity $\gamma(m)$, and shared-failure correlation $\rho$.
+Right: the competing baseline of scaling up a single agent, governed by a single-agent scaling law with exponent $\beta$.
+Our theory combines these elements into an effective per-layer gain $\alpha_{\rho}$, which induces an organization exponent $s$ for hierarchical aggregation.
+Budgeted synergy occurs precisely when $s>\beta$, meaning that scaling out via organization can outperform scaling up a single agent under the same total budget.
 By calibratable we do not mean a theory “about measurement.” We mean a coordination theory parameterized by a small set of environment-dependent effective quantities that can be estimated once (via targeted experiments) and then used to predict qualitative regimes.
@@
 Finally, we make precise why topology becomes decisive only because constraints bind: absent context limits, central aggregation is information-dominant (data processing inequality) (Cover and Thomas, 2006), and hierarchy is valuable only insofar as it bypasses fan-in limits without crossing into the collapsing regime.
+Figure 1 illustrates our core contribution.
 Beyond the binary core, a continuous-performance warm-up yields closed-form risks for star/chain/tree organizations and makes correlation and communication floors explicit.
@@
 Methodological stance: theory vs. benchmarking.
-This is a theory paper: our aim is to delineate regimes of behavior and identify the mechanisms that govern when multi-agent systems help, saturate, or collapse under realistic constraints.
+Our aim in this paper is to theoretically delineate regimes of behavior and identify the mechanisms that govern when multi-agent systems help, saturate, or collapse under realistic constraints.
 Large-scale matched-budget evaluations with real LLM agent stacks are expensive and can be confounded by prompt engineering and rapidly changing model/tool ecosystems.
@@
 Section 3 develops the continuous warm-up.
-Section 4 presents the binary theory, including the phase transition, small-signal amplification, and saturation.
-Section 5 compares topologies under context constraints and derives budget thresholds.
+Section 4 presents the binary theory, including the phase transition, small-signal amplification, and saturation; it also compares topologies under context constraints and derives budget thresholds.
 Section 5 turns the theory into design diagnostics.
@@
 2.1 Tasks and performance metrics
-We study two task primitives that cover many MAS evaluations and isolate the coordination bottlenecks above.
-Binary success/failure captures accuracy-like outcomes (pass rates, accept/reject decisions).
-Continuous performance on $[0,1]$ captures graded scoring or calibrated estimates and provides a smooth warm-up where many expressions admit closed forms.
-The continuous setting is not a substitute for the binary theory; it makes correlation and communication floors explicit without the nonlinearity introduced by majority vote.
+We study two task primitives that are deliberately simple but theoretically expressive.
+They are not meant to exhaust the diversity of real-world tasks; rather, they isolate the coordination bottlenecks that arise repeatedly inside more complex multi-agent pipelines.
+Binary success/failure captures decision-like outcomes (e.g., accept/reject, pass/fail, correctness of a proposed action), while continuous performance on $[0,1]$ captures graded judgments, scores, or calibrated estimates.
+Together, these two primitives cover a wide range of evaluation signals used in practice and allow clean analysis of aggregation, communication loss, and shared failures.
+From a theoretical standpoint, any richer task that can be decomposed into local scoring, communication, and aggregation steps must confront the same coordination constraints analyzed here.
+The continuous setting serves as a smooth warm-up where many quantities admit closed forms and correlation and communication floors are explicit.
+The binary setting is not a special case of the continuous one: the nonlinearity of majority vote introduces qualitatively new behavior, including a sharp amplification–collapse phase transition that has no analogue in linear aggregation.
+Many realistic agent workflows combine both primitives.
+For example, agents may exchange graded scores or confidence estimates internally, but the final system decision is binary.
+So analyzing both is necessary to capture the full coordination picture.
 Binary tasks.
 […]
-$\mu\;:=\;\mathbb{E}[\widehat{Y}Y]\in[-1,1],\qquad\Pr(\widehat{Y}=Y)=\frac{1+\mu}{2}.$
-The bias $\mu$ is convenient for multi-layer analysis because it composes cleanly under simple channel models and aggregation maps.
+$\mu\;:=\;\mathbb{E}[\widehat{Y}Y]\in[-1,1],\qquad\Pr(\widehat{Y}=Y)=\frac{1+\mu}{2}.$ (1)
+The bias $\mu$ is a natural summary statistic for weak agents: it measures how far performance is above chance and composes cleanly under simple channel models and aggregation maps.
+This makes it particularly suitable for analyzing whether small, local improvements are amplified or destroyed by hierarchical organization.
 Continuous tasks.
 Let $Y\in[0,1]$ and a system output $\widehat{Y}\in[0,1]$. We use mean squared error (MSE)
-$v\;:=\;\mathbb{E}\big[(\widehat{Y}-Y)^{2}\big],$
+$v\;:=\;\mathbb{E}\big[(\widehat{Y}-Y)^{2}\big],$ (2)
 and occasionally report a bounded performance score such as $\mathrm{Perf}=1/(1+v)$, which is monotone in $v$.
-The continuous analysis yields explicit floors; the binary analysis additionally introduces a sharp amplification–collapse transition.
+Continuous metrics capture graded information that is often exchanged within agent systems (scores, confidence levels, value estimates) and lead to linear aggregation dynamics.
+In this setting, the effect of correlation and communication loss appears as explicit error floors rather than phase transitions, providing intuition that complements the binary analysis.
 2.2 Single-agent capability and scaling
-Our central comparison is fixed-budget: the same total budget $B$ can be spent on fewer stronger agents (scale up) or more weaker agents (scale out).
-We model per-leaf compute by a single knob $x\geq 0$.
-In an LLM agent, $x$ can represent a token budget for reasoning, the number of samples or self-consistency votes, the number of tool calls, or any additive resource that trades off against parallelism.
-We summarize the effect of compute via a single-agent scaling law, motivated by empirical power-law relations between performance and compute observed for large language models and related systems (Kaplan et al., 2020; Hoffmann et al., 2022).
-For binary tasks, a leaf with compute $x$ produces a vote $\widehat{Y}$ with bias
+Our central comparison is fixed budget: the same total computational budget $B$ can be allocated either to scale up a single agent (making one agent stronger) or to scale out across many agents (running more agents in parallel).
+Understanding when scale out can outperform scale up under the same budget is the core question of this paper.
+We model per-leaf compute by a single nonnegative knob $x\geq 0$, representing the share of the per-task computational budget allocated to an individual agent.
+In LLM-based agents, $x$ may correspond to inference tokens devoted to reasoning, the number of samples or self-consistency votes, structured test-time inference procedures (e.g., generate–select or tournament-style algorithms), the number of tool calls, or any other additive per-task resource that trades off directly against parallelism.
+We summarize the effect of per-agent compute via a single-agent scaling law.
+Classical scaling laws are typically reported for training loss as a function of training compute (Kaplan et al., 2020; Hoffmann et al., 2022).
+Here, however, we use scaling laws in a different but related sense: as an effective description of how a fixed trained agent’s test-time performance improves when allocated additional per-task computational resources.
+Recent work on test-time or inference-time scaling shows that such improvements can follow systematic power-law or exponential trends in appropriate regimes and metrics, including accuracy, error probability, or calibrated scores (Chen et al., ; Levi, 2025; Snell et al., 2024; Wu et al., 2025).
+Motivated by these findings, we treat the scaling exponent as a per-task scaling exponent governing scale up, not as a statement about retraining the model.
+For binary success/failure tasks, a leaf agent allocated compute $x$ produces a vote $\widehat{Y}$ with bias
 $\mu_{0}(x)=g(x),$
-where $g(\cdot)$ is increasing and typically exhibits diminishing returns.
-For continuous tasks, a leaf produces an estimate $X$ with conditional variance $v_{0}(x)$, where $v_{0}(\cdot)$ decreases with $x$.
+where $g(\cdot)$ is increasing and exhibits diminishing returns. For continuous tasks, a leaf produces an estimate $X$ with conditional variance $v_{0}(x)$, where $v_{0}(\cdot)$ decreases with $x$.
 In our main theorems, we work in a regime where these functions are well-approximated by power laws:
@@
 with constants $k,c>0$ and a single-agent scaling exponent $\beta>0$.
-The exponent $\beta$ is not a universal constant.
-It depends on the model, prompting, tools, and the task family, and it is intended to be measured.
+Here “small-signal regime” refers to the operating range of $x$ in which (i) the local power-law approximation is accurate enough for comparison, and (ii) the induced leaf performance is not already saturated.
+In the binary analysis, this typically corresponds to weak leaf votes (small bias $\mu_{0}(x)$, i.e., accuracy only slightly above chance), which is precisely the regime where the majority update map is well-approximated by its derivative at the origin and where the organization exponent $s$ governs growth.
+In the continuous warm-up, it corresponds to the pre-saturation range where aggregation has not yet hit the communication/correlation floor.
+The exponent $\beta$ is not a universal constant: it depends on the model, prompting, tools, test-time procedure, and task family, and is intended to be measured in the operating regime of interest.
+Our theory does not require the power-law approximation to hold globally or asymptotically; it relies only on the existence of a local scaling exponent $\beta$ in the budget range where the scale-up versus scale-out comparison is made.
 2.3 Communication as a controllable lossy channel
@@
 For binary tasks and majority-style protocols, it is natural to treat each child as intending to transmit a single bit of information (its vote or decision), encoded into a message of length $m$.
-We model the receiver’s decoded bit as passing through a binary symmetric channel (BSC) with reliability $\gamma(m)\in(0,1]$: if a child has bias $\mu$ relative to $Y$, then after transmission the received bit has bias $\gamma(m)\mu$.
-Equivalently, $\gamma(m)=1-2\epsilon(m)$ where $\epsilon(m)$ is an effective flip probability.
+We model the receiver’s decoded bit as passing through a binary symmetric channel (BSC) with reliability $\gamma(m)\in(0,1]$:
+$\mu_{\text{recv}}=\gamma(m)\,\mu.$ (5)
+Equivalently, $\gamma(m)=1-2\epsilon(m)$, where $\epsilon(m)$ is an effective flip probability.
 Continuous: additive distortion.
 For continuous tasks, we model transmission as additive zero-mean distortion,
-$\widetilde{X}=X+\eta,\qquad\mathbb{E}[\eta]=0,\quad\mathrm{Var}(\eta)=\sigma_{c}^{2}(m),$
+$\widetilde{X}=X+\eta,\qquad\mathbb{E}[\eta]=0,\quad\mathrm{Var}(\eta)=\sigma_{c}^{2}(m),$ (6)
 where $\sigma_{c}^{2}(m)$ decreases with $m$. This captures compression loss, imperfect interpretation, and degradation introduced by summarization or constrained prompts.
@@
 Let $W$ denote the maximum number of tokens that an internal node can read and process. If each incoming message has length $m$, then any aggregation node can read at most
-$\text{fan-in}\leq\Big\lfloor\frac{W}{m}\Big\rfloor.$
+$\text{fan-in}\;\leq\;\Big\lfloor\frac{W}{m}\Big\rfloor.$ (7)
 This inequality is the reason depth matters: a star is capped by $N\leq\lfloor W/m\rfloor$ leaves, while a tree can scale to $N=b^{L}$ leaves by distributing the fan-in constraint across levels.
@@
 In particular, we focus on three canonical organizations that recur in practice and isolate the key trade-offs: star, chain, and hierarchical tree (Figure 1).
-[…]
-Figure 1: Canonical organizations under a context window constraint $W$.
-A star requires the root to read all messages, so feasibility enforces $Nm\leq W$.
-A hierarchical tree enforces the constraint locally ($bm\leq W$), enabling many more leaves at the cost of multi-hop communication and repeated aggregation.
-Chains avoid fan-in bottlenecks but compound communication loss.
 Star.
@@
 Trees are the simplest family that bypasses the fan-in cap ($bm\leq W$ locally) while still aggregating many signals; this is precisely the setting in which phase transitions and budget thresholds emerge.
+Figure 2: From general DAGs to spanning arborescences.
+A general DAG (a) may contain extra edges (dashed) that add redundancy or dependence, but any feed-forward organization contains a spanning arborescence that routes information from leaves to the output (b).
 It is also useful to view more general feed-forward organizations through these motifs.
@@
 Additional edges or multi-round interaction are valuable mainly when they effectively change the measurable parameters, for example by improving effective fidelity through redundancy or by reducing effective correlation through independent verification, rather than by rerouting the same information along multiple dependent paths.
+Figure 2 illustrates this reduction from general DAGs to a spanning arborescence.
 Aggregation rules.
@@
 Table 1: Key parameters of the framework and indicative estimation strategies. The theory treats these quantities as inputs; calibration is optional and necessarily coarse.
-[…]
 Figure 3: A possible use of the framework.
@@
 3.1 Model assumptions for the continuous setting
-Let $Y\in[0,1]$ denote the latent truth and let each leaf output an estimate $X\in[0,1]$.
-To keep the warm-up focused on organization rather than estimator bias, we assume leaf estimates are conditionally unbiased:
-$\mathbb{E}[X\mid Y]=Y.$
-We write the leaf residual as E:=XYE:=X-Y, so $\mathbb{E}[E\mid Y]=0$. Leaf quality is summarized by the conditional MSE,
-$v_{0}(x):=\mathbb{E}\big[(X-Y)^{2}\big]=\mathbb{E}\big[E^{2}\big],$
-which depends on per-leaf compute $x$. In the scaling regime of interest we model
+This warm-up instantiates the continuous-task primitive from Section 2 using the communication and shared-failure models already defined in Sections 2.3 and 2.4.
+The only additional assumptions are that leaf estimates are (approximately) unbiased and admit a compute-dependent MSE description.
+Let $Y\in[0,1]$ denote the latent target and let each leaf output an estimate $X\in[0,1]$. To keep the warm-up focused on organization rather than estimator bias, we assume conditional unbiasedness,
+$\mathbb{E}[X\mid Y]=Y.$ (8)
+Write the residual E:=XYE:=X-Y, so $\mathbb{E}[E\mid Y]=0$, and summarize leaf quality by the MSE
+$v_{0}(x):=\mathbb{E}\big[(X-Y)^{2}\big]=\mathbb{E}[E^{2}],$ (9)
+which depends on the per-leaf compute allocation $x$. In the operating range where we compare scale up versus scale out, we use the scaling approximation
 $v_{0}(x)\approx c\,x^{-\beta},$ (10)
-with constants $c>0$ and $\beta>0$, as introduced in Section 2.2.
-Communication distortion.
-An estimate sent across an edge as a message of length $m$ is received with additive distortion:
-$\widetilde{X}=X+\eta,\qquad\mathbb{E}[\eta]=0,\quad\mathrm{Var}(\eta)=\sigma_{c}^{2}(m),$ (4)
-where $\sigma_{c}^{2}(m)$ decreases with $m$.
-We treat the channel noise variables $\eta$ as independent across edges and independent of $Y$.
-This is a compact abstraction of imperfect summarization, constrained prompts, and the general fact that receivers do not reconstruct sender information perfectly.
-Shared errors as residual correlation.
-Within any aggregation group, child residuals are not assumed independent. Instead, we use an equal-correlation model: for children $i\neq j$ with residuals $E_{i},E_{j}$,
+with constants $c>0$ and $\beta>0$ (Section 2.2).
+Communication follows the additive distortion model from Section 2.3, with message length $m$; we treat edge distortions as independent across edges and independent of $Y$.
+Within any aggregation group, residuals are modeled by the equal-correlation structure from Section 2.4: for children $i\neq j$,
 $\mathrm{Var}(E_{i})=v,\qquad\mathrm{Cov}(E_{i},E_{j})=\rho v,$ (11)
-where $\rho\in[0,1)$ is the correlation coefficient.
-This matches the intended interpretation of $\rho$ as a measurable “shared-failure” strength: $\rho=0$ corresponds to independent residuals, while $\rho$ near 1 indicates strong groupthink.
-Aggregation by averaging.
-An internal node with fan-in $b$ receives $\widetilde{X}_{1},\dots,\widetilde{X}_{b}$ and outputs the average
+where $\rho\in[0,1)$ is the shared-failure correlation parameter.
+Finally, an internal node with fan-in $b$ outputs the average of its received inputs,
 $\widehat{Y}=\frac{1}{b}\sum_{i=1}^{b}\widetilde{X}_{i}.$ (12)
-Under the symmetry assumptions above, this is the MSE-optimal linear unbiased aggregator.
+Under these symmetry assumptions, averaging is the MSE-optimal linear unbiased aggregator.
 3.2 Closed-form risk for star, chain, and tree
@@
 This observation will reappear in the binary setting as a saturation effect: once the system has reached its fixed point, additional scale should be spent on changing the communication or correlation regime rather than adding depth.
-depth LLMSE $v_{L}$$v^{\star}(b,m,\rho)$$v^{\star}(b,m^{\prime},\rho)$shorter messageslonger messagesexponential shrink until saturation
-Figure 3: Conceptual behavior of the tree recursion (9).
-Increasing depth reduces error rapidly at first (by repeatedly averaging), but improvement saturates at a floor $v^{\star}$ determined by communication distortion and shared correlation.
-Longer messages (larger $m$) reduce $\sigma_{c}^{2}(m)$, lowering the floor, but also reduce feasible fan-in due to the context constraint $bm\leq W$.
+Figure 4: Continuous averaging recursion through a $b$-ary hierarchy with equal-correlation $\rho$.
+The MSE $v_{L}$ contracts approximately exponentially with depth $L$ until it reaches the communication-limited floor $v^{\star}(b,m,\rho)=\sigma_{c}^{2}(m)/\bigl((b-1)(1-\rho)\bigr)$.
+Longer messages reduce $\sigma_{c}^{2}(m)$ and therefore lower the attainable floor.
 3.4 Scale-out versus scale-up under a fixed budget
@@
 Substituting $v_{0}(x)\approx cx^{-\beta}$ and $N\approx B/(x+c_{0})$ into (21) yields
-$v_{L}\;\approx\;c\,x^{-\beta}\left(\frac{B}{x+c_{0}}\right)^{-t}=c\,B^{-t}\,x^{-\beta}(x+c_{0})^{t}.$
+$v_{L}\;\approx\;c\,x^{-\beta}\left(\frac{B}{x+c_{0}}\right)^{-t}=c\,B^{-t}\,x^{-\beta}(x+c_{0})^{t}.$ (22)
 For a fixed budget $B$, minimizing this expression over $x$ reduces to minimizing $x^{-\beta}(x+c_{0})^{t}$. A short calculus argument shows a qualitative threshold:
@@
 The binary case is qualitatively different.
-Averaging is linear, so errors evolve smoothly toward an explicit floor.
-Majority vote discards magnitude and retains only a sign; when the incoming signal is too weak relative to noise and correlation, repeated majority aggregation can wash out information rather than denoise it.
-Section 4 formalizes this amplification–collapse transition and shows how the same measurable quantities ($\rho$, message length $m$, fan-in $b$, and budget allocation) determine whether organization helps or hurts on success/failure tasks.
+Because majority vote is nonlinear, it can fundamentally change how errors propagate under repeated aggregation.
+Section 4 analyzes this effect and shows that hierarchical organization exhibits a sharp amplification–collapse transition on success/failure tasks.
 4 Binary Core: Majority Aggregation and Phase Transitions
-Binary success/failure tasks behave differently from the continuous warm-up because majority vote discards magnitude information and retains only a sign.
-That nonlinearity can either amplify weak signal across layers or wash it out entirely.
-As a result, hierarchical organization exhibits a sharp tipping point: below a threshold, deeper trees collapse to chance; above it, depth improves accuracy until it saturates.
-Real multi-agent pipelines often mix graded intermediate signals (scores, confidences, rankings) with hard decisions (accept/reject, pass/fail).
-We therefore view the continuous and binary analyses as two canonical primitives that appear within richer protocols; the same measurable constraints (communication fidelity, shared failures, and context-limited fan-in) govern both.
+We begin by characterizing how one majority layer transforms weak bias, then incorporate correlated errors and lossy communication to obtain a phase transition and budget-relevant exponents for deep trees.
 4.1 Majority maps and correlated aggregation
-From leaf votes to bias.
-[…]
-$\mu:=\mathbb{E}[\widehat{Y}Y],\qquad\Pr(\widehat{Y}=Y)=\frac{1+\mu}{2}.$
-In a tree, each internal node receives $b$ child votes and outputs an aggregated vote. To analyze such systems, we need to understand how bias transforms through one aggregation step.
+The central object in the binary analysis is the one-step aggregation map that describes how an internal node transforms the biases of its child votes into an output bias.
 Majority aggregation map.
@@
 $\mathrm{Maj}(V_{1},\dots,V_{b}):=\mathrm{sign}\!\Big(\sum_{i=1}^{b}V_{i}\Big)\in\{-1,+1\},$
-where ties are impossible because $b$ is odd. The induced majority map is the bias at the output:
+where ties are impossible because $b$ is odd; even $b$ can be handled by specifying a tie-breaking rule (e.g., random) with only minor notational changes.
+For the same reason, we will assume an odd $b$ in our future sections.
+The induced majority map is the bias at the output:
 $f_{b}(u)\;:=\;\mathbb{E}\big[\mathrm{Maj}(V_{1},\dots,V_{b})\,Y\big]=2\,\Pr\!\Big(\mathrm{Bin}(b,p)\geq\tfrac{b+1}{2}\Big)-1,\qquad p=\frac{1+u}{2}.$ (24)
@@
 Figure 5 visualizes $f_{b}(u)$ for several fan-ins, highlighting its concavity and the increasing (but sublinear) small-signal gain $f_{b}^{\prime}(0)$.
-input bias uuoutput bias $f_{b}(u)$f=uf=ub=3b=3b=5b=5b=7b=7b=11b=11
-Slope at origin: $f_{b}^{\prime}(0)=\frac{b\binom{b-1}{(b-1)/2}}{2^{b-1}}$    (grows as $O(\sqrt{b})$)
-Figure 4: The majority aggregation map $f_{b}(u)$ for odd fan-in $b\in\{3,5,7,11\}$.
-Each curve shows how input bias $u$ transforms into output bias after majority voting.
-All curves pass through $(0,0)$ and $(1,1)$, with slope $f_{b}^{\prime}(0)>1$ at the origin—majority amplifies weak signals.
-Larger fan-in produces stronger amplification, but the gain grows only as $O(\sqrt{b})$.
+Figure 5: Majority amplification map $f_{b}(u)$ for odd fan-in $b$ (shown for $b\in\{3,5,7,11\}$), where $u$ is the input bias and $f_{b}(u)=2\Pr(\mathrm{Bin}(b,(1+u)/2)\geq(b+1)/2)-1$.
+The diagonal $f(u)=u$ is shown for reference.
+The slope at the origin, $f_{b}^{\prime}(0)=\frac{b\binom{b-1}{(b-1)/2}}{2^{\,b-1}}$, grows on the order of $\sqrt{b}$, capturing how majority increasingly amplifies weak signals as fan-in grows.
 A tractable shared-failure model.
@@
 The key design message is already visible in (27): increasing $\rho$ moves $f_{b,\rho}$ closer to the identity map, weakening amplification and making deep organization harder.
-[…]
 Figure 6: The $\rho$-shared correlation model (Definition 2).
@@
 Moreover, $\mu^{\star}=1$ if $\gamma=1$, and $\mu^{\star}<1$ if $\gamma<1$.
-We assume odd fan-in $b$ to avoid ties in majority voting; even $b$ can be handled by specifying a tie-breaking rule (e.g., random) with only minor notational changes.
 Figure 7 illustrates the two regimes: when $\alpha_{\rho}<1$ the recursion map lies below the diagonal and $\mu_{t}$ collapses to 0, while when $\alpha_{\rho}>1$ the map crosses the diagonal at a stable fixed point $\mu^{\star}$, yielding amplification followed by saturation.
@@
 This matches the practitioner experience that groupthink and aggressive compression are the two most common reasons multi-agent scaling fails.
-$\mu_{t}$$T(\mu_{t})$$T(\mu)=\mu$$\mu^{\star}$$\alpha_{\rho}<1$$\alpha_{\rho}>1$
-(a) Recursion map (schematic)
-depth $t\mu_{t}$$\mu^{\star}$collapseamplify then saturate
-(b) Trajectories (schematic)
-Figure 6: A phase transition in deep majority-vote trees.
-Left: the recursion map $T(\mu)=f_{b,\rho}(\gamma\mu)$ compared to the diagonal.
-If $\alpha_{\rho}=T^{\prime}(0)<1$, the map stays below the diagonal and bias decays to 0.
-If $\alpha_{\rho}>1$, the map is initially above the diagonal, creating amplification, but boundedness and communication loss yield a stable fixed point $\mu^{\star}$.
-Right: corresponding trajectories from the same initialization.
+(a) Recursion map $T(\mu)$ vs. the diagonal.
+(b) Iterates $\mu_{t+1}=T(\mu_{t})$ from the same initialization.
+Figure 7: Binary phase transition under the recursion map $T(\mu)=f_{b,\rho}(\gamma\mu)$ where $f_{b,\rho}(u)=\rho u+(1-\rho)f_{b}(u)$.
+In the subcritical regime ($\alpha_{\rho}<1$) trajectories collapse to $\mu=0$.
+In the supercritical regime ($\alpha_{\rho}>1$) a stable fixed point $\mu^{\star}>0$ emerges and trajectories amplify then saturate at $\mu^{\star}$.
 4.3 Small-signal amplification and the organization exponent $s$
@@
 Theorem 5 (Small-signal amplification band)
-Fix $b,\rho,\gamma$ and let $T(u)=f_{b,\rho}(\gamma u)$ with $\alpha_{\rho}=T^{\prime}(0)$.
+Fix odd $b\geq 3$, $\rho\in[0,1)$, and channel reliability $\gamma=\gamma(m)\in(0,1]$.
+Let $T(u)=f_{b,\rho}(\gamma u)$ and $\alpha_{\rho}:=T^{\prime}(0)$.
 For any $\eta\in(0,1)$, there exists $\delta=\delta(\eta;b,\rho,\gamma)\in(0,1]$ such that for all $u\in[0,\delta]$,
@@
 For a full $b$-ary tree of depth $L$, the number of leaves is $N=b^{L}$. Theorem 5 implies that in the growth regime (before saturation),
-$\mu_{L}\approx\mu_{0}\,\alpha_{\rho}^{L}=\mu_{0}\,N^{s},\qquad s:=\frac{\log\alpha_{\rho}}{\log b}.$
+$\mu_{L}\approx\mu_{0}\,\alpha_{\rho}^{L}=\mu_{0}\,N^{s},\qquad s:=\frac{\log\alpha_{\rho}}{\log b}.$ (34)
 We call $s$ the organization exponent. It converts the per-layer gain $\alpha_{\rho}$ into a statement about how performance scales with the number of leaves under hierarchical aggregation.
 This exponent is the main interface between organization and budget.
-Later, in Section 5, we compare $s$ to the single-agent scaling exponent $\beta$ to decide whether scale-out can beat scale-up under a fixed budget.
+Later, in Section 4.5, we compare $s$ to the single-agent scaling exponent $\beta$ to decide whether scale-out can beat scale-up under a fixed budget.
 At this stage, the key point is conceptual: $s$ increases with better communication ($\gamma(m)$), decreases with shared failures ($\rho$), and is limited by fan-in through $f_{b}^{\prime}(0)$.
@@
 Proof sketch.
+Write the recursion as $\mu_{t+1}=T(\mu_{t})$ with $T(u)=f_{b,\rho}(\gamma u)$ and a stable fixed point $\mu^{\star}$ when $\alpha_{\rho}=T^{\prime}(0)>1$.
 There are two phases.
-First, while $\mu_{t}$ is small, the map grows approximately linearly with factor $\alpha_{\rho}>1$, so reaching a constant-size neighborhood of $\mu^{\star}$ takes $O(\log(1/\mu_{0}))$ steps.
-Second, concavity implies $T^{\prime}(\mu)\leq\kappa<1$ on a neighborhood below $\mu^{\star}$, yielding a contraction argument that reaches relative error $\varepsilon$ in $O(\log(1/\varepsilon))$ additional steps.
-An explicit construction and constants appear in Appendix E.
+First, while $\mu_{t}$ remains in the small-signal band, Theorem 5 gives $(1-\eta)\alpha_{\rho}^{t}\mu_{0}\leq\mu_{t}\leq\alpha_{\rho}^{t}\mu_{0}$, so reaching a constant-sized neighborhood of $\mu^{\star}$ takes $O(\log(1/\mu_{0}))$ steps.
+Second, stability implies $T^{\prime}(\mu^{\star})<1$; by continuity there exists a neighborhood below $\mu^{\star}$ on which $T^{\prime}(\mu)\leq\kappa<1$, so the recursion is a contraction and the error shrinks geometrically, reaching relative error $\varepsilon$ in $O(\log(1/\varepsilon))$ additional steps.
+Appendix E provides an explicit construction and constants.
 $\square$
@@
 In the subcritical regime, adding depth is actively harmful; the only viable route to scale-out is to move the system across the threshold $\alpha_{\rho}>1$ by improving $m$, reducing $\rho$, or increasing fan-in $b$ within the context constraint.
-Section 5 uses the organization exponent $s$, the threshold $\alpha_{\rho}>1$, and the clipped objective (27) to derive topology and budget phase diagrams under context limits.
+Section 4.5 uses the organization exponent $s$, the threshold $\alpha_{\rho}>1$, and the clipped objective (35) to derive topology and budget phase diagrams under context limits.
 Section 5 turns these results into a theory-guided design algorithm that outputs monotone communication design curves and compute allocations under a fixed budget.
-5 Topology and Budget Phase Diagrams
-Sections 34 characterized how information transforms through a single aggregation layer.
-This section turns those layer-wise results into system-level guidance: when centralized aggregation (star) is feasible, when hierarchy (tree) is necessary, and how these choices interact with a fixed budget $B$ under a context window constraint $W$.
-The resulting phase diagrams organize a large design space into regimes governed by the environment parameters of Section 2 (the single-agent scaling exponent $\beta$, shared-failure correlation $\rho$, and communication fidelity $\gamma(m)$) together with the context constraint $W$.
-We focus on the binary setting and use the recursion and organization exponent developed in Section 4; when helpful, we connect back to the continuous warm-up to build intuition.
-5.1 Star as an upper bound without constraints
+4.5 Topology and Budget Phase Diagrams
+Building on the binary recursion and organization exponent from Section 4, we translate the layer-wise results into system-level guidance under a fixed budget $B$ and context window $W$.
+We ask when a context-limited star is feasible, when hierarchy is necessary, and when scaling out can beat scaling up before saturation.
+The resulting regimes can be summarized as phase diagrams in the measurable environment parameters $\beta$, $\rho$, and $\gamma(m)$ (Section 2), together with the context constraint $W$.
+4.5.1 Star as an upper bound without constraints
 A recurring empirical pattern is that hierarchies sometimes outperform naive centralization.
@@
 Once $W$ binds, a star can only incorporate a bounded number of leaves regardless of budget, and that is precisely where hierarchy becomes a meaningful design choice.
-5.2 Why chains degrade without new evidence
+4.5.2 Why chains degrade without new evidence
 Chains are appealing in practice because they resemble deliberation: one agent’s output becomes the next agent’s input.
@@
 Design takeaway.
-Absent new evidence, chains are not a mechanism for synergy.
-If a task is inherently sequential, chains may still be necessary, but then one must model the arrival of new information explicitly.
-If intermediate agents inject new information or perform error-correcting verification, the effective bias can increase; Proposition 8 isolates the relay-only communication effect.
-Our phase diagrams therefore focus on star versus tree, which represent the two competing strategies for aggregating many leaf signals under context constraints.
-5.3 Context-limited star versus hierarchical trees
+A relay-only chain strictly attenuates signal (Proposition 8); it helps only when intermediate steps add new evidence (tools, retrieval, verification) rather than passing along the same information.
+We therefore focus on star versus tree as the two scalable evidence-aggregation strategies under context limits.
+4.5.3 Context-limited star versus hierarchical trees
 The context window $W$ turns Proposition 7 from a theoretical upper bound into a practical obstacle.
@@
 Figure 8 illustrates this trade-off: the star saturates early due to context limits, while the tree can continue scaling until it reaches its fixed-point ceiling $\mu^{\star}$.
-[…]
-Figure 7: Schematic performance comparison under a fixed context window $W$.
-Star (orange) aggregates NN leaves but saturates when NN hits the ceiling $N_{\max}=\lfloor W/m\rfloor$.
-Tree (blue) distributes the constraint locally ($bm\leq W$), enabling $N=b^{L}$ leaves.
-When $s>\beta$, trees overtake at budget $B_{\mathrm{crit}}$ and continue with slope $s$ until reaching $\mu^{\star}$.
+Figure 8: Star vs.
+tree vs.
+single-agent performance under a fixed context window $W$.
+A star can increase fan-in NN until it hits the context ceiling $N_{\max}=\lfloor W/m\rfloor$ (defining $B_{\mathrm{sat}}$), after which further budget can only improve per-leaf compute (scaling like $B^{\beta}$).
+A hierarchical tree can continue increasing the effective number of leaves by adding depth, exhibiting a growth exponent $s=\log(\alpha_{\rho})/\log b$ in the growth regime; when $s>\beta$, trees eventually dominate beyond $B_{\mathrm{crit}}$.
 This benefit is real only if information survives the hierarchy. The binary phase transition from Section 4.2 provides the feasibility condition: a deep tree can only maintain nontrivial accuracy when
@@
 Even when $\alpha_{\rho}>1$, deeper trees saturate at a fixed point $\mu^{\star}$, so trees should be viewed as a tool for scaling up to a regime of strong performance, not for unlimited improvement.
-[…]
-Figure 8: A schematic binary phase diagram for hierarchical trees at fixed fan-in $b$ and single-agent scaling exponent $\beta$.
-The lower boundary $\alpha_{\rho}=1$ (Theorem 4) separates regimes where deep hierarchies collapse to chance from regimes where they amplify accuracy.
-Within the amplifying regime, the stronger condition $\alpha_{\rho}>b^{\beta}$ (equivalently $s>\beta$) marks where scaling out can beat scaling up in the budgeted growth regime.
-Both boundaries move upward as $\rho$ increases.
-Since $\gamma(m)\leq 1$, the $\alpha_{\rho}=b^{\beta}$ boundary may exit the feasible range at large $\rho$; the schematic clips it at $\gamma=1$, truncating the synergy region.
+Figure 9: Phase regions in $(\rho,\gamma)$ for fixed fan-in $b$ and single-agent scaling exponent $\beta$.
+The boundary $\alpha_{\rho}=1$ separates collapse ($\alpha_{\rho}\leq 1$) from amplification.
+The boundary $\alpha_{\rho}=b^{\beta}$ separates mere amplification (no synergy, $s\leq\beta$) from synergy ($s>\beta$), where hierarchical scale-out can outpace single-agent scaling.
+Design levers such as increasing message length $m$ (raising $\gamma$) or reducing shared-failure correlation $\rho$ move systems toward the synergy region.
 Figure 9 summarizes the logic in a single picture.
@@
 The next subsection turns (39) into an explicit budget threshold.
-5.4 Budget thresholds for synergy: conditions on $s$ and $\beta$
-We now connect topology to budget.
-The question is not whether a deep tree can amplify weak votes in isolation (Section 4), but whether it can do so more efficiently than simply making a single agent stronger.
-This is where the comparison between the organization exponent $s$ and the single-agent scaling exponent $\beta$ becomes decisive.
+4.5.4 Budget thresholds for synergy: conditions on $s$ and $\beta$
+The layer-wise analysis of Section 4 yields an organization exponent $s$ that governs how signal grows with depth before saturation.
+Here we translate that picture into a fixed-budget comparison: under the same total budget $B$, when can scaling out with a tree outperform scaling up a single agent?
 Setup (growth regime).
-We work in the small-signal regime where the single-agent scaling law is well approximated by
-$\mu_{0}(x)=g(x)\approx kx^{\beta},$
-and where the tree has not yet saturated (so Theorem 5 applies). For a full $b$-ary tree of depth $L$ with $N=b^{L}$ leaves, the small-signal prediction gives
-$\mu_{L}\approx\mu_{0}(x)\,\alpha_{\rho}^{L}=\mu_{0}(x)\,N^{s},\qquad s=\frac{\log\alpha_{\rho}}{\log b}.$ (32)
-Under the budget model of Section 2.5, the total cost is approximately
-$B\approx N\big(x+c_{0}(b,m)\big),\qquad c_{0}(b,m)=2m\frac{b}{b-1},$
-so $N\approx B/(x+c_{0})$. Substituting into (32) yields the design objective
-$\mu_{\mathrm{grow}}(B;b,m,x)\;\approx\;k\,x^{\beta}\left(\frac{B}{x+c_{0}(b,m)}\right)^{s}.$ (33)
-This objective is interpretable: $x^{\beta}$ captures scale-up, and $(B/(x+c_{0}))^{s}$ captures scale-out through the number of leaves.
+Assume the per-leaf scaling curve admits a local exponent $\beta$ in the operating range (Section 2.2), so that $\mu_{0}(x)=g(x)\approx kx^{\beta}$ when leaf bias is small.
+In the amplification band of Theorem 5 (i.e., before the recursion approaches $\mu^{\star}$), a depth-$L$ full $b$-ary tree satisfies
+$\mu_{L}\;\approx\;\mu_{0}(x)\,\alpha_{\rho}(b,m)^{L}\;=\;\mu_{0}(x)\,N^{s(b,m,\rho)},\qquad N=b^{L},$ (40)
+where $\alpha_{\rho}(b,m)$ is defined in (29) and $s(b,m,\rho)=\log\alpha_{\rho}(b,m)/\log b$ as in (34).
+With the tree budget model $B\approx N\big(x+c_{0}(b,m)\big)$ from Section 2.5, we have $N\approx B/(x+c_{0})$, giving the growth surrogate
+$\mu_{\mathrm{grow}}(B;b,m,x)\;\approx\;k\,x^{\beta}\left(\frac{B}{x+c_{0}(b,m)}\right)^{s(b,m,\rho)}.$ (41)
+We use (41) only as a pre-saturation diagnostic; in Section 4.4 we clip predictions at the fixed point $\mu^{\star}(b,m)$.
 Theorem 9 (Closed-form compute allocation in the growth regime)
-Fix $(b,m,\rho)$ and suppose $s(b,m,\rho)>\beta$. Then the growth objective (33) is maximized at
-$x^{\star}\;=\;\frac{\beta}{s-\beta}\,c_{0}(b,m),$ (34)
+Fix $(b,m,\rho)$ and suppose $s(b,m,\rho)>\beta$. Then the objective (41) is maximized at
+$x^{\star}(b,m)=\frac{\beta}{s(b,m,\rho)-\beta}\,c_{0}(b,m),$ (42)
 and the resulting optimized growth prediction scales as
-$\mu_{\mathrm{grow}}^{\star}(B;b,m)\;\approx\;k\,\kappa(s,\beta)\,c_{0}(b,m)^{\beta-s}\,B^{s},\qquad\kappa(s,\beta):=\beta^{\beta}(s-\beta)^{\,s-\beta}s^{-s}.$ (35)
-If $s\leq\beta$, the objective (33) is maximized by pushing $x$ as large as allowed (scale up rather than scale out).
+$\mu_{\mathrm{grow}}^{\star}(B;b,m)\;\approx\;k\,\kappa\big(s(b,m,\rho),\beta\big)\,c_{0}(b,m)^{\beta-s(b,m,\rho)}\,B^{s(b,m,\rho)},$ (43)
+where $\kappa(s,\beta)=\beta^{\beta}(s-\beta)^{s-\beta}s^{-s}$. If $s(b,m,\rho)\leq\beta$, then (41) is (weakly) increasing in $x$, so the growth surrogate favors scale up over scale out.
 Proof sketch.
-For $s>\beta$, maximize $\log\mu_{\mathrm{grow}}=\beta\log x+s\log B-s\log(x+c_{0})$ over $x>0$.
-Setting the derivative to zero gives (34).
-Substituting back yields (35).
-If $s\leq\beta$, the derivative is nonnegative for all $x$, so the objective increases with $x$.
-Details are routine and given in Appendix D.
+The optimization is one-dimensional: maximize $\log\mu_{\mathrm{grow}}=\beta\log x-s\log(x+c_{0})+s\log B$.
+Setting the derivative to zero yields (42), and substituting gives (43).
+This is the same calculus as in the continuous warm-up (Section 3.4), with the continuous exponent $t$ replaced by the binary exponent $s$; Appendix D.2 gives the full derivation.
 $\square$
 Synergy against a single agent.
-A single-agent baseline that spends the full budget on compute achieves (in the same small-signal regime)
-$\mu_{\mathrm{single}}(B)\approx k\,B^{\beta}.$ (36)
-Comparing (35) and (36) yields a simple conclusion: in the growth regime, a hierarchy can only outpace the single agent when $s>\beta$.
-Moreover, when $s>\beta$, there is an explicit budget threshold beyond which the optimized tree dominates the single-agent scaling curve.
+Spending the whole budget on a single agent yields
+$\mu_{\mathrm{single}}(B)\;=\;\mu_{0}(B)\;\approx\;k\,B^{\beta}.$ (44)
+Comparing (43) with (44) shows that exponent-level scale-out synergy in the growth regime requires $s>\beta$ (equivalently $\alpha_{\rho}>b^{\beta}$).
 Corollary 10 (A budget threshold for scale-out synergy (growth regime))
-Fix $(b,m,\rho)$ with $s>\beta$, and define $\kappa(s,\beta)$ as in (35).
-Ignoring saturation, a sufficient condition for the optimized tree prediction $\mu_{\mathrm{grow}}^{\star}(B;b,m)$ to exceed the single-agent prediction $\mu_{\mathrm{single}}(B)$ is
-$B\;\geq\;B_{\mathrm{crit}}(b,m)\;:=\;c_{0}(b,m)\,\kappa(s,\beta)^{-\frac{1}{s-\beta}}.$ (37)
+Fix $(b,m,\rho)$ with $s(b,m,\rho)>\beta$. If
+$B\;\geq\;B_{\mathrm{crit}}(b,m)\;\approx\;c_{0}(b,m)\,\kappa\big(s(b,m,\rho),\beta\big)^{-1/(s(b,m,\rho)-\beta)},$ (45)
+then the optimized growth prediction (43) exceeds the single-agent scaling ($4$).
 Interpretation.
-Equation (37) is a concrete, testable statement.
-It predicts that scale-out synergy is harder when coordination is expensive (large $c_{0}$), when the organization exponent barely beats the scaling exponent (small $s-\beta$), or when the system is close to the amplification boundary (small $s$).
-In Section 4.4 we introduced clipped objectives to account for saturation; when saturation matters, the threshold still indicates whether there exists a budget range in which scale-out should help before the fixed point dominates.
-[…]
-Figure 9: A schematic budget phase diagram under a fixed context window $W$.
-A star cannot incorporate more than $N_{\max}(m)$ leaves, so beyond a point its gains come primarily from making leaves stronger (scale-up), leading to a $B^{\beta}$-type growth until task-level saturation.
-A supercritical tree with organization exponent $s>\beta$ can exhibit a steeper $B^{s}$ growth regime before saturating at $\mu^{\star}$.
-The budget threshold $B_{\mathrm{crit}}$ marks where optimized scale-out can overtake single-agent scaling in the growth regime.
-5.5 A universal upper bound on amplification exponents
+Equation (45) makes the trade-off explicit: synergy is harder when coordination is expensive (large $c_{0}$) or when $s$ only barely exceeds $\beta$.
+Saturation can truncate the growth regime, but the threshold still indicates whether a budget window exists in which scale-out should help before the fixed point $\mu^{\star}$ dominates.
+4.5.5 A universal upper bound on amplification exponents
 The organization exponent $s$ governs how quickly a majority-vote hierarchy can amplify weak signal as the number of leaves grows.
@@
 $\square$
-Implications for synergy.
-The bound $s\leq 1/2$ has two immediate consequences.
-First, it explains why scale-out is fragile on some tasks.
-If the single-agent scaling exponent satisfies $\beta\geq 1/2$, then the inequality $s>\beta$ is impossible under majority-vote hierarchies with one-bit messages.
-In that regime, a designer should not expect budget-exponent gains from adding agents; improvement must come from making agents stronger, from richer communication than a single bit, or from protocols that effectively reduce shared failures.
-Second, the bound clarifies the role of diversity.
-Even if $\beta<1/2$, a large $\rho$ reduces $\alpha_{\rho}$ and pushes $s$ downward, shrinking or eliminating the synergy region in Figure 8.
-This is precisely the regime in which practitioners observe that “more agents does not help,” and it suggests a concrete remedy: invest budget not in more of the same agents, but in mechanisms that reduce shared error modes, which in our framework means reducing the effective $\rho$.
-Summary of the phase-diagram logic.
-Taken together, the results above reduce topology choice to a few quantitative checks.
-Under context constraints, deep trees are only viable in the supercritical regime $\alpha_{\rho}>1$; exponent-level budgeted synergy in the growth range requires $s>\beta$ (equivalently $\alpha_{\rho}>b^{\beta}$); and within the majority-vote tree family we always have the universal limit $s\leq 1/2$.
-Even when these conditions hold, depth helps only up to saturation: once the recursion is near $\mu^{\star}$, additional budget is typically better spent improving communication fidelity (larger $\gamma(m)$) or reducing shared failures (smaller $\rho$) than on additional depth.
-Section 6 turns these checks into a compact design diagnostic within the model.
+Phase-diagram summary and implications.
+The preceding results collapse topology choice into a small set of quantitative checks.
+Under context constraints, deep trees are only viable in the supercritical regime $\alpha_{\rho}(b,m)>1$ (otherwise bias collapses to chance as depth grows).
+In the growth regime, budgeted synergy against scale-up further requires $s(b,m,\rho)>\beta$, equivalently $\alpha_{\rho}>b^{\beta}$ (Section 4.5.4).
+Within one-bit majority-vote hierarchies we also have the universal cap $s\leq 1/2$ (Theorem 11), which explains why scale-out is fragile on some tasks: if $\beta\geq 1/2$, exponent-level synergy is impossible without richer communication, heterogeneity that lowers the effective $\rho$, or protocols that change the effective aggregation map.
+Even when $s>\beta$, gains persist only until saturation at $\mu^{\star}(b,m)$; beyond that point, additional budget is typically better spent improving fidelity (larger $m$, hence larger $\gamma(m)$) or reducing shared failures (smaller $\rho$) than on additional depth.
+Figure 9 summarizes these regimes, and Section 5 turns the checks into concrete design diagnostics within the model.
 5 Design Diagnostics Within the Model
-The results so far characterize when different organizations can help under a fixed budget: a context-limited star saturates (Section 5.3), chains compound loss (Section 5.2), and trees are viable only in the supercritical regime $\alpha_{\rho}>1$ with small-signal growth governed by an organization exponent $s$ (Section 4).
+The results so far characterize when different organizations can help under a fixed budget: a context-limited star saturates (Section 4.5.3), chains compound loss (Section 4.5.2), and trees are viable only in the supercritical regime $\alpha_{\rho}>1$ with small-signal growth governed by an organization exponent $s$ (Section 4).
 Here we translate these results into design diagnostics within the model: given a budget $B$, a context window $W$, and a small set of effective environment quantities, we ask whether hierarchy is feasible, whether scale-out can outpace scale-up before saturation, and which constraint is binding when it cannot.
@@
 We focus on binary success/failure tasks, where hierarchies can either amplify or collapse; the continuous setting admits analogous (and smoother) diagnostics via the closed-form recursions of Section 3.
-6.1 Closed-form compute allocation: optimal $x^{\star}$ in the growth regime
-Fix a candidate tree architecture specified by a fan-in $b$ and message length $m$, with local feasibility constraint $bm\leq W$.
-Let $\rho$ be the calibrated shared-failure correlation and $\gamma(m)$ be the calibrated communication reliability.
-Define the per-layer weak-signal gain
-$\alpha_{\rho}(b,m):=\gamma(m)\big(\rho+(1-\rho)f_{b}^{\prime}(0)\big),\qquad s(b,m):=\frac{\log\alpha_{\rho}(b,m)}{\log b}\quad\text{(for fixed calibrated }\rho\text{)}.$
-Theorem 4 implies that deep hierarchies are only viable if $\alpha_{\rho}(b,m)>1$.
-Assume we are in this supercritical regime and consider budgets for which the system remains in the growth regime before saturation.
-Under the small-signal scaling approximation $g(x)\approx kx^{\beta}$ and the tree budget model $B\approx N(x+c_{0}(b,m))$ with $c_{0}(b,m)=2m\frac{b}{b-1}$, Section 5.4 derived the growth objective
-$\mu_{\mathrm{grow}}(B;b,m,x)\;\approx\;k\,x^{\beta}\left(\frac{B}{x+c_{0}(b,m)}\right)^{s(b,m)}.$ (40)
-The next fact is the inner workhorse of the design procedure: once $(b,m)$ are fixed, the optimal compute per leaf has a closed form.
-Proposition 12 (Growth-regime compute allocation)
-Fix $(b,m)$ with $s(b,m)>\beta$. Then the objective (40) is maximized at
-$x^{\star}(b,m)=\frac{\beta}{s(b,m)-\beta}\,c_{0}(b,m),$ (41)
-and the resulting optimized growth prediction scales as
-$\mu_{\mathrm{grow}}^{\star}(B;b,m)\;\approx\;k\,\kappa\big(s(b,m),\beta\big)\,c_{0}(b,m)^{\beta-s(b,m)}\,B^{s(b,m)},$ (42)
-where $\kappa(s,\beta)=\beta^{\beta}(s-\beta)^{s-\beta}s^{-s}$. If $s(b,m)\leq\beta$, then (40) is (weakly) increasing in $x$, so the growth approximation favors scale-up over scale-out.
-Proposition 12 is a restatement of Theorem 9 in design form.
-It makes the trade-off explicit: when coordination is expensive (large $c_{0}$), each leaf should be stronger (larger $x^{\star}$); when organization is efficient (large $s-\beta$), each leaf can be weaker, freeing budget to add more leaves.
-In practice, given $B$, the recommended number of leaves and depth are
-$N^{\star}(B;b,m)\;\approx\;\left\lfloor\frac{B}{x^{\star}(b,m)+c_{0}(b,m)}\right\rfloor,\qquad L^{\star}(B;b,m)\;\approx\;\left\lfloor\log_{b}N^{\star}(B;b,m)\right\rfloor.$ (43)
-Depth $L^{\star}$ should not exceed the mixing depth needed to approach saturation (Theorem 6); see Section 6.4.
+5.1 Compute allocation in the growth regime
+Fix a feasible fan-in $b$ and message length $m$ with $bm\leq W$.
+Let $\alpha_{\rho}(b,m)$ be defined as in (29), and write the associated organization exponent as $s(b,m):=\log\alpha_{\rho}(b,m)/\log b$.
+In the supercritical regime $\alpha_{\rho}(b,m)>1$ and before saturation, the binary small-signal prediction takes the growth form (41).
+Theorem 9 yields a closed-form compute allocation $x^{\star}(b,m)$ (eq.
+(42)) and an optimized growth curve $\mu_{\mathrm{grow}}^{\star}(B;b,m)$ (eq.
+(43)) whenever $s(b,m)>\beta$.
+If $s(b,m)\leq\beta$, the growth surrogate favors spending budget to scale up rather than scale out.
+Given a budget $B$, we then allocate
+$N^{\star}(B;b,m)\approx\left\lfloor\frac{B}{x^{\star}(b,m)+c_{0}(b,m)}\right\rfloor,\qquad L^{\star}(B;b,m)\approx\left\lfloor\log_{b}N^{\star}(B;b,m)\right\rfloor,$ (48)
+and cap $L^{\star}$ using the mixing-depth guarantee (Theorem 6) when operating near saturation.
 5.2 Monotone communication design: why $m^{\star}(B)$ increases with budget
@@
 $\mathcal{M}_{b}:=\{m\in\mathbb{Z}_{+}:\ bm\leq W\}.$
-For each $m\in\mathcal{M}_{b}$, define the growth-optimal prediction from Proposition 12,
+For each $m\in\mathcal{M}_{b}$, define the growth-optimal prediction from Theorem 9,
 $\underline{\mu}_{b}(m;B):=\mu_{\mathrm{grow}}^{\star}(B;b,m)=k\,\kappa(s_{b,m},\beta)\,c_{0}(b,m)^{\beta-s_{b,m}}\,B^{s_{b,m}},\qquad s_{b,m}:=s(b,m).$ (49)
@@
 A full proof, stated in terms of increasing differences and upper envelopes, is given in Appendix I.
-$t=\log B$$\log\underline{\mu}_{b}(m;B)$upper envelopemm smallmm mediummm large
-(a) Envelope over $\log B$
-$t=\log B$$m_{b}^{\star}(B)$$m_{1}$$m_{2}$$m_{3}$
-(b) Monotone design curve
-Figure 10: Why the optimal message length increases with budget in the growth regime (fixed fan-in $b$).
-After optimizing compute allocation $x$, each candidate message length $m$ induces a line $\log\underline{\mu}_{b}(m;B)=s_{b,m}\log B+q_{b}(m)$ with slope $s_{b,m}$ increasing in $m$.
-The maximum over $m$ is the upper envelope of these lines (left), implying that an optimal $m_{b}^{\star}(B)$ can be chosen as a nondecreasing, piecewise-constant function of $B$ (right).
+(a) Upper envelope over candidate message lengths.
+(b) Resulting $m_{b}^{\star}(B)$ (monotone step function).
+Figure 10: Upper-envelope argument behind the monotone message-length design curve (fixed fan-in $b$).
+For each feasible message length $m$ (numbers shown are tokens in this example), we first optimize the per-leaf compute allocation $x$ in the growth-regime surrogate, yielding a closed-form prediction of the form $\log\underline{\mu}_{b}(m;B)=s_{b,m}\,\log B+q_{b}(m)$, i.e., a line in $t=\log B$ whose slope $s_{b,m}$ increases with $m$ because communication fidelity $\gamma(m)$ improves with longer messages.
+(a) We plot these candidate lines for a representative subset of feasible $m$ values and highlight their pointwise maximum (thick curve), i.e., the upper envelope.
+Red markers indicate budgets where two candidates tie (line intersections), and the vertical dashed lines project these switch points onto the $t$-axis.
+(b) The maximizing message length $m_{b}^{\star}(B)$ is therefore a nondecreasing step function of $B$, jumping exactly at the envelope breakpoints.
+Algorithm 1 constructs the full envelope efficiently by scanning candidates in increasing $m$ (increasing slope) and maintaining only the active lines and their intersection thresholds.
 What the theorem does and does not claim.
@@
 The output is a model-based prediction and a set of diagnostics (which constraint is binding), rather than a claim that the abstraction optimizes real agent stacks end-to-end.
+Figure 11 summarizes the workflow of design-diagnostics.
+Inputs / measured quantities
+budget $B$, context $W$, single-agent exponent $\beta$ (or $g(\cdot)$), fidelity curve $\gamma(m)$, shared-failure correlation $\rho$
+Feasible design grid
+choose odd fan-in $b$ and message length $m$ subject to $bm\leq W$
+Amplification diagnostics
+compute $\alpha_{\rho}(b,m)$ (29), compute $s(b,m,\rho)$ (34), check $\alpha_{\rho}>1$ and $s>\beta$
+Budget split and depth
+compute $x^{\star}$ (42), compute $N^{\star},L^{\star}$ (48), cap depth using Theorem 6
+Predict and compare
+growth prediction $\mu^{\star}_{\mathrm{grow}}$ (43), clip at saturation $\mu^{\star}$ (Sec. 4.4), compare to star and single baselines
+Outputs
+recommended $(b,m,x^{\star},L^{\star})$, predicted regime: growth vs. saturation diagnostic: which constraint binds
+Figure 11: A design-diagnostics pipeline within the model.
+Measured constraints ($B,W$) and calibrated parameters ($\beta$, $\gamma(m)$, $\rho$) determine which $(b,m)$ are feasible, whether depth amplifies ($\alpha_{\rho}>1$), and whether scale-out can beat scale-up in the growth regime ($s>\beta$).
+The closed-form allocation $x^{\star}$ and implied $(N^{\star},L^{\star})$ yield a predicted performance that is clipped at the saturation fixed point $\mu^{\star}$ and compared against star and single-agent baselines.
 Compare against centralization.
@@
 For exponent-level budgeted synergy in the growth regime, also discard $s(b,m)\leq\beta$.
-For remaining candidates, set the per-leaf compute to the closed form $x^{\star}(b,m)$ in (41) (Proposition 12) and choose the implied leaf count and depth via (43), with depth interpreted through the mixing-depth guarantee of Theorem 6.
+For remaining candidates, set the per-leaf compute to the closed form $x^{\star}(b,m)$ in (42) (Theorem 9) and choose the implied leaf count and depth via (48), with depth interpreted through the mixing-depth guarantee of Theorem 6.
 For fixed $b$, Algorithm 1 yields the monotone budget-indexed communication curve $m_{b}^{\star}(B)$, so the outer search can be performed efficiently.
@@
 Synthetic experiments provide a controlled setting where the ground-truth parameters are known and the phase boundary can be probed directly.
-Our goal here is not to optimize absolute performance, but to verify that the qualitative prediction—amplification versus collapse governed by $\alpha_{\rho}$—is recovered under the generative model analyzed in Section 4.
+Our goal here is not to optimize absolute performance, but to verify that the qualitative prediction, i.e., amplification versus collapse governed by $\alpha_{\rho}$, is recovered under the generative model analyzed in Section 4.
 We include one lightweight Monte Carlo study under the $\rho$-shared binary model.
@@
 This minimalism makes the regime boundaries transparent, but it also defines what the theory does and does not claim.
-8.1 Relationship to prior results
+While our results provide a rigorous asymptotic foundation for agent scaling, the practical utility of these bounds hinges on the empirical estimation of the shared-failure correlation $\rho$ and the organization exponent $s$.
+In practice, $\rho$ serves as a ‘homogeneity metric’ for a given model ensemble; a high $\rho$ suggests that the agents are inheriting the same inductive biases or training data flaws, thereby collapsing the effective fan-in $b$.
+Future work should focus on developing diagnostic probes—short, high-entropy test batteries—to calibrate these parameters for specific LLM families.
+By doing so, the Phase Transition $\alpha_{\rho}$ can be transformed from a theoretical threshold into a real-time ‘stopping rule’ for determining whether further tree depth will yield diminishing returns or catastrophic signal decay.
+Relationship to prior results
 Our binary phase transition echoes reconstruction thresholds for broadcasting on trees (Kesten and Stigum, 1966; Evans et al., 2000; Mossel, 2001), where information can or cannot propagate from leaves to the root depending on a single scalar.
@@
 Together, these results suggest that sharp thresholds, whether in inter-agent communication, intra-agent skill selection, or hierarchical aggregation, may be a recurring motif in bounded-resource AI systems.
-8.2 What the framework explains (and what it does not)
+What the framework explains (and what it does not)
 The framework is designed to explain qualitative regime changes that are otherwise hard to diagnose: why stars saturate under context limits, why chains compound loss, why trees can either amplify or collapse depending on $\alpha_{\rho}$, and why correlated failures impose floors that make large ensembles behave like “one agent with shared blind spots.” It also clarifies what additional sophistication must accomplish: any useful complexity must effectively increase communication fidelity, reduce shared failures, or relax the fan-in constraint.
@@
 Parameters such as $\rho$ and $\gamma(m)$ are effective summaries that can vary with depth, prompts, tools, and task difficulty; real communication is far richer than a one-bit abstraction; and practical budget accounting can differ across implementations.
-For these reasons, we view the theory as a coordination diagnostic—predicting which constraint is binding and which direction should help—rather than as a source of precise point predictions.
+For these reasons, we view the theory as a coordination diagnostic, i.e., predicting which constraint is binding and which direction should help, rather than as a source of precise point predictions.
 Crucially, the alignment between our derived bottlenecks and independent large-scale observations (Kim et al., 2025) suggests that the effective parameters defined here capture the structural essence of the scaling problem, rendering the theory robust to changes in the underlying base models.
-8.3 How stronger protocols fit the framework
+How stronger protocols fit the framework
 Protocols such as debate, critique, verification, or self-consistency can be interpreted as mechanisms that change the effective parameters of the model (Irving et al., 2018; Christiano et al., 2018; Du et al., 2024; Wang et al., 2023).
@@
 The theory suggests a concrete way to evaluate such protocols: measure how much they move the effective bottlenecks per unit budget, and check whether they move the system across the supercritical boundary $\alpha_{\rho}>1$ or the synergy boundary $s>\beta$.
-8.4 Extensions beyond our scope
+Extensions beyond our scope
 Several extensions are natural.
@@
 Acknowledgments and Disclosure of Funding
-The authors gratefully acknowledge funding from Canada CIFAR AI Chair Program and the Canada NSERC Discovery Grant (RGPIN-2021-03115).
+The authors gratefully acknowledge funding from Canada CIFAR AI Chair Program and the Canada NSERC Discovery Grant program.
 Proof Roadmap and Dependency Graph
@@
 E.2 Clipped objectives and approximation guarantees
-The clipped proxy $\widehat{\mu}_{L}=\min\{\mu^{\star},\mu_{0}\alpha_{\rho}^{L}\}$ in (27) is most naturally interpreted for the regime $\mu_{0}\leq\mu^{\star}$, which includes the small-signal settings used in Section 5.4.
+The clipped proxy $\widehat{\mu}_{L}=\min\{\mu^{\star},\mu_{0}\alpha_{\rho}^{L}\}$ in (35) is most naturally interpreted for the regime $\mu_{0}\leq\mu^{\star}$, which includes the small-signal settings used in Section 4.5.4.
 In that case, monotonicity and Theorem 4 imply $\mu_{L}\leq\mu^{\star}$ for all $L$, while Lemma 16 implies $\mu_{L}\leq\mu_{0}\alpha_{\rho}^{L}$.
@@
 F.4 Tree vs. star dominance beyond a threshold
-The comparison between a context-limited star and a supercritical tree in Section 5 is driven by two facts: (i) the star cannot use more than $N_{\max}(m)$ leaves for fixed $m$, hence eventually must spend additional budget on scale-up (which grows like $B^{\beta}$ in the small-signal regime), while (ii) a tree with exponent $s>\beta$ admits a growth regime with $B^{s}$ scaling (Appendix D) before saturating at $\mu^{\star}$ (Appendix E).
+The comparison between a context-limited star and a supercritical tree in Section 4.5 is driven by two facts: (i) the star cannot use more than $N_{\max}(m)$ leaves for fixed $m$, hence eventually must spend additional budget on scale-up (which grows like $B^{\beta}$ in the small-signal regime), while (ii) a tree with exponent $s>\beta$ admits a growth regime with $B^{s}$ scaling (Appendix D) before saturating at $\mu^{\star}$ (Appendix E).
 This yields a provable intermediate budget range where trees can dominate under context constraints, formalized through the sufficient condition $B\geq B_{\mathrm{crit}}$ in Corollary 10.
@@
 These templates are meant to support qualitative diagnostics: whether context is binding, whether measured correlation suggests limited gains, and whether the calibrated $\widehat{\alpha}_{\rho}$ places a hierarchy in the amplifying or collapsing regime.
-In that sense, the phase diagrams and design curves from Sections 56 provide concrete hypotheses about which organization should help under matched budgets and which lever (communication fidelity versus shared-failure reduction) is most likely to matter when it does not.
+In that sense, the phase diagrams and design curves from Sections 4.55 provide concrete hypotheses about which organization should help under matched budgets and which lever (communication fidelity versus shared-failure reduction) is most likely to matter when it does not.
 Establishing quantitative predictiveness in realistic deployments, where prompts, tools, and dependence structure may drift with depth and task difficulty, is an important direction for future work.